A Complete Bibliography of Publications in Biometrics: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 October 2023 Version 1.07

Title word cross-reference

$119.95 [518]. $96 [257]. F [28]. G [91]. l_{2} [658]. N [58]. P [644, 499, 575]. t [55].

-estimation [91]. -hacking [499]. -mixture [58]. -splines [644]. -test [55, 28]. -type [658]. -value [575].

0 [518].

1 [252]. 10.1201 [425]. 10.1201/978 [425]. 10.1201/978-1-003-20346-9 [425]. 121.00 [38]. 130.00 [260]. 131 [299]. 19 [291, 295, 242, 686, 688, 687, 365, 411, 660].

2015 [77]. 2017 [375]. 2019 [228, 117, 229, 40]. 2020 [155]. 2021 [422, 424, 465]. 2022 [611, 466, 612]. 268 [379]. 276 [296]. 293 [340]. 2nd [706, 297].

330 [339]. 390 [341]. 3rd [613, 154].

426 [298]. 457 [297]. 460 [338].

523 [300]. 592. [380].

608 [337].

7 [707, 467]. 71 [77]. 711 [259]. 73 [375]. 75 [228, 229]. 79.96 [6].

9 [425, 227]. 978 [707, 467, 518, 227]. 978-0-367-73432-9 [227]. 978-1-003-20346-9 [425]. 978-1-118-94889-7 [378]. 978-1-118-94890-3 [378]. 978-1-119-23857-7 [467]. 978-1-138-04898-0 [518]. 978-1-138-10640-6 [380]. 978-3-030-62403-3 [377]. 978-3-030-62404-0 [377]. 978-3-030-73791-7 [707]. 978-3-030-97678-1 [705]. 978-3-030-97679-8 [705]. 99.99 [258].

absence [608, 85]. absolute [415]. abundance [31, 251]. abundances [415]. Accelerated [487]. accommodating [372]. Accounting [502, 192, 615]. accuracy [608, 16, 110, 128, 97, 550]. accurate [575]. Acknowledgments [156, 301, 468]. activities [147]. activity [644]. acyclic [210]. Adachi [297]. ADAPT [691]. ADAPT-R [691]. Adaptive [180, 475, 23, 474, 472, 471, 476, 633, 473, 515, 481, 54, 324, 565, 516, 17, 189, 33, 443, 308, 599, 644, 504, 20, 47, 448, 201, 600, 24, 453, 232]. additive [389, 619, 483, 317, 539]. Addressing [401]. adjacent [446]. adjusted [229, 135, 253, 4]. Adjusting [653, 52]. adjustment [291, 295, 485, 604, 294, 477, 292, 373, 293]. adolescents [225]. adult [149]. adults [225]. aerial [325]. after [385, 383, 384, 382, 386]. AG [376, 466, 118]. against [686]. Age [687, 623]. age-at-onset [623]. Age-related [687]. ages [32]. aggregate [249, 646]. aggregated [364]. agriculture [40]. Ahmad [466]. al [115, 116]. Albert [338]. Alejandro [341]. alert [652]. Alex [294]. algorithm [446]. algorithm-assisted [446]. Algorithms [259, 254, 250]. Alicia [611]. alive [627]. Allaire [39]. allele [348]. allele-based [348]. allocation [23, 671]. allowance [103]. already [411]. alteration [286]. alternating [50]. alternative [599, 26]. Alzheimer [192, 344, 126, 698, 99, 125, 672]. analyses [43, 512, 223]. analysi [139, 116]. Analysis [244, 435, 259, 316, 1, 71, 228, 703, 151, 207, 389, 187, 9, 28, 417, 674, 517, 5, 38, 145, 206, 368, 175, 594, 567, 598, 447, 40, 239, 240, 220, 256, 285, 136, 566, 272, 653, 636, 675, 593, 640, 408, 493, 11, 196, 299, 632, 121, 355, 587, 495, 502, 363, 307, 115, 421, 370, 140, 506, 407, 184, 409, 616, 691, 360, 138, 36, 339, 628, 351, 454, 358, 361, 52, 406, 29, 246, 654, 107, 440, 241, 87, 219, 679, 498, 543, 109, 50, 48, 268, 209, 425, 365, 387, 445, 434, 633, 581, 541, 72, 10, 423, 404, 410, 188, 451]. analysis [585, 492, 534, 26, 44, 405, 457, 595, 664, 238, 536, 152, 137, 380, 466, 297, 261, 340]. analytically [398]. analytics [378]. analyze [147]. Analyzing [166, 357, 649, 605, 174]. Anastasios [117, 385, 383, 384]. anatomic [185]. ANCOVA [114, 113]. Andrew [80]. Anh [424]. animal [191]. animals [100]. Ann [153]. annual [237]. Anthony [67, 66, 68, 69]. antidepressants [276]. antipsychotic [216]. Application [391, 361, 99, 322, 218, 582, 197, 9, 235, 229, 446, 198, 329, 683, 388, 447, 239, 352, 272, 463, 250, 283, 353, 196, 344, 46, 237, 532, 596, 644, 442, 698, 543, 365, 445, 600, 169, 570, 649, 4, 97, 284, 252]. Applications [394, 441, 28, 424, 549, 16, 273, 227, 347, 15, 379, 125, 660, 371, 705]. approach [703, 281, 270, 648, 16, 432, 427, 663, 220, 15, 397, 233, 495, 502, 370, 429, 652, 215, 431, 428, 430, 364, 603, 509, 278, 387, 546, 288, 578, 188, 554, 22, 354, 12, 419, 174, 337]. approaches [691]. Approval [162, 165, 163, 164]. Approximate [231, 86]. Approximation [43, 450]. architecture [683]. area [676, 452]. areas [331]. armed [24]. arthritis [32, 693]. artificial [256, 155]. ascent [418]. Ashis [518]. Ashkan [527, 530, 529]. Assessing [597, 62, 61, 63, 689, 59, 64, 60, 274, 319, 607, 466]. assessment [229, 327, 477, 160, 159, 161, 4]. assignment [691, 503]. assisted [446, 247, 96]. associated [147, 190, 83, 108, 672]. Associating [700]. association [320, 402, 513, 399, 170, 140, 610, 341, 72, 348, 457, 550, 71]. Assuming [310, 249]. assumption [22]. Aston [261]. asymmetric [584]. asymptomatic [687]. Asymptotic [360]. Asynchronous [637]. attributes [394]. augmented [142, 658]. Augustin [429, 431, 428, 430]. Aurore [422]. Automated [679]. automatic [441]. autopolyploid [657]. auxiliary [485, 345, 85]. average [9, 88, 198, 17, 362, 459]. averaging [285, 650, 488].

B [228, 117, 160]. Babette [465]. background [686]. Baker [152]. balance [319, 673]. balancing [673]. bands [287]. Banerjee [518]. Barah [517]. Bartlett [114]. barycenter [552]. based [181, 228, 55, 562, 701, 80, 5, 229, 438, 368, 648, 242, 659, 668, 565, 444, 79, 81, 162, 165, 485, 285, 576, 221, 226, 398, 363, 370, 140, 652, 360, 308, 163, 620, 351, 454, 164, 366, 20, 29, 452, 624, 610, 96, 167, 72, 202, 183, 544, 103, 4, 128, 459, 348, 666, 564, 297]. baseline [308]. basis [16]. Basu [518]. Batch [217, 253, 702]. batch-effect [253]. batch-mark [702]. Bayes [188, 535, 630]. Bayesian [80, 337, 81, 218, 629, 311, 389, 187, 173, 417, 368, 594, 683, 685, 79, 352, 327, 216, 309, 328, 704, 15, 397, 101, 217, 233, 200, 196, 632, 682, 19, 443, 168, 34, 532, 511, 215, 510, 172, 83, 631, 308, 596, 127, 126, 73, 628, 538, 442, 82, 369, 231, 219, 684, 633, 600, 288, 486, 534, 533, 232, 234, 354, 595, 536, 601, 537, 143, 376, 339, 6]. Beat [67, 66, 68, 69]. before [411]. Behavior [425]. behaviour [326]. Ben [408, 407, 409, 406]. benefit [95, 599, 96, 274]. Benkeser [294]. Berlin [299]. beta [394]. beta-diversity [394]. Beth [80]. between [311, 513, 11, 394, 135, 205]. between-subject [394]. beyond [339]. Bhattacharyya [517]. bi [284]. bi-level [284]. Bias [57, 1, 320, 43, 607, 653, 115, 370, 499, 512, 116]. Biased [336, 335, 133, 242, 32, 191]. Biclustering [37]. bidimensionally [13]. big [387]. Binacox [441]. binary [291, 295, 294, 121, 292, 361, 452, 57, 25, 545, 288, 293, 630]. Binbing [379]. binning [186]. binomial [573, 670]. bio [162, 165, 163]. bio-creep [162, 165, 163]. bioconductor [341]. biocreep [164]. bioimaging [11]. biomarker [475, 474, 477, 443, 95, 472, 603, 471, 476, 96, 314, 473, 108, 274]. biomarker-assisted [96]. biomarker-treatment [314]. biomarkers [308, 182, 313]. Biometry [298]. biomonitoring [434]. biosimilars [73]. biostatistics [612, 613, 377]. birth [177, 290]. birth-and-death [177]. birthweight [366]. bisulfite [199]. Bivariate [452, 343, 392, 135, 87, 282, 72, 371]. blockwise [418]. Boca [337, 517, 155, 117, 38, 260, 380, 422, 41, 259, 298, 40, 338, 612, 613, 300, 296, 339, 379, 6, 153, 258, 425, 341, 257, 261]. bone [276]. Book [340, 611, 337, 519, 517, 155, 117, 38, 260, 380, 422, 707, 706, 41, 424, 259, 376, 298, 466, 40, 338, 193, 612, 613, 467, 227, 378, 39, 300, 299, 226, 705, 297, 154, 296, 118, 339, 379, 465, 6, 153, 518, 258, 425, 341, 257, 377, 261, 423]. Boosting [669]. bootstrap [444, 652]. boreal [442]. borrow [478]. Botev [300]. both [320]. bound [181]. bounded [57]. bounds [76]. Brain [286, 272, 200, 29, 195, 679, 681, 453, 672]. breast [147, 333]. Brian [160, 6]. bridge [580]. Bridging [459, 26]. broaden [189]. Broemeling [339]. Brumback [465]. Building [12]. Bukhari [466]. bulk [107]. burden [225]. Burgess [706].

C [228, 586]. CACE [152, 151]. Caceres [341]. Caiado [153]. calcium [594]. calculation [104, 575]. calculations [684]. calibrated [233, 603, 312]. calibrating [449]. calibration [123, 317, 209]. can [673]. cancer [147, 289, 243, 683, 461, 352, 285, 227, 333, 34, 532, 454, 253, 680, 697]. candidate [697]. Cao [424]. capture [181, 325, 45, 251, 112, 395, 111]. capture-recapture [45, 251, 112, 395, 111]. care [106]. Carlo [324]. Carone [160]. Carriquiry [611]. Carson [226]. CASANOVA [490]. Case [167, 589, 105, 648, 176, 170, 591, 500, 631, 590, 588, 592, 123, 317, 231, 44, 405, 643]. case-cohort [643]. case-control [648, 176, 170, 500, 123, 317, 167, 44, 405]. case-crossover [589, 105, 591, 590, 588, 592, 231]. categorical [219]. Causal [311, 264, 266, 263, 267, 514, 265, 465, 359, 589, 703, 523, 515, 173, 706, 522, 663, 566, 516, 305, 189, 494, 398, 662, 307, 591, 524, 172, 590, 20, 588, 592, 82, 555, 545, 521, 525, 223, 210, 564]. causality [519]. causally [567]. CD4 [252]. CDN$ [38]. CEDAR [674]. Ceja [425]. cell [417, 556, 551, 552, 609, 107, 371, 252]. cellular [417]. Censored [133, 85, 389, 245, 621, 388, 617, 316, 647, 494, 391, 279, 440, 624, 610, 698, 539, 561, 278]. censoring [187, 372, 242, 414, 27, 626, 87, 315, 312]. censorings [143]. Center [658]. Center-augmented [658]. cerebral [679]. chain [420]. chained [281]. Cham [376, 705, 118]. Chan [466]. Chang [155]. Change [640]. Change-plane [640]. changepoint [250]. changes [412, 600]. Chapman [337, 519, 155, 117, 38, 260, 380, 422, 706, 298, 338, 300, 296, 339, 379, 465, 6, 518, 258, 341, 257]. characteristic [287]. characteristics [140]. characterizing [683, 216]. Chasan [612]. Chasan-Taber [612]. Chen [228, 139, 138, 137]. Chester [338]. Child [290]. childhood [329, 98]. China [242]. choice [129]. Choleksky [14]. Chollet [39]. choosing [413]. Christopher [154]. Chu [139, 138, 137]. Chuan [139, 138, 137]. circular [323, 221]. City [611]. class [16, 256, 316, 495, 275, 288]. classes [513]. Classification [148, 412, 289, 378, 457, 560, 153]. classifiers [313]. clinical [481, 229, 690, 23, 220, 460, 639, 227, 653, 480, 478, 505, 700, 506, 215, 360, 73, 504, 518, 561, 96, 253, 201, 314, 680, 24, 4, 232]. clonal [185]. Closed [570, 331, 112, 111]. clumping [269]. Cluster [236, 204, 229, 245, 104, 271, 647, 121, 436, 355, 46, 665, 689, 184, 366, 585, 4, 459]. cluster-adjusted [229, 4]. cluster-based [366]. cluster-randomized [245, 459]. clustered [207, 389, 621, 357, 316, 367, 540, 97]. Clustering [558, 153, 701, 403, 239, 37, 186, 172, 442, 288, 544]. clusters [130]. co [121]. co-primary [121]. coefficient [347, 102, 434]. coefficients [539]. Coherent [545]. cohort [249, 242, 462, 123, 643]. Colin [41]. collaborative [20]. collected [596]. colocalization [11]. combination [460, 397, 575]. combine [387]. Combining [249, 645, 21, 546]. commensurate [537]. Comment [115, 116]. common [56, 268]. communication [674]. communities [678]. community [47]. comparability [249]. comparable [699]. comparative [23]. Comparing [686, 55, 684]. comparison [276]. comparisons [331]. compartmental [83]. compendium [376]. Competing [336, 513, 392, 620, 335, 618, 232]. complex [533, 349]. complex-valued [533]. complier [494]. component [55, 235, 414, 593, 408, 407, 409, 406, 268, 581, 404, 410]. component-wise [414]. components [442]. composite [275]. composition [551]. Compositional [246, 396, 587, 234]. compound [578]. computation [171]. computational [466]. Concave [668]. concentrations [652]. concerning [229, 4]. Concordance [617]. concurrent [149]. conditional [245, 321, 108, 12]. conditionally [452]. conditions [412]. confidence [322, 573, 287, 48, 282, 202, 124, 518]. confirmatory [220]. confounders [175, 52]. confounding [368, 132, 432, 427, 566, 429, 431, 428, 430]. connectivity [286, 126, 29, 195]. connectome [453]. Consequences [310, 55]. considerations [482, 690]. Consistent [66, 678, 67, 388, 65, 70, 68, 69]. constrained [205]. constraint [144, 96]. constraints [171]. Constructing [124]. construction [556, 210]. contaminated [648, 509]. contamination [167]. Continuous [582, 475, 333, 474, 640, 149, 477, 443, 484, 326, 472, 112, 509, 471, 476, 448, 651, 473, 585, 111]. continuous-time [112, 111]. Contrasting [638]. control [180, 648, 176, 505, 170, 363, 500, 123, 317, 167, 44, 405]. controls [179]. conversions [86]. convolutions [533]. Cook [38]. Copas [370]. Copas-model [370]. coprimary [585]. copula [669, 460, 542, 233, 577, 609, 624, 72]. copula-type [460]. copy [250]. Cormack [100]. correct [175, 304]. corrected [573, 283, 509]. Correcting [688]. Correction [77, 228, 229, 375, 1, 115, 57, 116]. Corrections [105]. correlated [609, 574, 630]. correlation [343, 576, 395, 371]. correspondence [667]. corridors [48]. cost [330]. cost-effectiveness [330]. count [204, 141, 9, 606, 699, 90, 356, 371]. counties [225]. counts [270, 450, 169, 252]. coupled [417]. course [9]. covariance [388, 456, 14, 208, 268, 578, 284, 664]. Covariate [248, 477, 291, 295, 319, 604, 189, 294, 135, 292, 442, 439, 201, 209, 356, 633, 51, 673, 293, 664, 667, 199]. covariate-adaptive [201]. covariate-adjusted [135]. covariate-dependent [633, 664]. Covariate-driven [248]. covariate-induced [439]. covariate-informed [442]. covariates [133, 30, 621, 677, 333, 491, 251, 391, 205, 509, 651, 501, 213]. cover [38, 260, 259, 118, 6, 258, 257, 261]. coverage [548]. COVID [466, 294, 339, 292, 293, 291, 295, 242, 686, 688, 687, 507, 365, 411, 660]. COVID-19 [466, 294, 339, 292, 293, 291, 295, 242, 686, 688, 687, 365, 411, 660]. COVID-19-induced [507]. Cox [441, 625, 418, 346, 255, 53, 278, 51]. CRC [337, 517, 155, 38, 260, 380, 422, 41, 424, 259, 298, 40, 338, 612, 613, 227, 300, 296, 339, 379, 465, 6, 258, 425, 257, 261, 117, 153, 341]. creep [162, 165, 163]. criteria [228, 5, 413]. critical [510]. CRM [149]. Cross [235, 550, 104, 623, 391, 208, 584]. Cross-component [235]. cross-covariance [208]. cross-sectional [104, 623, 391]. Cross-trait [550]. cross-validation [584]. crossover [589, 105, 591, 638, 89, 590, 588, 592, 231]. cumulative [420, 52, 615]. Cun [260]. Cun-Hui [260]. cure [277, 345, 87, 379]. current [212, 35]. curve [331, 676, 667]. curves [235, 287, 56, 654]. cut [441]. cut-point [441]. cutoff [313]. cutting [566]. cutting-edge [566]. Cynthia [518].

D [160, 339, 542]. D-vine [542]. daily [701, 411]. damage [693]. Daniel [294]. Dash [466]. Data [55, 260, 259, 20, 1, 572, 94, 218, 21, 228, 211, 141, 389, 31, 563, 197, 187, 9, 235, 173, 212, 337, 517, 5, 145, 244, 249, 222, 206, 198, 646, 586, 41, 368, 621, 424, 594, 280, 598, 323, 556, 142, 388, 343, 298, 513, 466, 40, 338, 23, 248, 239, 399, 352, 433, 240, 328, 617, 542, 692, 272, 463, 622, 357, 250, 636, 186, 15, 413, 478, 593, 101, 283, 606, 14, 196, 316, 699, 226, 344, 637, 505, 494, 552, 297, 251, 170, 394, 558, 665, 462, 115, 609, 670, 34, 540, 638, 506, 215, 392, 185, 279, 596, 191, 628]. data [358, 603, 47, 246, 90, 107, 440, 624, 698, 539, 241, 518, 584, 53, 437, 57, 543, 109, 50, 116, 48, 278, 365, 356, 387, 464, 680, 681, 541, 72, 642, 618, 261, 290, 411, 188, 554, 22, 574, 166, 359, 371, 544, 390, 350, 315, 213, 559, 649, 97, 284, 457, 354, 664, 605, 702, 630, 252, 238, 536, 449, 537, 236, 174, 666, 38, 300, 299]. Data-adaptive [20]. datasets [685]. David [294]. Davidian [117, 385, 383, 384]. dealing [582]. death [177, 414, 514, 620]. Debiased [501]. decision [704, 480, 578]. decisions [257]. decline [411]. Decomposition [577, 141]. decompositions [14]. Deep [39]. definite [353]. degree [195]. Delaigle [422]. delayed [105, 33, 90, 688]. Delivering [699]. demonstration [254]. dense [453]. densities [328]. Density [323, 276]. dependence [693, 135, 575]. dependent [621, 346, 366, 509, 87, 439, 633, 694, 51, 664]. depression [15]. derived [199]. Design [598, 444, 690, 26, 589, 519, 639, 227, 18, 217, 33, 443, 591, 590, 73, 110, 588, 592, 317, 253, 232, 459, 601]. Design-based [444]. design-With [227]. Designing [74]. designs [105, 402, 482, 490, 176, 475, 474, 130, 89, 308, 472, 471, 476, 25, 473, 643]. detect [314]. detection [412, 441, 28, 586, 329, 594, 250, 455, 286, 640, 655, 221, 532, 150, 496, 213]. detections [395]. Determination [365, 537]. Developing [212]. development [218, 155, 401, 467]. device [162, 165, 163, 164, 166]. Dhruba [517]. diabetes [254]. diagnostic [608, 16, 508, 128]. diagnostics [21]. dialysis [102]. Diaz [294]. dietary [229, 4]. difference [527, 530, 573, 655, 529, 528, 526, 531]. difference-in-differences [527, 530, 529, 528, 526, 531]. differences [481, 527, 530, 529, 528, 526, 531]. different [185]. differential [120, 286, 128]. diffusion [216, 353]. digital [325]. Dimension [616, 403, 677, 458]. dimensional [441, 707, 132, 240, 593, 355, 558, 307, 670, 82, 246, 555, 680, 488, 544, 390, 559, 560, 12]. dimensions [311, 557, 99]. direct [675, 484, 178, 419]. directed [210]. direction [147]. directional [179]. Dirichlet [214, 671, 688]. Dirk [300]. Discovering [98]. discovery [179, 453, 535]. discrete [518]. discretely [177]. discretized [509]. discriminant [358]. Discussion [589, 523, 527, 67, 80, 139, 264, 62, 530, 475, 385, 522, 266, 383, 384, 474, 66, 529, 408, 294, 61, 63, 160, 591, 429, 407, 409, 138, 524, 431, 68, 163, 590, 472, 292, 428, 164, 69, 430, 406, 60, 265, 528, 473, 293, 137, 81, 476]. Discussions [165, 432, 525, 410]. disease [21, 270, 192, 401, 699, 46, 506, 126, 698, 99, 125, 508, 453, 108, 451, 605, 672, 344, 193]. disease-associated [108]. disease-related [453]. diseases [339, 365]. disorders [361]. Distance [29, 388, 221]. Distance-based [29, 221]. distances [183]. distributed [674, 3, 375, 694, 24]. distribution [271, 623, 439]. Distributional [408, 407, 409, 406, 404, 669, 410]. distributions [142, 93, 445, 128]. dive [338]. diverging [501]. diverse [408, 407, 409, 406, 404, 410]. diversity [394]. divide [387]. divide-and-combine [387]. DNA [445, 199]. DOI [425]. Domain [572]. dose [327, 460, 33]. dose-finding [460]. dose-schedule [33]. Double [657, 88, 143]. double-index [88]. Doubly [77, 132, 2, 555]. Doubly-robust [77, 2]. driven [248, 626]. Dropout [336, 228, 5, 335]. drug [582, 155, 460, 467, 397, 535]. drugs [216, 601]. due [105]. Dupont [429, 431, 428, 430]. duration [365]. during [126]. D’Urso [153]. dyadic [363]. Dylan [527, 530, 529]. Dynamic [176, 480, 99, 77, 562, 701, 220, 2, 371, 666, 117]. dynamically [478]. dynamics [191, 681].

early [327, 535]. early-stage [535]. easygoing [299]. eBook [705, 377, 378]. ecology [40]. economic [686]. ed [613]. Eddy [611]. edge [566]. Edited [422]. edition [706, 467, 297, 154]. Editor [151, 152]. editors [303, 8, 158, 470]. Eds [611]. EEG [681, 328]. effect [519, 269, 17, 494, 362, 363, 662, 76, 689, 253, 459, 419, 564]. Effective [66, 67, 80, 79, 81, 65, 70, 68, 69, 35]. effectiveness [330]. effects [597, 204, 139, 173, 372, 145, 88, 198, 645, 132, 91, 418, 273, 136, 333, 604, 675, 397, 101, 305, 189, 102, 398, 443, 540, 484, 138, 178, 334, 82, 610, 545, 680, 641, 486, 359, 615, 667, 199, 137]. efficacy [306, 607, 385, 120, 383, 273, 384, 382, 386]. efficiency [30, 145, 607, 505, 123, 317]. Efficient [412, 567, 273, 462, 691, 182, 641, 674, 198, 130]. eigenvalue [554]. Elastic [478, 654]. elderly [506]. electroencephalogram [681]. electronic [320, 198, 329, 648, 254, 489, 168, 696, 167, 122, 649]. Elizabeth [153]. elliptical [630]. EM-based [370]. embedding [678]. EMBRACE [370]. Emerson [163, 164]. Emiko [429, 431, 430]. Emil [423]. Empirical [228, 5, 364, 129]. Empirical-likelihood-based [228, 5]. encouragement [459]. endemic [270]. endemic-epidemic [270]. endpoint [436, 571]. endpoints [229, 245, 121, 618, 585, 4, 601]. enrichment [475, 690, 474, 308, 472, 471, 476, 473]. Enrique [425]. enrolled [504]. Ensemble [186, 560]. ensembles [131]. entry [626]. envelope [403]. environment [538, 10, 451]. environmental [298, 511, 510, 150]. epidemic [270, 242, 450, 225]. epidemics [466]. epidemiology [612]. epigenomic [412]. equation [192, 258, 210]. equations [281, 636]. equilibrium [657]. Equivalence [56]. Eric [523, 117, 522, 524]. Erica [77]. error [572, 244, 422, 105, 43, 444, 283, 462, 304, 571, 696, 509, 57, 209, 568, 569]. error-contaminated [509]. error-prone [462, 696]. errors [323, 93, 557]. errors-in-variables [557]. Ertefaie [515, 527, 530, 529, 528]. estimands [627, 638]. estimate [645, 91, 507]. Estimated [574]. estimates [123]. Estimating [94, 222, 88, 385, 383, 551, 144, 384, 225, 676, 510, 89, 448, 195, 382, 1, 109, 330, 666, 438, 446, 420, 328, 636, 115, 205, 317, 687, 116, 681, 386]. Estimation [245, 242, 14, 484, 336, 561, 582, 77, 523, 563, 562, 84, 80, 30, 198, 646, 177, 323, 461, 608, 414, 343, 79, 81, 522, 657, 433, 91, 345, 622, 655, 221, 355, 494, 251, 662, 665, 462, 619, 623, 370, 237, 391, 135, 524, 172, 596, 208, 620, 678, 20, 255, 82, 452, 439, 335, 2, 365, 580, 521, 525, 290, 578, 486, 393, 315, 615, 312, 128, 564]. estimator [444, 660]. estimators [181, 372, 565, 485, 362, 398, 171]. evaluate [601]. Evaluating [513, 604, 279, 276, 212, 304, 599]. Evaluation [203, 120]. event [291, 295, 84, 28, 463, 622, 593, 294, 436, 215, 292, 626, 50, 72, 534, 293, 174]. event-driven [626]. event-related [28, 534]. events [23, 146, 627, 135, 87]. evidence [189, 370, 296]. Exact [573]. Exact-corrected [573]. examples [298]. excess [277, 507]. exchangeability [197]. exercises [707]. expanded [416]. experimental [340]. experiments [181, 347, 675, 491, 609, 615, 459, 667]. Exploiting [189]. exposure [132, 689, 334, 509, 694]. exposure-time [689]. expression [517, 192, 551, 34, 350]. extended [702]. extension [100]. external [249, 646, 505, 364, 449]. Extracting [453].

F [611, 80, 38]. factor [147, 309, 670, 188, 169, 354, 210]. factorial [490]. factorization [248, 13]. factors [306, 190, 296]. failure [91, 487, 574, 315]. Fair [106]. fairness [549]. false [420, 179]. familywise [572, 571]. Fan [260]. Fast [685, 92]. fatal [411]. Faulkner [80]. favor [479]. Feature [390, 553, 455, 413, 355, 558, 559, 234]. features [454, 108, 12]. Feng [163, 164]. field [150, 679]. fields [208]. Fieller [282]. Fienberg [611]. filter [246]. Finding [359, 460, 215]. fine [673]. fingerprint [97]. finite [668]. finite-support [668]. Fisher [575]. fit [62, 450, 61, 63, 59, 64, 60, 624]. fitting [270]. FL [337, 155, 117, 38, 260, 380, 422, 41, 259, 298, 40, 338, 612, 613, 300, 296, 339, 379, 6, 153, 258, 341, 257, 261]. Flammini [466]. Flexible [460, 609, 224]. Florida [517]. Flory [527, 530, 529, 528]. fMRI [665, 533, 284]. forestry [298]. forests [190, 221, 19]. formal [589, 591, 590, 588, 592]. formula [421, 223]. forward [481]. forward-looking [481]. Foundations [260, 193]. fraction [87]. fractions [275]. frailty [146, 347, 316, 620, 87]. frailty-based [620]. framework [572, 21, 31, 659, 3, 375, 655, 87, 584, 492, 559]. frameworks [90]. Francesco [466]. Francis [517]. Francois [39]. Franz [299]. free [396, 202]. frequency [498]. Frequentist [650, 311]. function [497, 400, 574, 318]. function-on-function [400]. function-on-scalar [497]. Functional [483, 580, 541, 664, 572, 141, 235, 28, 593, 200, 344, 637, 237, 532, 596, 125, 48, 496, 581, 72, 236]. Functions [1, 115, 208, 224, 116, 48, 356]. Fundamentals [707, 465]. fusion [368, 285, 351, 241].

G [706, 375, 223]. g-formula [223]. gamma [360]. Gandrud [154]. Gangopadhyay [518]. Garcia [425]. Gaussian [173, 233, 321, 552, 577, 650, 351, 224, 452, 350, 534, 236]. Gauvreau [613]. GEE [54]. Gene [72, 556, 551, 538, 634, 10, 350, 517]. Gene-based [72]. gene-environment [538, 10]. General [626, 31, 659, 176, 655, 27, 421, 627, 209, 559]. generalizability [579]. Generalized [663, 500, 183, 568, 332, 28, 92, 636, 493, 670, 688, 501, 348, 575, 12]. generalized-Dirichlet-multinomial [688]. generation [242, 188]. genetic [706, 170, 610, 188, 348]. genetically [192]. genetics [441, 369]. genome [289, 140, 554, 550]. genome-wide [140, 550]. Georgia [139, 138, 137]. Geostatistical [353, 21]. Germany [299]. Ghosh [6]. Gilbert [160]. Giovanna [116]. Giovanni [62, 61, 63, 60]. Gittins [481]. given [191]. Global [16, 361]. Globaltest [570]. GmbH [299]. gold [608]. Gonzalez [341]. good [311]. goodness [62, 61, 63, 59, 64, 60, 624]. goodness-of-fit [62, 624]. Grace [422]. grant [612]. graphical [173, 244, 200, 650, 351, 108, 350, 284]. graphs [665, 210]. greater [74].

Gröbner [16]. Group [436, 214, 604, 101, 196, 326, 308, 580, 496, 128]. group-tested [128]. Grouped [636]. groups [55]. growth [235, 333]. Guangquan [337]. guarantees [548, 455]. guided [131]. Guo [408, 407, 409, 406]. Gustafson [422]. gut [463].

H [429, 431, 430]. hacking [499]. Haining [337]. Haitao [139, 138, 137]. Hall [337, 155, 38, 260, 380, 422, 298, 338, 300, 296, 339, 379, 465, 6, 518, 258, 257, 519, 117, 706, 341]. Hall/CRC [337, 155, 38, 260, 380, 422, 298, 338, 300, 296, 339, 379, 465, 6, 258, 257, 117, 341]. Handbook [380, 422, 261]. handling [320]. Hard [38, 260, 259, 118, 6, 258, 257, 261]. harm [695]. having [600]. hazard [703, 523, 277, 522, 524, 86, 209, 521, 525]. hazards [211, 244, 621, 619, 317, 509, 22, 103]. hbk [518]. health [340, 320, 198, 329, 648, 254, 489, 397, 606, 168, 696, 167, 122, 106]. healthcare [155]. Heather [613]. Heinz [67, 66, 68, 69]. Hennessy [527, 530, 529, 528]. heritability [32]. Hernan [590, 591]. Hernando [261]. heterogeneity [646, 401, 285, 458, 367, 46, 689, 547, 351, 454, 364, 107]. heterogeneous [461, 542, 665, 641, 359]. Hi [586]. Hi-C [586]. Hidden [289, 692]. Hierarchical [454, 683, 216, 360, 90, 224, 651, 605, 672]. high [629, 311, 441, 707, 608, 132, 240, 355, 558, 307, 670, 557, 82, 246, 555, 680, 488, 544, 390, 559, 560, 667]. high-dimensional [441, 707, 132, 240, 355, 558, 307, 670, 246, 680, 488, 544, 390, 559, 560]. high-throughput [667]. Highly [130, 565, 17]. Hilbert [637]. Histopathological [285, 454]. historical [478]. history [325, 38, 290]. HIV [397, 462, 252]. HIV-1 [252]. HLA [697]. HLA-I [697]. Holloway [117]. homogeneity [352, 93, 651]. homophily [363]. Hong [139, 138, 137]. Horseshoe [80, 79, 81]. Horseshoe-based [80, 79, 81]. Horvitz [221]. Hosmer [62, 61, 63, 64, 60, 59]. Hospital [704]. hospitalization [102]. HPV [607]. Huang [264, 266, 265]. Hui [260]. Human [506]. Huntington [519]. Huntington-Klein [519]. Hvitfeldt [423]. hypersphere [14]. hypoglycemia [190]. hypotheses [32, 26]. hypothesis [388]. hypothetical [638].

I/II [232]. identifiability [16]. identification [46, 662]. Identifying [306, 374, 190, 652, 108, 672]. II [232]. Iickho [705]. illness [105, 414, 620]. illness-death [414, 620]. Illustration [691]. image [239, 309]. image-on-image [309]. imaging [594, 285, 272, 15, 593, 353, 454, 369, 48]. imaging-based [285]. immune [683]. impact [201, 35, 74]. imperfect [489, 508]. implementation [41, 379]. implementations [223]. implications [310]. importance [324, 160, 109, 159, 161]. improve [289, 491, 602, 123, 317, 546, 122]. Improved [622, 360]. Improving [291, 30, 294, 579, 505, 191, 292, 293, 276, 295]. imputation [281, 321, 492]. incidence [450, 391]. incidences [686]. incomplete [638]. incorporated [350]. incorporating [281, 327, 628, 290]. Increasing [607]. incubation [242]. independence [310, 11, 47]. independent [408, 407, 409, 626, 406, 404, 410]. Index [481, 88, 511, 205, 274]. indications [73, 601]. indicators [187]. indices [617]. indirect [484, 178, 419]. individual [306, 145, 182, 109, 508, 605, 449]. individual-level [605]. individualized [197, 446, 144, 89, 483, 330, 274]. individuals [395]. induced [509, 439, 507]. infant [218]. infection [148, 83, 411]. infections [411]. infectious [339, 365, 508]. infer [107]. Inference [576, 677, 49, 504, 610, 465, 311, 515, 55, 320, 197, 9, 173, 706, 269, 685, 490, 254, 247, 176, 416, 338, 433, 663, 273, 18, 250, 675, 516, 17, 305, 699, 505, 662, 27, 514, 670, 627, 95, 602, 36, 631, 191, 366, 20, 231, 555, 618, 202, 574, 534, 97]. inferences [567]. Inferring [411, 397, 195, 554]. Infinite [692]. inflated [463, 172, 169, 371, 569]. influence [274]. Information [7, 42, 78, 119, 157, 194, 230, 262, 302, 342, 381, 426, 469, 520, 614, 350, 249, 485, 345, 489, 478, 628, 85, 364, 624, 179]. Information-incorporated [350]. informative [368, 175, 247, 184, 596, 541, 188]. informed [512, 442]. informer [535]. inhomogeneities [374]. initiation [438]. initiative [125]. inpatient [506]. institutions [322]. instrument [464]. Instrumental [523, 522, 494, 524, 521, 525, 459]. Instrumented [527, 529, 528, 526, 531, 530]. instruments [305]. Integrating [495, 680, 609, 449]. Integration [437, 424, 552, 34, 584, 238]. Integrative [631, 13, 405, 616]. intelligence [155]. intensive [433]. inter [75]. inter-unit [75]. interaction [582, 628, 10, 451, 359, 536]. interactions [205, 538, 314]. Interactive [226]. interest [611]. interference [372, 663, 27]. Interim [503]. intermittent [218, 206]. International [707, 377]. interpolator [444]. interpretable [567, 498]. interpretation [589, 591, 590, 588, 592]. intersection [332]. intersection-union [332]. interval [389, 245, 621, 573, 316, 647, 494, 391, 440, 610, 698, 282, 278, 315]. interval-censored [389, 245, 621, 316, 647, 494, 391, 440, 610, 698, 278]. intervals [322, 360, 518, 202, 124]. intervention [597, 229, 4, 564]. interventions [273]. intractable [398]. intratumor [107]. introduction [519, 297, 299, 259]. invalid [305]. Inverse [372, 53, 565, 653, 362, 140, 255, 660, 312]. inverse-probability-weighted [565]. inverse-variance [660]. Investigating [366]. involving [73]. irregularly [654, 393]. Irrizarry [259]. ISBN [380, 707, 467, 227, 378, 705, 518, 377]. Island [39]. Ismay [338]. isoform [551]. isotonic [445]. Issue [7, 42, 78, 119, 157, 194, 230, 262, 302, 342, 381, 426, 469, 520, 614]. issues [467]. Italy [582]. item [321]. Iterated [197]. iteratively [491]. Ivan [294].

J [38, 39, 6, 377]. Jackknife [488]. James [527, 80, 530, 529, 591, 590]. Jean [163, 164]. Jerald [38]. Jersey [467]. Jian [408, 407, 409, 406]. Jianqing [260]. Jianrong [227]. Jodi [294]. Johannes [707]. John [467, 378, 261]. Joint [556, 463, 344, 107, 618, 312, 457, 252, 643, 549, 146, 15, 647, 502, 693, 224, 87, 125, 315, 419, 174]. Jolly [100, 142]. Jonathan [114]. Jorge [153]. Joseph [39]. Juan [341]. Judith [611]. Jugal [517]. Juha [298]. Julia [423].

Kalita [517]. Kang [408, 407, 409, 406]. Kaplan [103]. kernel [637, 700, 356, 533]. Khan [340]. kidney [418]. Kim [424, 338]. Kim-Anh [424]. Kimberlee [613]. Kimmel [139, 138, 137]. kinetic [416]. Klein [519]. knockoff [246]. Kohei [297]. Kroese [300]. Kronthaler [299]. Kumar [517, 466].

L [228, 611, 62, 61, 63, 60]. L.L [229]. Lê [424]. Laber [117]. Labs [707]. lag [3, 375, 694]. lagged [561]. language [254]. Lappi [298]. Large [9, 62, 685, 456, 61, 63, 59, 64, 60, 390]. large-scale [456, 390]. Lasso [516, 515, 565, 17, 501]. late [680, 232]. late-onset [232]. Latent [219, 496, 702, 147, 325, 16, 513, 309, 256, 310, 671, 495, 577, 168, 670, 693, 99, 681, 288, 553]. Lauri [298]. Lawless [38]. layers [498]. Learning [109, 96, 517, 446, 162, 165, 658, 39, 300, 671, 665, 160, 163, 164, 241, 425, 641, 159, 161, 423]. Learning-based [96, 162, 165, 163, 164]. learns [491]. least [77, 258, 2, 125]. Lederer [707]. left [617, 278]. left-truncated [617, 278]. Lemeshow [62, 61, 63, 60, 62, 61, 63, 59, 60]. length [133, 242, 506, 448]. length-biased [133, 242]. lesion [532]. Letter [151, 152]. level [628, 508, 284, 605]. levels [252]. leverage [537]. Leveraging [602, 646]. Li [337, 260, 229]. life [38, 191]. lifetime [438]. like [221]. likelihood [228, 84, 5, 668, 256, 251, 619, 364, 171, 278]. likelihood-based [668]. likelihoods [148]. limited [366]. limits [213]. Linbo [523, 522, 524]. Lindquist [261]. linear [177, 396, 450, 3, 375, 635, 316, 637, 400, 500, 557, 358, 85, 507, 209, 496, 501, 12]. link [224, 630]. linked [13]. Lisa [612]. Liu [229]. load [252]. local [435]. localized [498]. location [57]. lockdowns [411]. locked [195]. log [103]. log-rank [103]. Logistic [493, 62, 480, 671, 196, 61, 63, 59, 64, 60, 583]. logistic-tree [671]. London [517, 425]. Longford [257]. Longitudinal [1, 203, 218, 228, 5, 206, 41, 175, 280, 598, 433, 240, 542, 463, 636, 604, 14, 637, 115, 168, 540, 392, 634, 20, 539, 241, 219, 543, 116, 545, 541, 393, 544, 315, 252, 419]. longitudinally [686]. look [447]. looking [481]. loss [20]. loss-based [20]. lost [91]. low [366, 679, 498]. low-field [679]. low-rank [498]. Ltd [340, 467]. Luedtke [294]. Lunardon [116]. lung [683]. Lyle [339]. lymph [333].

M [77, 228, 611, 591, 590, 518, 377, 374]. M-RA [374]. MacFarland [377]. machine [517, 162, 165, 300, 160, 163, 164, 425, 159, 161, 423]. machines [287]. Magee [80]. Maharaj [153]. making [480, 257]. malaria [148, 679]. mammogram [593]. manifolds [280]. Manning [39]. mapping [193, 46]. Marcel [376]. Marcello [613]. Marco [160]. Margaret [611]. Marginal [621, 184, 228, 270, 5, 373, 486, 312]. marginally [233]. Marie [117, 424, 385, 383, 384]. Mark [155, 702]. marked [166]. marker [304, 547, 174]. markers [203, 279, 546]. Markov [420, 692]. marriage [311]. Martin [261]. martingale [492]. Martinussen [523, 522, 524]. matched [319, 661, 276]. matching [270, 491, 97]. Mathematical [300]. matrices [353, 13, 268]. matrix [239, 272, 286, 15, 200, 14, 665, 578, 354, 297]. Matrix-based [297]. matrix-normal [354]. matrix-valued [15]. matrix-variate [286, 665]. Mattie [613]. MAW [552]. max [331]. max-t [331]. Maximum [251, 619, 84, 9, 195, 171]. MCMC [49]. mean [438, 43, 444, 91, 479, 448, 129, 48, 393, 595, 664, 318]. mean-squared [43]. means [282]. measure [46]. measurement [244, 422, 105, 283, 93, 304, 209, 568, 569]. measurements [451]. measurements-environment [451]. measures [190, 95, 603, 686, 199, 695]. Mechanisms [336, 335]. mediation [264, 266, 566, 263, 267, 307, 82, 265, 680, 332]. mediator [680, 419]. mediators [307]. medical [162, 165, 118, 163, 164, 649]. medicare [506]. Medicine [117, 155, 612, 74]. Mehmet [258]. Mehmetoglu [258]. Mehtätalo [298]. Meier [103]. memory [611]. Menardi [116]. Mendelian [43, 662, 464, 660, 44, 706]. Menon [518]. mental [397, 606]. meta [151, 139, 145, 567, 136, 653, 502, 370, 138, 36, 109, 152, 137, 380, 340]. meta-analysi [139]. meta-analysis [151, 145, 567, 136, 653, 502, 370, 138, 36, 109, 152, 137, 380, 340]. meta-regression [502]. metabolomics [570]. metastases [333]. method [206, 192, 418, 493, 149, 532, 688, 199, 672]. Methods [340, 117, 424, 379, 54, 706, 567, 227, 413, 557, 36, 304, 53, 300, 6]. methylation [199]. metric [599]. Michael [77, 62, 294, 61, 63, 60]. microbial [415]. Microbiome [671, 463, 413, 283, 394, 47, 246, 169, 234]. migration [417]. migratory [100]. Miguel [591, 590]. minimum [20]. Minin [80]. misaligned [699, 354]. misclassification [320, 201, 128, 568]. misclassified [509]. Misdiagnosis [695]. Misdiagnosis-related [695]. missing [218, 187, 206, 435, 692, 251, 602, 361, 583, 51, 649]. missingness [228, 5, 583]. misspecification [310]. mitigate [680]. mixed [101, 635, 102, 19, 577, 662, 540, 507, 568, 601, 210]. mixed-scale [19, 662]. mixOmics [424]. mixture [218, 629, 374, 403, 92, 58, 321, 552, 87, 695]. mixtures [214, 511, 510, 487, 694]. MNAR [30]. mobile [197]. modalities [408, 407, 409, 406, 404, 410]. Model [625, 544, 218, 629, 228, 147, 281, 325, 441, 133, 5, 646, 403, 549, 414, 45, 416, 401, 216, 309, 345, 328, 460, 285, 542, 622, 357, 333, 286, 480, 101, 647, 102, 344, 676, 394, 577, 650, 619, 400, 609, 370, 652, 360, 172, 596, 351, 346, 20, 317, 369, 509, 224, 452, 539, 687, 87, 561, 507, 278, 356, 445, 633, 600, 680, 488, 681, 508, 415, 202, 108, 169, 350, 284, 605, 536, 143]. Model-based [544, 652, 351]. model-free [202]. Modeling [277, 280, 392, 371, 311, 31, 192, 659, 463, 15, 101, 353, 102, 367, 168, 670, 693, 126, 487, 90, 99, 258, 545, 694, 533, 284, 354, 252, 199, 419, 174, 193]. Modelling [499, 326, 466, 135, 337]. Models [336, 181, 270, 197, 173, 212, 38, 62, 244, 269, 41, 621, 645, 142, 92, 16, 548, 450, 456, 485, 3, 375, 91, 58, 146, 418, 692, 378, 310, 347, 333, 635, 200, 98, 316, 321, 637, 552, 19, 61, 63, 650, 540, 275, 500, 511, 49, 59, 64, 557, 83, 631, 483, 191, 620, 110, 379, 373, 52, 60, 255, 395, 231, 624, 651, 219, 335, 125, 50, 209, 496, 583, 72, 288, 501, 393, 315, 615, 312, 534, 568, 702, 630, 569, 12, 643, 210, 695, 422]. modern [338]. modes [83]. modification [62, 61, 63, 59, 64, 60]. modifications [162, 165, 163, 164]. modified [14]. molecular [347, 451]. moment [270]. moments [493]. monitoring [175, 189, 503]. Monte [324]. Moodie [77]. mortality [237, 596, 507, 290, 411]. Morton [139, 138, 137]. motor [218]. movement [326]. MRI [196, 532, 679]. MRS [143]. multi [462, 168, 609, 638, 150, 24, 457, 332, 143]. multi-armed [24]. multi-environmental [150]. multi-national [462]. multi-omics [609]. multi-period [638]. multi-risks [143]. multi-SNP [332]. multi-state [168]. multi-view [457]. multiarm [514]. multiblock [248]. multicenter [360]. Multidimensional [644, 451, 193]. multidrug [109]. multidrug-resistant [109]. multigroup [71]. Multikink [543]. multilevel [102]. Multimodal [238, 536]. Multinomial [58, 94, 688, 702]. multiphenotype [513]. multiple [21, 412, 281, 331, 9, 229, 206, 567, 214, 420, 639, 692, 98, 699, 121, 511, 184, 691, 83, 279, 126, 73, 373, 437, 546, 508, 641, 618, 503, 492, 4, 405, 605, 601, 313]. Multiplicity [373]. multisite [605]. Multisource [552, 197]. multistate [207, 38]. Multivariate [442, 90, 387, 218, 629, 139, 235, 319, 598, 136, 692, 604, 14, 344, 297, 138, 36, 208, 440, 99, 219, 288, 252, 137, 424]. multiview [399, 358, 415]. Multiwave [696]. multiway [403]. mutation [700, 185].

N [80, 474, 429, 431, 430]. national [418, 462]. Nattino [62, 61, 63, 60]. natural [254, 657]. Nature [340, 376, 466, 297, 118]. nearest [444]. negative [363]. neighbor [444]. nested [176, 91, 52, 123, 317, 393]. Net [274, 95]. network [597, 556, 399, 286, 370, 506, 678, 698, 109, 108, 568, 210]. networks [200, 540, 126, 195]. Neuenschwander [67, 66, 68, 69]. Neural [540, 698]. neurocognitive [680]. neurodevelopment [98]. neuroimaging [408, 407, 409, 406, 125, 680, 642, 261, 404, 410, 238, 536]. neuron [644]. neurostimulator [217]. next [188]. NHANES [596]. Nicholas [257]. Nick [519]. Nicola [116]. Nicole [429, 431, 430]. Nigel [475, 472, 473]. Noah [160, 163, 164]. nocturnal [190]. node [333]. nodewise [576]. noise [450]. noisy [244, 594]. Non [623, 3, 375, 233, 236]. non-Gaussian [233, 236]. non-linear [3, 375]. Non-parametric [623]. noncompliance [459]. nonequidistant [168]. Nonhomogeneous [420, 207]. nonignorability [435]. nonignorable [649]. noninferiority [573]. Noniterative [485]. Nonlinear [307, 497, 443, 191]. nonmixture [345]. nonmonotone [443, 361]. Nonnegative [141, 269]. Nonparametric [207, 563, 414, 343, 565, 271, 272, 160, 627, 237, 651, 439, 159, 161, 128, 564, 80, 594, 645, 79, 81, 328, 273, 397, 172, 127, 487, 356, 486, 559, 595, 41]. nonparanormal [576]. nonproportional [211, 103]. nonresponse [321]. nonsystematic [189]. nontargeted [607]. normal [200, 671, 140, 282, 354]. normally [24]. note [571]. notes [254]. Novel [25, 481, 432, 427, 566, 429, 532, 431, 428, 430, 660, 199]. nuclear [417]. null [271, 26]. nulls [661]. number [250, 355, 678, 501]. numbers [83]. NY [39].

O [41]. obesity [329]. Objective [142]. observation [541]. observational [703, 173, 222, 319, 645, 661, 579, 296, 512, 334, 109, 359, 666]. observations [218, 207, 93, 393]. observed [177, 242, 323, 450]. Obtaining [313]. occupancy [685, 110]. occurrence [357]. odds [462, 561, 86]. O’Hagan [67, 66, 70, 68, 69]. Ohio [225]. Oijen [376]. Ombao [261]. Omic [341]. omics [609, 82]. oncology [327, 220, 149]. One [84, 551, 655]. One-step [84, 655]. online [22]. onset [105, 32, 623, 232]. open [659, 45]. open-population [45]. Operating [140, 287]. operator [581]. opioid [225]. Optimal [639, 110, 86, 222, 243, 438, 446, 700, 365, 122, 202, 330, 535, 313]. Optimality [211, 455]. Optimization [402, 217]. optimize [33]. optimized [96]. optimizing [217, 483]. Order [97]. Order-restricted [97]. ordinal [181, 291, 295, 294, 76, 292, 293]. origin [32, 539]. Ornstein [219]. osteoporosis [447]. other [466]. outbreak [242]. Outcome [515, 516, 320, 269, 91, 489, 15, 17, 602, 366, 448, 561, 419]. Outcome-adaptive [515, 516]. outcome-dependent [366]. outcomes [204, 291, 295, 84, 435, 513, 396, 593, 98, 294, 647, 33, 700, 76, 514, 275, 184, 172, 292, 634, 361, 610, 25, 545, 24, 288, 451, 660, 232, 293]. outlier [329, 455, 150]. overdispersion [94].

P [77, 337, 466, 300]. p. [376]. PA [149]. PA-CRM [149]. package [424]. Pagano [613]. Pair [656]. Pair-switching [656]. paired [127]. pairs [510]. pairwise [278, 581]. Pal [408, 407, 409, 406]. pan [34]. pan-cancer [34]. pandemics [466]. panel [212, 356]. Pani [466]. Pankaj [517]. Parameter [177, 416, 336, 277, 335]. parameters [56, 124]. Parametric [378, 223, 645, 623]. parent [32]. parent-of-origin [32]. Park [705]. Partial [125, 372, 576, 258, 268]. partially [450, 602, 496]. partitioning [532]. partly [316]. pathway [628, 238]. patient [401, 599, 109]. patients [418, 102, 687]. Pattern [701]. Pattern-based [701]. Paul [422, 296]. PCA [283]. pediatric [149, 680]. pedigrees [402]. peer [363]. Penalized [620, 50, 282, 192, 329, 3, 375, 285, 351, 364, 314, 660, 252]. Peng [379]. Pennell [62, 61, 63, 60]. people [397]. peptide [697]. peptides [697]. performance [209]. perinatal [510]. period [242, 638]. Permutation [490, 36]. perspective [330, 517]. Peter [160]. pharmacodynamics [327]. pharmacokinetics [327]. phase [598, 327, 273, 493, 149, 462, 25, 232, 449]. phenotypes [192, 71]. Pierpaolo [153]. plane [640]. Plant [40]. platform [600, 601]. plot [221]. plotly [226]. point [441, 334, 166]. points [534]. Poisson [214, 283, 360, 169, 569]. Poisson-gamma [360]. policies [162, 165, 163, 164]. Polya [659]. polygenic [349]. polyglutamine [416]. Pool [446]. pooled [603, 434]. pools [648]. population [181, 582, 31, 80, 646, 567, 45, 79, 81, 485, 367, 676, 215, 191, 504, 364, 554, 74]. population-based [485]. populations [659, 461, 549, 657, 458, 178, 112, 111]. positive [420, 353]. positive-definite [353]. possibly [443]. Post [250, 631, 502]. post-randomization [502]. Post-selection [250]. Post-selective [631]. Poststratification [241].

Posttreatment [566, 243]. potential [599, 448]. potentials [28, 534]. Power [104, 121, 334, 585, 291, 295, 9, 245, 675, 491, 294, 292, 103, 293]. powered [324]. powerful [179]. pp [340, 337, 38, 260, 380, 707, 259, 298, 338, 300, 299, 297, 296, 118, 339, 379, 6, 518, 258, 341, 257, 261]. Pradhan [518]. Precision [117, 291, 295, 155, 294, 292, 293, 74]. predicting [697]. Prediction [172, 212, 254, 548, 489, 480, 676, 34, 360, 698, 349, 550]. prediction-set [548]. predictions [549]. predictive [306, 401, 443, 182, 378]. Predictively [67, 66, 65, 70, 68, 69]. predictors [435, 396, 102, 587, 672]. PrediXcan [192]. preexperimental [537]. preference [459]. preference-based [459]. Presence [336, 646, 461, 85, 364, 335, 508, 174, 143]. presentation [697]. Press [517, 424, 259, 40, 612, 227, 465, 425, 261, 41, 613]. prevalence [21]. preventive [612, 686]. primary [249, 121, 571]. principal [55, 593, 638, 268, 581]. Principles [613]. Prior [66, 55, 67, 659, 142, 65, 70, 68, 69, 35, 188, 26]. Prioritizing [697]. priors [374, 368, 478, 234, 537]. Probabilistic [368]. probabilities [221, 26]. probability [372, 565, 653, 362, 255, 53, 312, 705]. problem [566, 134]. procedure [247, 491, 127, 179, 24]. procedures [211, 599]. process [392, 224, 486, 534]. processes [582, 207, 374, 177, 99, 166]. processing [254]. profile [75]. profiling [704]. progesterone [543]. prognostic [443]. programs [73]. progression [126]. projection [560]. prone [462, 696]. propensity [206, 88, 663, 362]. properties [311, 444]. proportional [244, 621, 622, 275, 509, 561, 22]. proportions [463]. proposals [612]. prospective [170]. prostate [532]. protein [34, 445]. proteins [416, 697]. pseudo [677, 278]. pseudo-covariates [677]. pseudo-likelihood [278]. psoriatic [32, 693]. Pte [340]. public [611]. publication [653, 499]. Publishing [707, 377, 370]. pure [123, 317, 449]. Pursuing [367]. pursuit [352].

Q [229]. quadratic [574]. Quantification [35, 625, 695]. Quantifying [419, 333, 370, 667]. Quantile [213, 649, 133, 683, 543, 488, 694, 642, 486]. quantitative [140]. quarantine [365].

R [80, 707, 424, 338, 227, 226, 154, 296, 341, 377, 691, 41, 259, 298, 40, 378, 39, 465, 258, 425, 423]. RA [374]. radiogenomics [631]. Radislav [300]. Rafael [259]. Random [333, 560, 190, 145, 11, 102, 705, 251, 208, 583, 284]. random-effects [145]. Randomization [27, 113, 706, 43, 502, 662, 599, 114, 464, 24, 660, 44]. randomized [597, 204, 703, 113, 291, 295, 245, 567, 461, 385, 383, 104, 384, 675, 294, 121, 436, 689, 691, 215, 292, 361, 382, 386, 561, 114, 201, 314, 122, 503, 585, 615, 293, 459]. rank [229, 343, 576, 140, 498, 103, 4]. rank-based [229, 576, 140, 4]. ranked [447]. ranking [322]. ranks [322]. rare [188]. rate [572, 548, 622, 571, 179]. rates [237]. ratio [703, 523, 522, 462, 421, 524, 624, 561, 282, 521, 525, 554, 459]. ratio-based [624]. ratios [86]. Raton [337, 517, 155, 117, 38, 260, 380, 422, 41, 259, 298, 40, 338, 612, 613, 300, 296, 339, 379, 6, 153, 258, 425, 341, 257, 261]. Re [449]. Re-calibrating [449]. reaction [515, 516]. read [169]. Reader [515, 516]. readmission [506]. reassessment [149]. recapture [181, 45, 251, 112, 395, 111]. Receiver [287]. record [648, 254]. record-based [648]. records [320, 198, 329, 489, 168, 696, 167, 122, 649]. Recruitment [360]. recurrence [243]. recurrent [23, 146, 622, 627, 135, 87, 50]. Reducing [456, 607]. reduction [55, 403, 657, 458, 362, 370, 616]. Referees [156, 301, 468]. reference [271]. referenced [21]. regime [438, 220]. regimen [77, 327, 2]. Regimes [117, 562, 446, 33, 89, 202, 666]. regions [374]. registration [235]. registries [653]. Regression [587, 363, 391, 204, 389, 562, 133, 30, 62, 329, 175, 683, 668, 447, 396, 456, 485, 625, 497, 309, 669, 272, 576, 378, 677, 101, 606, 493, 196, 632, 637, 502, 700, 61, 63, 400, 275, 56, 510, 59, 64, 205, 308, 346, 85, 60, 246, 440, 53, 543, 209, 356, 445, 314, 434, 580, 488, 496, 694, 72, 642, 486, 51, 213, 534, 649, 234, 667, 672, 106]. Regression-based [363, 562]. regressions [674]. regularization [658, 544]. Regularized [239, 576, 678, 672]. regulated [192]. Reich [6]. reinfection [148]. Rejoinder [295, 432, 81, 165, 267, 516, 64, 70, 112, 592, 476, 336, 386, 116, 114, 525, 161, 410, 531, 152]. Relapse [148]. related [28, 687, 453, 534, 695]. relatedness [185]. relationship [677]. relationships [99]. Relative [145, 587, 76, 415]. reliability [75, 183]. remedies [118]. removal [58]. Repeated [190, 603]. Replication [296]. Report [8, 158, 303, 470]. reported [606]. reporting [90, 688]. reproducibility [667]. reproducible [413, 154]. reproducing [637]. reproductive [83]. rerandomization [656]. Resampling [202]. Resampling-based [202]. Research [154, 519, 462, 518]. researchers [118]. residual [438]. resistant [109]. resource [366]. resource-limited [366]. respiratory [352, 699]. response [481, 668, 272, 682, 308, 599, 504, 642, 24, 569]. response-adaptive [481, 599, 504, 24]. responses [233, 632, 219, 553, 568]. resting [328, 284]. resting-state [328, 284]. Restricted [400, 318, 438, 91, 479, 129, 288, 97, 595, 666]. restriction [318]. results [413, 128]. Retrospective [170]. Review [340, 611, 337, 519, 517, 155, 117, 38, 260, 380, 422, 707, 706, 41, 424, 259, 376, 298, 466, 40, 338, 612, 613, 467, 227, 378, 39, 300, 299, 226, 705, 297, 154, 296, 118, 339, 379, 465, 6, 153, 518, 258, 425, 341, 257, 377, 261, 423, 193]. Richard [139, 38, 40, 138, 137]. Riemannian [280]. right [388, 617, 53]. right-censored [388, 617]. right-truncated [53]. Riley [139, 138, 137]. Risk [489, 212, 549, 420, 573, 144, 102, 123, 86, 96, 232, 349, 449]. risks [264, 513, 266, 263, 267, 699, 392, 620, 317, 265, 618, 143]. RMRF [190]. RNA [9, 556, 445, 371, 252]. RNA-seq [9]. Robert [337]. Robins [591, 590]. Robust [198, 433, 18, 17, 115, 304, 538, 1, 116, 546, 581, 77, 206, 180, 567, 648, 132, 256, 305, 662, 691, 555, 2, 464, 202, 348]. Robustness [113, 114]. ROC [331, 676, 131, 97]. ROC-guided [131]. roles [333]. Rosenbaum [296]. Rosenblum [294]. RStudio [299]. rule [330, 481]. rules [144, 483, 96, 122, 274]. Runze [260]. Rwanda [366]. Ryoung [705].

S [527, 530, 467, 529]. Salanti [139, 138, 137]. Sally [139, 138, 137]. Sample [482, 461, 66, 421, 103, 67, 43, 134, 495, 623, 65, 70, 68, 69, 366, 334, 684, 464, 35, 537]. sampled [654]. samples [62, 447, 61, 63, 59, 64, 127, 123, 60]. sampling [133, 402, 324, 58, 273, 221, 32, 462, 500, 596, 110, 366, 696, 25]. Sandeep [518]. Santos [466]. SBART [389]. Scalable [256, 200, 433]. scalar [497]. scale [9, 456, 19, 662, 390]. scan [586]. scanning [651]. Scharfstein [294]. schedule [33]. schemes [366]. Schmidli [67, 66, 68, 69]. Science [259, 338, 300, 260]. scientifically [288]. scope [189]. Score [583, 206, 88, 324, 92, 663, 362, 170, 509]. scores [55, 349]. Scott [163, 164]. Screening [557, 420, 333, 682, 182, 314, 390, 553]. Sean [527, 530, 529]. Sebastian [67, 66, 68, 69]. Seber [142, 100]. secondary [192, 368, 571, 71]. sectional [104, 623, 391]. sector [686]. SEER [352]. Segal [294]. Segmented [667]. selected [124]. selection [572, 228, 515, 320, 562, 5, 54, 192, 247, 497, 146, 250, 455, 413, 516, 101, 233, 635, 196, 682, 558, 34, 49, 538, 20, 182, 440, 555, 315, 559, 234, 536]. Selective [548, 370, 631]. self [606]. self-reported [606]. semi [198, 392, 620]. semi-competing [392, 620]. semi-supervised [198]. semicompeting [264, 266, 263, 267, 265]. Semiparametric [389, 269, 646, 345, 606, 321, 19, 135, 208, 539, 268, 209, 694, 393, 563, 485, 240, 622, 316, 647, 134, 394, 215, 171, 445, 643]. Senn [467]. Sensitivity [220, 512, 703, 435, 368, 566, 442, 361, 492]. sensor [197]. Seokho [705]. separable [484]. September [77, 375]. seq [9]. sequences [666]. sequencing [556, 107, 188, 554, 371, 199]. Sequential [54, 604, 491, 436, 691, 308, 191, 503]. sequentially [563, 12]. Sergio [258]. Series [153, 692, 237, 560]. set [447, 548, 610]. set-based [610]. sets [11, 535]. setting [211]. settings [366]. severities [699]. Shahjahan [340]. Shahn [591, 590]. Shannon [117]. shape [171]. Shapiro [80]. Shared [336, 19, 335]. sharing [56]. Sharp [76]. Shelter [39]. Shen [228]. shiny [226]. Shortreed [515]. shrinkage [495, 82]. Sievert [226]. sign [581]. signal [28, 247]. signals [533]. significance [28, 271]. Silge [423]. Silke [299]. similarities [495]. Simon [706, 429, 431, 430, 160, 163, 164, 253]. Simple [150]. Simplifying [608]. Simulation [398]. Simulation-based [398]. simulations [324]. Simultaneous [322, 329, 455, 355, 665, 440, 48, 315, 403, 580]. simultaneously [601]. Singapore [340, 297]. single [556, 347, 552, 609, 205, 107, 371]. single-cell [556, 552, 609, 107, 371]. single-index [205]. sites [185]. size [181, 582, 80, 482, 461, 79, 81, 647, 421, 130, 184, 334, 684, 35, 103, 537]. Sizes [66, 67, 104, 65, 70, 68, 69]. skew [630]. skew-elliptical [630]. skewed [218, 632, 445]. slab [101]. Small [366, 139, 136, 138, 452, 137, 527, 530, 529, 528]. small-area [452].

Small-sample [366]. SMART [684]. SMIM [492]. smoothed [509]. smoothing [329]. Smykla [611]. SNP [332]. social [597]. soft [389]. software [162, 165, 163, 164]. solution [566]. Solutions [634]. somatic [700, 185]. Song [705]. Sons [467, 378]. sources [367]. space [100, 480, 637, 191, 681]. space-for-time [100]. spaced [542, 393]. Sparse [358, 94, 280, 37, 344]. sparsely [654]. sparsity [355]. Spatial [204, 432, 693, 429, 431, 428, 430, 395, 354, 629, 374, 337, 329, 427, 45, 444, 352, 663, 309, 310, 532, 369, 581, 427, 40]. spatial-temporal [337]. spatially [21, 699]. spatiotemporal [533]. Spearman [343]. specific [682, 647]. specificity [608]. specified [452]. Spectra [498]. spectral [328, 633]. Speeding [324]. SpeLLL [498]. spending [106]. spike [101]. spikes [594]. spillover [204, 675]. spline [252]. splines [644]. Springer [340, 611, 707, 376, 466, 299, 705, 297, 118, 377]. Springer-Verlag [299]. squared [43, 444]. squares [77, 258, 2, 125]. stability [413]. stacked [281]. stage [690, 220, 675, 191, 253, 314, 535]. staggered [626]. Stallard [475, 474, 472, 473]. standard [608]. Stanley [62, 61, 63, 60]. Stata [258]. state [328, 480, 168, 191, 681, 284]. state-space [191]. stationary [534]. Statistical [320, 117, 254, 675, 95, 481, 517, 247, 300, 6, 584, 199, 260, 338, 467, 227, 118]. statistics [707, 586, 550, 449, 611, 257]. status [212]. stay [506]. steepest [418]. step [84, 186, 655, 49]. Stephen [611, 139, 706, 467, 138, 137]. stepped [482, 104, 18, 689, 130]. stepped-wedge [689]. Stijn [523, 522, 524]. stimulus [195]. stimulus-locked [195]. stochastic [450, 416, 220, 273, 178, 564]. strains [607]. strategies [189, 448, 684]. strategy [638]. Stratified [418, 100]. stratum [638]. streams [22]. Structural [615, 210, 192, 91, 126, 52, 393, 312, 258]. structure [98, 665, 307, 554]. Structured [10, 458, 495, 510, 358, 288, 234, 568]. structures [395, 664, 672]. Stuart [152]. studies [703, 320, 319, 175, 648, 645, 402, 598, 16, 661, 23, 566, 604, 15, 493, 579, 514, 95, 140, 644, 512, 361, 334, 82, 610, 684, 437, 253, 167, 341, 26, 459, 405, 550, 449, 71, 340, 296]. Studio [154]. Study [459, 139, 447, 162, 165, 385, 383, 136, 463, 384, 344, 138, 631, 163, 164, 382, 386, 125, 44, 137]. studying [25]. subclusters [482]. subcommunity [671]. subgraph [453]. subgraphs [453]. subgroup [658, 640, 655, 682]. subgroup-specific [682]. subgroups [641, 359]. Subhadip [408, 407, 409, 406]. Subhendu [466]. subject [54, 433, 93, 394, 213, 568]. subsample [247]. subsequent [220]. Subset [635]. subspace [456]. substance [361]. substantive [281]. subunit [647]. subunit-specific [647]. Sufficient [458, 147]. Sujata [466]. Sujit [6]. sum [324, 633]. summary [145, 464, 290, 550, 449]. summary-data [464]. Supervised [593, 198, 423]. support [668, 287, 580]. surgery [222]. surprising [209]. surrogacy [634, 546]. surrogate [203, 602, 279, 304, 547]. surveillance [270, 243, 329, 605]. surveys [325, 596]. survival [211, 147, 389, 563, 187, 244, 490, 388, 343, 352, 345, 489, 227, 344, 34, 616, 308, 628, 52, 85, 131, 87, 129, 439, 437, 50, 209, 387, 223, 22, 492, 51, 390, 615, 213, 595, 419, 318, 143]. survival-time-dependent [51]. susceptibility [693, 510]. switching [326, 615, 656]. Switzerland [376, 466, 118]. Syed [466]. symmetric [134]. symptomatic [687]. Synthesizing [364]. systems [605].

T [117, 257, 377, 331]. Taber [612]. Tailored [243]. Taimre [300]. Tang [229]. Tanur [611]. target [567, 676, 602]. Targeted [109, 84, 20, 641]. Tathagata [518]. Taylor [517]. Tchetgen [523, 522, 524]. techniques [160, 159, 161]. temporal [337, 99, 600]. temporally [100]. Tensor [403, 642, 353, 632]. terminal [146, 174]. terms [35]. test [481, 55, 331, 28, 62, 229, 324, 92, 16, 93, 61, 63, 59, 64, 373, 60, 47, 624, 583, 103, 4, 348, 332]. tested [128]. Testing [661, 399, 136, 32, 11, 138, 185, 547, 137, 211, 214, 608, 388, 271, 639, 101, 436, 127, 570, 22, 139]. tests [319, 608, 132, 657, 420, 98, 170, 512, 651, 464, 508, 103]. text [423]. Thall [118]. their [99]. theory [704, 705]. therapeutic [73]. therapy [634]. there [55]. third [467]. Thomas [300]. Thompson [706, 261, 221]. Thompson-like [221]. threshold [564]. threshold-based [564]. thresholding [248]. throughput [629, 667]. Tian [41]. Tidyverse [338]. time [582, 291, 295, 701, 277, 133, 84, 9, 438, 621, 242, 608, 385, 91, 383, 100, 418, 551, 542, 692, 463, 384, 593, 294, 436, 689, 479, 693, 237, 484, 215, 326, 308, 292, 504, 112, 52, 487, 509, 448, 539, 195, 129, 382, 386, 561, 498, 694, 72, 574, 51, 315, 111, 293, 595, 560, 318, 153]. time-course [9]. time-dependent [621, 509, 694]. time-frequency [498]. time-heterogeneous [542]. time-interaction [582]. time-lagged [561]. time-to-cure [277]. time-to-event [291, 295, 84, 463, 593, 294, 436, 215, 292, 72, 293]. time-varying [133, 91, 418, 52, 539, 195]. times [175, 541]. timing [222]. Ting [527, 530, 529]. tissue [107]. tools [305]. Torben [523, 522, 524]. toxicity [682, 215]. Tractable [630]. trait [140, 550]. traits [349]. trajectories [701, 80, 417, 79, 81, 411]. trajectory [256]. transcriptomics [629]. transformation [646, 396, 316, 140]. transformation-free [396]. transition [148]. Translocation [586]. transmissibility [687]. transmission [83]. transplant [418]. transport [700]. Transporting [567, 178, 676]. Treatment [117, 306, 77, 481, 562, 438, 446, 88, 198, 269, 180, 645, 23, 220, 144, 604, 15, 640, 17, 189, 682, 494, 362, 662, 76, 689, 479, 506, 89, 205, 483, 182, 448, 2, 109, 314, 641, 122, 202, 330, 359, 615, 312, 274, 666]. treatments [291, 295, 91, 220, 294, 292, 293]. tree [659, 671, 495, 510, 672, 666, 313]. tree-based [666]. tree-structured [495]. trees [389, 131, 633, 359]. trend [237, 373]. trial [597, 229, 227, 653, 579, 33, 443, 691, 215, 150, 561, 4, 232, 601]. trials [204, 481, 113, 291, 295, 245, 567, 461, 690, 220, 104, 460, 573, 273, 639, 227, 478, 294, 121, 149, 436, 505, 689, 638, 691, 360, 292, 626, 504, 361, 114, 201, 314, 600, 24, 122, 503, 585, 293, 74]. truncated [563, 617, 514, 53, 278]. truncation [346, 439]. Tsiatis [385, 383, 384, 117]. Tsung [264, 266, 265]. tuberculosis [109]. Tukey [373]. tumors [185]. twin [328]. Two [214, 305, 628, 314, 43, 598, 690, 220, 273, 675, 593, 493, 11, 134, 462, 684, 25, 253, 464, 449]. two-dimensional [593]. Two-group [214]. Two-level [628]. two-phase [598, 273, 493, 462, 25, 449]. two-sample [43, 134, 464]. Two-stage [314, 690, 220, 675, 253]. type [254, 460, 658, 551]. types [289].

Uhlenbeck [219]. UK [411]. Ultra [240, 555]. Ultra-high [555]. ultrahigh [12]. unbalanced [542]. unblinded [385, 383, 384, 382, 386]. uncertainty [625, 508]. underreported [270]. underrepresented [549]. undirected [650]. unequal [113, 130, 114]. unequally [542]. unified [572, 120, 492]. unimodal [271]. union [332]. unit [75]. unknown [307]. unmeasured [368]. updating [22]. Upper [181]. US$ [6]. US$113.63 [118]. USA [258]. USD [260, 261]. use [460, 361]. Using [147, 289, 481, 389, 55, 320, 291, 295, 701, 145, 706, 329, 368, 424, 254, 607, 447, 466, 40, 256, 551, 704, 227, 378, 653, 677, 333, 17, 493, 294, 14, 316, 321, 682, 579, 362, 148, 19, 462, 160, 34, 540, 532, 652, 691, 185, 631, 596, 292, 373, 442, 52, 688, 99, 258, 109, 425, 618, 159, 161, 122, 166, 492, 293, 672, 419, 537, 666, 313, 377]. utilities [682]. utility [547]. Utilizing [413].

vaccination [607, 686]. vaccine [372, 385, 120, 383, 273, 384, 382, 386, 253]. vaccines [697]. Vaisman [300]. validation [95, 634, 696, 584]. value [655, 575]. valued [15, 533]. values [51, 313]. Vansteelandt [523, 522, 524]. Variable [515, 562, 497, 146, 516, 523, 192, 247, 522, 104, 310, 101, 233, 494, 160, 34, 524, 538, 440, 555, 521, 525, 159, 161, 315, 459]. variables [54, 668, 502, 705, 577, 557, 452]. Variance [362, 255, 29, 660]. variances [433]. variant [188]. variants [706]. variate [286, 665]. variation [250, 577]. varies [504]. Varying [347, 434, 133, 91, 418, 102, 52, 539, 195]. Varying-coefficient [434]. vector [287]. vegetation [442]. Venturini [258]. Verlag [299]. version [388]. versus [145, 170]. via [77, 374, 586, 338, 285, 37, 566, 576, 653, 186, 286, 655, 558, 510, 351, 678, 487, 2, 581, 544, 533, 210]. view [457]. vine [542]. violators [446]. visualization [226]. Vivek [518]. Vladimir [80].

W [229, 114, 377]. Wallace [77]. Wang [228, 523, 522, 115, 524]. warping [701]. Weak [464, 247, 661]. Weak-instrument [464]. weak-signal-assisted [247]. wearable [166]. web [226]. web-based [226]. Weber [67, 66, 68, 69]. wedge [482, 104, 18, 689, 130]. weigh [701]. weigh-in [701]. Weight [123, 317]. Weighted [175, 77, 372, 565, 362, 255, 2, 660, 103]. weighting [653, 53]. weights [312]. Welham [424]. well [673]. Wellington [466]. Were [411]. Wesley [261]. which [491]. while [627]. while-alive [627]. whole [554]. whose [307]. wide [140, 550]. wildlife [659]. Wiley [467, 378]. William [611]. Williamson [160]. win [275, 421]. win-fractions [275]. window [129]. windows [510]. wise [414]. WiSER [433]. within [597, 433]. within-subject [433]. without [249]. Wood [429, 431, 428, 430]. work [673]. Writing [612]. Wu [41, 227, 408, 407, 409, 406].

Xin [41]. xiv [376]. Xue [228].

Yates [377]. Ye [527, 530, 529, 528]. Yen [264, 266, 265]. Yen-Tsung [264, 266, 265]. Yi [422]. yield [74]. yields [413]. Ying [408, 407, 409, 406]. Yingwei [379]. Yong [139, 138, 137]. Yoon [705]. York [611, 519, 706, 424, 465]. Yu [379].

Zach [591, 590]. Zdravko [300]. Zero [169, 569, 269, 142, 463, 172, 371]. zeroaugmented [142]. Zero-inflated [169, 569, 463, 172, 371]. Zhang [228, 260]. Zoe [424]. Zöllner [299]. Zou [260].

References

Wang:2005:REF

[1] You-Gan Wang, Xu Lin, and Min Zhu. Robust estimating functions and bias correction for longitudinal data analysis. Biometrics, 61(3):684–691, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See comments [115] and rejoinder [116].

Wallace:2015:DRD

[2] Michael P. Wallace and Erica E. M. Moodie. Doubly-robust dynamic treatment regimen estimation via weighted least squares. Biometrics, 71 (3):636–644, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correlation [77].

Gasparrini:2017:PFD

[3] Antonio Gasparrini, Fabian Scheipl, Ben Armstrong, and Michael G. Kenward. A penalized framework for distributed lag non-linear models. Biometrics, 73(3):938–948, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [375].

Zhang:2019:CAR

[4] Wei Zhang, Aiyi Liu, Larry L. Tang, and Qizhai Li. A cluster-adjusted rank-based test for a clinical trial concerning multiple endpoints with application to dietary intervention assessment. Biometrics, 75(3):821–830, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [229].

Chen:2019:ELB

[5] Chixiang Chen, Biyi Shen, Lijun Zhang, Yuan Xue, and Ming Wang. Empirical-likelihood-based criteria for model selection on marginal analysis of longitudinal data with dropout missingness. Biometrics, 75(3):

950–965, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [228].

Tang:2020:BRB

[6] Becky Tang and Amy H. Herring. Book review: Bayesian statistical methods, Reich, Brian J. Ghosh, Sujit K., Boca Raton, FL: Chapman and Hall/ CRC, 2019, Hard cover. pp. 288. US$ 79.96. Biometrics, 71(3):671–672, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2020:IIa

[7] Anonymous. Issue information. Biometrics, 76(1):1–4, March 2020. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2020:RE

[8] Anonymous. Report of the Editors — 2019. Biometrics, 76(1):5–8, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cao:2020:LSM

[9] Meng Cao, Wen Zhou, F. Jay Breidt, and Graham Peers. Large scale maximum average power multiple inference on time-course count data with application to RNA-seq analysis. Biometrics, 76(1):9–22, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2020:SGE

[10] Mengyun Wu, Qingzhao Zhang, and Shuangge Ma. Structured geneenvironment interaction analysis. Biometrics, 76(1):23–35, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lavancier:2020:TIB

[11] Frédéric Lavancier, Thierry Pécot, Liu Zengzhen, and Charles Kervrann. Testing independence between two random sets for the analysis of colocalization in bioimaging. Biometrics, 76(1):36–46, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2020:BGL

[12] Qi Zheng, Hyokyoung G. Hong, and Yi Li. Building generalized linear models with ultrahigh dimensional features: A sequentially conditional approach. Biometrics, 76(1):47–60, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2020:IFB

[13] Jun Young Park and Eric F. Lock. Integrative factorization of bidimensionally linked matrices. Biometrics, 76(1):61–74, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2020:ECM

[14] Keunbaik Lee, Hyunsoon Cho, Min-Sun Kwak, and Eun Jin Jang. Estimation of covariance matrix of multivariate longitudinal data using modified choleksky and hypersphere decompositions. Biometrics, 76(1):75–86, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jiang:2020:BAJ

[15] Bei Jiang, Eva Petkova, Thaddeus Tarpey, and R. Todd Ogden. A Bayesian approach to joint modeling of matrix-valued imaging data and treatment outcome with applications to depression studies. Biometrics, 76(1):87–97, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Duan:2020:GIL

[16] Rui Duan, Ming Cao, Yang Ning, Mingfu Zhu, Bin Zhang, Aidan McDermott, Haitao Chu, Xiaohua Zhou, Jason H. Moore, Joseph G. Ibrahim, Daniel O. Scharfstein, and Yong Chen. Global identifiability of latent class models with applications to diagnostic test accuracy studies: A Gröbner basis approach. Biometrics, 76(1):98–108, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ju:2020:RIA

[17] Cheng Ju, David Benkeser, and Mark J. van der Laan. Robust inference on the average treatment effect using the outcome highly adaptive lasso. Biometrics, 76(1):109–118, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hughes:2020:RIS

[18] James P. Hughes, Patrick J. Heagerty, Fan Xia, and Yuqi Ren. Robust inference for the stepped wedge design. Biometrics, 76(1):119–130, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Linero:2020:SMS

[19] Antonio R. Linero, Debajyoti Sinha, and Stuart R. Lipsitz. Semiparametric mixed-scale models using shared Bayesian forests. Biometrics, 76 (1):131–144, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schnitzer:2020:DAL

[20] Mireille E. Schnitzer, Joel Sango, Steve Ferreira Guerra, and Mark J. van der Laan. Data-adaptive longitudinal model selection in causal inference with collaborative targeted minimum loss-based estimation. Biometrics, 76(1):145–157, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Amoah:2020:GFC

[21] Benjamin Amoah, Peter J. Diggle, and Emanuele Giorgi. A geostatistical framework for combining spatially referenced disease prevalence data from multiple diagnostics. Biometrics, 76(1):158–170, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xue:2020:OUA

[22] Yishu Xue, HaiYing Wang, Jun Yan, and Elizabeth D. Schifano. An online updating approach for testing the proportional hazards assumption with streams of survival data. Biometrics, 76(1):171–182, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gao:2020:ATA

[23] Jingya Gao, Pei-Fang Su, Feifang Hu, and Siu Hung Cheung. Adaptive treatment allocation for comparative clinical studies with recurrent events data. Biometrics, 76(1):183–196, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williamson:2020:RAR

[24] S. Faye Williamson and Sofía S. Villar. A response-adaptive randomization procedure for multi-armed clinical trials with normally distributed outcomes. Biometrics, 76(1):197–209, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2020:NTP

[25] Le Wang, Matthew L. Williams, Yong Chen, and Jinbo Chen. Novel two-phase sampling designs for studying binary outcomes. Biometrics, 76 (1):210–223, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zeng:2020:DAB

[26] Donglin Zeng, Zhiying Pan, and D. Y. Lin. Design and analysis of bridging studies with prior probabilities on the null and alternative hypotheses.

Biometrics, 76(1):224–234, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Loh:2020:RIG

[27] Wen Wei Loh, Michael G. Hudgens, John D. Clemens, Mohammad Ali, and Michael E. Emch. Randomization inference with general interference and censoring. Biometrics, 76(1):235–245, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Causeur:2020:FGT

[28] David Causeur, Ching-Fan Sheu, Emeline Perthame, and Flavia Rufini. A functional generalized F -test for signal detection with applications to event-related potentials significance analysis. Biometrics, 76(1):246–256, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shinohara:2020:DBA

[29] Russell T. Shinohara, Haochang Shou, Marco Carone, Robert Schultz, Birkan Tunc, Drew Parker, Melissa Lynne Martin, and Ragini Verma. Distance-based analysis of variance for brain connectivity. Biometrics, 76 (1):257–269, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Che:2020:IEE

[30] Menglu Che, Peisong Han, and Jerald F. Lawless. Improving estimation efficiency for regression with MNAR covariates. Biometrics, 76(1):270– 280, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Besbeas:2020:GFM

[31] Panagiotis Besbeas and Byron J. T. Morgan. A general framework for modeling population abundance data. Biometrics, 76(1):281–292, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lakhal-Chaieb:2020:THP

[32] Lajmi Lakhal-Chaieb, Richard J. Cook, and Yujie Zhong. Testing the heritability and parent-of-origin hypotheses for ages at onset of psoriatic arthritis under biased sampling. Biometrics, 76(1):293–303, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2020:ATD

[33] Ruitao Lin, Peter F. Thall, and Ying Yuan. An adaptive trial design to optimize dose-schedule regimes with delayed outcomes. Biometrics, 76 (1):304–315, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maity:2020:BDI

[34] Arnab Kumar Maity, Anirban Bhattacharya, Bani K. Mallick, and Veerabhadran Baladandayuthapani. Bayesian data integration and variable selection for pan-cancer survival prediction using protein expression data. Biometrics, 76(1):316–325, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wiesenfarth:2020:QPI

[35] Manuel Wiesenfarth and Silvia Calderazzo. Quantification of prior impact in terms of effective current sample size. Biometrics, 76(1):326–336, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Noma:2020:PIM

[36] Hisashi Noma, Kengo Nagashima, and Toshi A. Furukawa. Permutation inference methods for multivariate meta-analysis. Biometrics, 76(1):337– 347, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Helgeson:2020:BSC

[37] Erika S. Helgeson, Qian Liu, Guanhua Chen, Michael R. Kosorok, and Eric Bair. Biclustering via sparse clustering. Biometrics, 76(1):348–358, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2020:BRM

[38] Li-Pang Chen. Book review: Multistate models for the analysis of life history data, Cook, Richard J., Lawless, Jerald F., Boca Raton, FL: Chapman and Hall/CRC, 2020. Hard cover. pp. 441. CDN$ 121.00. Biometrics, 76(1):359–360, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2020:BRD

[39] Sehee Kim. Book review: Deep learning with R, Chollet, François, Allaire, Joseph J., Shelter Island, NY: Manning. Biometrics, 76(1):361–362, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fortin:2020:BRS

[40] Marie-Josée Fortin. Book review: Spatial data analysis in ecology and agriculture using R, Plant, Richard E., Boca Raton, FL: CRC Press, 2019.

Biometrics, 76(1):362, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chowdhury:2020:BRN

[41] Mohammed Chowdhury. Book review: Nonparametric models for longitudinal data: With implementation in R, Wu, Colin O., Tian, Xin, Boca Raton, FL: CRC Press. Biometrics, 76(1):363–364, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2020:IIb

[42] Anonymous. Issue information. Biometrics, 76(2):365–368, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Deng:2020:ABM

[43] Lu Deng, Han Zhang, Lei Song, and Kai Yu. Approximation of bias and mean-squared error in two-sample Mendelian randomization analyses. Biometrics, 76(2):369–379, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2020:MRA

[44] Han Zhang, Jing Qin, Sonja I. Berndt, Demetrius Albanes, Lu Deng, Mitchell H. Gail, and Kai Yu. On Mendelian randomization analysis of case-control study. Biometrics, 76(2):380–391, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Efford:2020:SOP

[45] Murray G. Efford and Matthew R. Schofield. A spatial open-population capture-recapture model. Biometrics, 76(2):392–402, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2020:HMC

[46] Pei-Sheng Lin and Jun Zhu. A heterogeneity measure for cluster identification with application to disease mapping. Biometrics, 76(2):403–413, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2020:AIT

[47] Yaru Song, Hongyu Zhao, and Tao Wang. An adaptive independence test for microbiome community data. Biometrics, 76(2):414–426, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2020:SCC

[48] Yueying Wang, Guannan Wang, Li Wang, and R. Todd Ogden. Simultaneous confidence corridors for mean functions in functional data analysis of imaging data. Biometrics, 76(2):427–437, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Michelot:2020:IMS

[49] Théo Michelot, Paul G. Blackwell, Simon Chamaillé-Jammes, and Jason Matthiopoulos. Inference in MCMC step selection models. Biometrics, 76(2):438–447, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2020:PSM

[50] Lili Wang, Kevin He, and Douglas E. Schaubel. Penalized survival models for the analysis of alternating recurrent event data. Biometrics, 76(2):448– 459, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yi:2020:CRS

[51] Yanyao Yi, Ting Ye, Menggang Yu, and Jun Shao. Cox regression with survival-time-dependent missing covariate values. Biometrics, 76(2):460– 471, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Seaman:2020:ATV

[52] Shaun Seaman, Oliver Dukes, Ruth Keogh, and Stijn Vansteelandt. Adjusting for time-varying confounders in survival analysis using structural nested cumulative survival time models. Biometrics, 76(2):472–483, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vakulenko-Lagun:2020:IPW

[53] Bella Vakulenko-Lagun, Micha Mandel, and Rebecca A. Betensky. Inverse probability weighting methods for Cox regression with right-truncated data. Biometrics, 76(2):484–495, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2020:SAV

[54] Zimu Chen, Zhanfeng Wang, and Yuan chin Ivan Chang. Sequential adaptive variables and subject selection for GEE methods. Biometrics, 76(2): 496–507, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bedrick:2020:DRP

[55] Edward J. Bedrick. Data reduction prior to inference: Are there consequences of comparing groups using a t-test based on principal component scores? Biometrics, 76(2):508–517, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mollenhoff:2020:ERC

[56] Kathrin Möllenhoff, Frank Bretz, and Holger Dette. Equivalence of regression curves sharing common parameters. Biometrics, 76(2):518–529, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Walker:2020:BCB

[57] Nelson B. Walker, Trevor J. Hefley, and Daniel P. Walsh. Bias correction of bounded location error in binary data. Biometrics, 76(2):530–539, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Haines:2020:MMM

[58] Linda M. Haines. Multinomial N-mixture models for removal sampling. Biometrics, 76(2):540–548, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nattino:2020:AGF

[59] Giovanni Nattino, Michael L. Pennell, and Stanley Lemeshow. Assessing the goodness of fit of logistic regression models in large samples: A modification of the Hosmer–Lemeshow test. Biometrics, 76(2):549–560, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [60, 61, 62, 63] and rejoinder [64].

Shmueli:2020:DAG

[60] Galit Shmueli. Discussion on “Assessing the goodness of fit of logistic regression models in large samples: A modification of the Hosmer– Lemeshow test” by Giovanni Nattino, Michael L. Pennell, and Stanley Lemeshow. Biometrics, 76(2):561–563, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [59, 64].

Liu:2020:DAGa

[61] Ivy Liu and Daniel Fernández. Discussion on “Assessing the goodness of fit of logistic regression models in large samples: A modification of the Hosmer–Lemeshow test” by Giovanni Nattino, Michael L. Pennell, and Stanley Lemeshow. Biometrics, 76(2):564–568, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [59, 64].

Chen:2020:DAG

[62] Li-Ching Chen and Jiun-Yi Wang. Discussion of “Assessing the goodnessof-fit of logistic regression models in large samples: A modification of the Hosmer–Lemeshow test,” by Giovanni Nattino, Michael L. Pennell, and Stanley Lemeshow. Biometrics, 76(2):569–571, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [59].

Liu:2020:DAGb

[63] Dandan Liu and Bryan E. Shepherd. Discussion on “Assessing the goodness of fit of logistic regression models in large samples: A modification of the Hosmer–Lemeshow test” by Giovanni Nattino, Michael L. Pennell, and Stanley Lemeshow. Biometrics, 76(2):572–574, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [59, 64].

Nattino:2020:RAG

[64] Giovanni Nattino, Michael L. Pennell, and Stanley Lemeshow. Rejoinder to “Assessing the goodness of fit of logistic regression models in large samples: A modification of the Hosmer–Lemeshow test”. Biometrics, 76 (2):575–577, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [59, 60, 61, 62, 63].

Neuenschwander:2020:PCP

[65] Beat Neuenschwander, Sebastian Weber, Heinz Schmidli, and Anthony O’Hagan. Predictively consistent prior effective sample sizes. Biometrics, 76(2):578–587, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [66, 67, 68, 69] and rejoinder [70].

Kaizer:2020:DPC

[66] Alexander Kaizer and John Kittelson. Discussion on “Predictively Consistent Prior Effective Sample Sizes” by Beat Neuenschwander, Sebastian Weber, Heinz Schmidli, and Anthony O’Hagan. Biometrics, 76(2):588– 590, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [65, 70].

Biswas:2020:DPC

[67] Atanu Biswas and Jean-François Angers. Discussion on “Predictively consistent prior effective sample sizes,” by Beat Neuenschwander, Sebastian Weber, Heinz Schmidli, and Anthony O’Hagan. Biometrics, 76(2):591– 594, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [65, 70].

Park:2020:DPC

[68] Yeonhee Park and Ruitao Lin. Discussion on “Predictively consistent prior effective sample sizes,” by Beat Neuenschwander, Sebastian Weber, Heinz Schmidli, and Anthony O’Hagan. Biometrics, 76(2):595–598, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [65, 70].

Rosner:2020:DPC

[69] Gary L. Rosner and Peter Müller. Discussion on “Predictively consistent prior effective sample sizes,” by Beat Neuenschwander, Sebastian Weber, Heinz Schmidli, and Anthony O’Hagan. Biometrics, 76(2):599–601, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [65, 70].

Neuenschwander:2020:RPC

[70] Beat Neuenschwander, Sebastian Weber, Heinz Schmidli, and Anthony O’Hagan. Rejoinder to “Predictively consistent prior effective sample sizes,” by Beat Neuenschwander, Sebastian Weber, Heinz Schmidli, and Anthony O’Hagan. Biometrics, 76(2):602–605, June 2020. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [65, 66, 67, 68, 69].

Zhou:2020:ASP

[71] Fan Zhou, Haibo Zhou, Tengfei Li, and Hongtu Zhu. Analysis of secondary phenotypes in multigroup association studies. Biometrics, 76(2): 606–618, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wei:2020:GBA

[72] Yue Wei, Yi Liu, Tao Sun, Wei Chen, and Ying Ding. Gene-based association analysis for bivariate time-to-event data through functional regression with copula models. Biometrics, 76(2):619–629, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Psioda:2020:BDB

[73] Matthew A. Psioda, Kuolung Hu, Yang Zhang, Jean Pan, and Joseph G. Ibrahim. Bayesian design of biosimilars clinical programs involving multiple therapeutic indications. Biometrics, 76(2):630–642, June 2020. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2020:DPM

[74] Ying-Qi Zhao and Michael L. LeBlanc. Designing precision medicine trials to yield a greater population impact. Biometrics, 76(2):643–653, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2020:PIU

[75] Kevin He, Claudia Dahlerus, Lu Xia, Yanming Li, and John D. Kalbfleisch. The profile inter-unit reliability. Biometrics, 76(2):654–663, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2020:SBR

[76] Jiannan Lu, Yunshu Zhang, and Peng Ding. Sharp bounds on the relative treatment effect for ordinal outcomes. Biometrics, 76(2):664–669, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2020:CDR

[77] Anonymous. Correction to “Doubly-robust dynamic treatment regimen estimation via weighted least squares,” by Michael P. Wallace and Erica E. M. Moodie; 71 (3), 636–644, September 2015. Biometrics, 76(2):670, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [2].

Anonymous:2020:IIc

[78] Anonymous. Issue information. Biometrics, 76(3):673–676, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Faulkner:2020:HBB

[79] James R. Faulkner, Andrew F. Magee, Beth Shapiro, and Vladimir N. Minin. Horseshoe-based Bayesian nonparametric estimation of effective population size trajectories. Biometrics, 76(3):677–690, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [80] and rejoinder [81].

Cappello:2020:DHB

[80] Lorenzo Cappello, Swarnadip Ghosh, and Julia A. Palacios. Discussion on “Horseshoe-based Bayesian nonparametric estimation of effective population size trajectories” by James R. Faulkner, Andrew F. Magee, Beth Shapiro, and Vladimir N. Minin. Biometrics, 76(3):691–694, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [79].

Faulkner:2020:RDH

[81] James R. Faulkner, Andrew F. Magee, Beth Shapiro, and Vladimir N. Minin. Rejoinder for discussion on “Horseshoe-based Bayesian nonparametric estimation of effective population size trajectories”. Biometrics, 76 (3):695–699, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [79].

Song:2020:BSE

[82] Yanyi Song, Xiang Zhou, Min Zhang, Wei Zhao, Yongmei Liu, Sharon L. R. Kardia, Ana V. Diez Roux, Belinda L. Needham, Jennifer A. Smith, and Bhramar Mukherjee. Bayesian shrinkage estimation of high dimensional causal mediation effects in omics studies. Biometrics, 76(3):700– 710, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ozanne:2020:BCM

[83] Marie V. Ozanne, Grant D. Brown, Angela J. Toepp, Breanna M. Scorza, Jacob J. Oleson, Mary E. Wilson, and Christine A. Petersen. Bayesian compartmental models and associated reproductive numbers for an infection with multiple transmission modes. Biometrics, 76(3):711–721, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2020:OST

[84] Weixin Cai and Mark J. van der Laan. One-step targeted maximum likelihood estimation for time-to-event outcomes. Biometrics, 76(3):722–733, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sheng:2020:CLR

[85] Ying Sheng, Yifei Sun, Detian Deng, and Chiung-Yu Huang. Censored linear regression in the presence or absence of auxiliary survival information. Biometrics, 76(3):734–745, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

VanderWeele:2020:OAC

[86] Tyler J. VanderWeele. Optimal approximate conversions of odds ratios and hazard ratios to risk ratios. Biometrics, 76(3):746–752, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tawiah:2020:BJF

[87] Richard Tawiah, Geoffrey J. McLachlan, and Shu Kay Ng. A bivariate joint frailty model with mixture framework for survival analysis of recurrent events with dependent censoring and cure fraction. Biometrics, 76(3): 753–766, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2020:EAT

[88] David Cheng, Abhishek Chakrabortty, Ashwin N. Ananthakrishnan, and Tianxi Cai. Estimating average treatment effects with a double-index propensity score. Biometrics, 76(3):767–777, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nguyen:2020:EIT

[89] Crystal T. Nguyen, Daniel J. Luckett, Anna R. Kahkoska, Grace E. Shearrer, Donna Spruijt-Metz, Jaimie N. Davis, and Michael R. Kosorok. Estimating individualized treatment regimes from crossover designs. Biometrics, 76(3):778–788, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stoner:2020:MHF

[90] Oliver Stoner and Theo Economou. Multivariate hierarchical frameworks for modeling delayed reporting in count data. Biometrics, 76(3):789–798, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hagiwara:2020:ESN

[91] Yasuhiro Hagiwara, Tomohiro Shinozaki, and Yutaka Matsuyama. Gestimation of structural nested restricted mean time lost models to estimate effects of time-varying treatments on a failure time outcome. Biometrics, 76(3):799–810, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Duan:2020:FST

[92] Rui Duan, Yang Ning, Shuang Wang, Bruce G. Lindsay, Raymond J. Carroll, and Yong Chen. A fast score test for generalized mixture models. Biometrics, 76(3):811–820, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2020:THD

[93] DongHyuk Lee, Soumendra N. Lahiri, and Samiran Sinha. A test of homogeneity of distributions when observations are subject to measurement errors. Biometrics, 76(3):821–833, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Afroz:2020:EOS

[94] Farzana Afroz, Matt Parry, and David Fletcher. Estimating overdispersion in sparse multinomial data. Biometrics, 76(3):834–842, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Marsh:2020:SIN

[95] Tracey L. Marsh, Holly Janes, and Margaret S. Pepe. Statistical inference for net benefit measures in biomarker validation studies. Biometrics, 76 (3):843–852, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2020:LBB

[96] Yanqing Wang, Ying-Qi Zhao, and Yingye Zheng. Learning-based biomarker-assisted rules for optimized clinical benefit under a risk constraint. Biometrics, 76(3):853–862, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2020:ORI

[97] Wei Zhang, Larry L. Tang, Qizhai Li, Aiyi Liu, and Mei-Ling Ting Lee. Order-restricted inference for clustered ROC data with application to fingerprint matching accuracy. Biometrics, 76(3):863–873, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

LaLonde:2020:DSM

[98] Amy LaLonde, Tanzy Love, Sally W. Thurston, and Philip W. Davidson. Discovering structure in multiple outcomes models for tests of childhood neurodevelopment. Biometrics, 76(3):874–885, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tadde:2020:DMM

[99] Bachirou O. Taddé, Hélène Jacqmin-Gadda, Jean-François Dartigues, Daniel Commenges, and Cécile Proust-Lima. Dynamic modeling of multivariate dimensions and their temporal relationships using latent processes: Application to Alzheimer’s disease. Biometrics, 76(3):886–899, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hance:2020:TSE

[100] Dalton J. Hance, Russell W. Perry, John M. Plumb, and Adam C. Pope. A temporally stratified extension of space-for-time Cormack–Jolly–Seber for migratory animals. Biometrics, 76(3):900–912, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joyner:2020:MEM

[101] Chase N. Joyner, Christopher S. McMahan, Joshua M. Tebbs, and Christopher R. Bilder. From mixed effects modeling to spike and slab variable selection: a Bayesian regression model for group testing data. Biometrics, 76(3):913–923, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2020:MME

[102] Yihao Li, Danh V. Nguyen, Esra Kürüm, Connie M. Rhee, Yanjun Chen, Kamyar Kalantar-Zadeh, and Damla Sentürk. A multilevel mixed effects varying coefficient model with multilevel predictors and random effects for modeling hospitalization risk in patients on dialysis. Biometrics, 76(3): 924–938, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yung:2020:SSP

[103] Godwin Yung and Yi Liu. Sample size and power for the weighted log-rank test and Kaplan–Meier based tests with allowance for nonproportional hazards. Biometrics, 76(3):939–950, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Harrison:2020:PCC

[104] Linda J. Harrison, Tom Chen, and Rui Wang. Power calculation for crosssectional stepped wedge cluster randomized trials with variable cluster sizes. Biometrics, 76(3):951–962, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coull:2020:CME

[105] Brent A. Coull, Seokho Lee, Glen McGee, Justin Manjourides, Murray A. Mittleman, and Gregory A. Wellenius. Corrections for measurement error due to delayed onset of illness for case-crossover designs. Biometrics, 76 (3):963–972, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zink:2020:FRH

[106] Anna Zink and Sherri Rose. Fair regression for health care spending. Biometrics, 76(3):973–982, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2020:JAS

[107] Wei Sun, Chong Jin, Jonathan A. Gelfond, Ming-Hui Chen, and Joseph G. Ibrahim. Joint analysis of single-cell and bulk tissue sequencing data to infer intratumor heterogeneity. Biometrics, 76(3):983–994, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xie:2020:IDA

[108] Shanghong Xie, Xiang Li, Peter McColgan, Rachael I. Scahill, Donglin Zeng, and Yuanjia Wang. Identifying disease-associated biomarker network features through conditional graphical model. Biometrics, 76(3): 995–1006, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2020:ETI

[109] Guanbo Wang, Mireille E. Schnitzer, Dick Menzies, Piret Viiklepp, Timothy H. Holtz, and Andrea Benedetti. Estimating treatment importance in multidrug-resistant tuberculosis using targeted learning: an observational individual patient data network meta-analysis. Biometrics, 76(3): 1007–1016, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reich:2020:OSD

[110] Henry T. Reich. Optimal sampling design and the accuracy of occupancy models. Biometrics, 76(3):1017–1027, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2020:CTC

[111] Wei Zhang and Simon J. Bonner. On continuous-time capture-recapture in closed populations. Biometrics, 76(3):1028–1033, September 2020. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [112].

Schofield:2020:RCT

[112] Matthew R. Schofield and Richard J. Barker. Rejoinder to “On continuous-time capture-recapture in closed populations”. Biometrics, 76(3):1034–1035, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [111].

Bartlett:2020:RAR

[113] Jonathan W. Bartlett. Robustness of ANCOVA in randomized trials with unequal randomization. Biometrics, 76(3):1036–1038, September 2020.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [114].

Wang:2020:RRA

[114] Bingkai Wang, Elizabeth L. Ogburn, and Michael Rosenblum. Rejoinder to “Robustness of ANCOVA in randomized trials with unequal randomization” by Jonathan W. Bartlett. Biometrics, 76(3):1039, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [113].

Lunardon:2020:CWA

[115] Nicola Lunardon and Giovanna Menardi. Comment on “Wang et al. (2005), Robust estimating functions and bias correction for longitudinal data analysis”. Biometrics, 76(3):1040–1042, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1].

Wang:2020:RCW

[116] You-Gan Wang, Xu Lin, and Min Zhu. Rejoinder to “Comment on ‘Wang et al. (2005), Robust estimating functions and bias correction for longitudinal data analysis’ by Nicola Lunardon and Giovanna Menardi”. Biometrics, 76(3):1043–1044, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1].

Chen:2020:BRD

[117] Li-Pang Chen. Book review: Dynamic Treatment Regimes: Statistical Methods for Precision Medicine Tsiatis, Anastasios A. Davidian, Marie Holloway, Shannon T. Laber, Eric B. Boca Raton, FL: Chapman and Hall/CRC, 2019. Biometrics, 76(3):1045–1046, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muller:2020:BRS

[118] Peter Müller. Book review: Statistical remedies for medical researchers. Thall, Peter Cham, Switzerland: Springer Nature Switzerland AG, 2020, Hard cover. pp. 291. US$113.63. Biometrics, 76(3):1047–1048, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2020:IId

[119] Anonymous. Issue information. Biometrics, 76(4):1049–1052, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gabriel:2020:UED

[120] Erin E. Gabriel, Michael C. Sachs, Dean A. Follmann, and Therese M-L. Andersson. A unified evaluation of differential vaccine efficacy. Biometrics, 76(4):1053–1063, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2020:PAC

[121] Dateng Li, Jing Cao, and Song Zhang. Power analysis for cluster randomized trials with multiple binary co-primary endpoints. Biometrics, 76(4): 1064–1074, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2020:UEH

[122] Peng Wu, Donglin Zeng, Haoda Fu, and Yuanjia Wang. On using electronic health records to improve optimal treatment rules in randomized trials. Biometrics, 76(4):1075–1086, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shin:2020:WCI

[123] Yei Eun Shin, Ruth M. Pfeiffer, Barry I. Graubard, and Mitchell H. Gail. Weight calibration to improve the efficiency of pure risk estimates from case-control samples nested in a cohort. Biometrics, 76(4):1087–1097, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2020:CCI

[124] Haibing Zhao and Xinping Cui. Constructing confidence intervals for selected parameters. Biometrics, 76(4):1098–1108, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2020:PLS

[125] Yue Wang, Joseph G. Ibrahim, and Hongtu Zhu. Partial least squares for functional joint models with applications to the Alzheimer’s disease neuroimaging initiative study. Biometrics, 76(4):1109–1119, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peterson:2020:BMM

[126] Christine B. Peterson, Nathan Osborne, Francesco C. Stingo, Pierrick Bourgeat, James D. Doecke, and Marina Vannucci. Bayesian modeling of multiple structural connectivity networks during the progression of Alzheimer’s disease. Biometrics, 76(4):1120–1132, December 2020. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pereira:2020:BNT

[127] Luz Adriana Pereira, Daniel Taylor-Rodríguez, and Luis Gutiérrez. A Bayesian nonparametric testing procedure for paired samples. Biometrics, 76(4):1133–1146, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2020:NED

[128] Wei Zhang, Aiyi Liu, Qizhai Li, and Paul S. Albert. Nonparametric estimation of distributions and diagnostic accuracy based on group-tested results with differential misclassification. Biometrics, 76(4):1147–1156, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tian:2020:ECT

[129] Lu Tian, Hua Jin, Hajime Uno, Ying Lu, Bo Huang, Keaven M. Anderson, and Lj Wei. On the empirical choice of the time window for restricted mean survival time. Biometrics, 76(4):1157–1166, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matthews:2020:HES

[130] John N. S. Matthews. Highly efficient stepped wedge designs for clusters of unequal size. Biometrics, 76(4):1167–1176, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2020:RGS

[131] Yifei Sun, Sy Han Chiou, and Mei-Cheng Wang. ROC-guided survival trees and ensembles. Biometrics, 76(4):1177–1189, December 2020. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dukes:2020:DRT

[132] Oliver Dukes, Vahe Avagyan, and Stijn Vansteelandt. Doubly robust tests of exposure effects under high-dimensional confounding. Biometrics, 76(4):1190–1200, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2020:CQR

[133] Zexi Cai and Tony Sit. Censored quantile regression model with timevarying covariates under length-biased sampling. Biometrics, 76(4):1201– 1215, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2020:SST

[134] Moming Li, Guoqing Diao, and Jing Qin. On symmetric semiparametric two-sample problem. Biometrics, 76(4):1216–1228, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ning:2020:SME

[135] Jing Ning, Chunyan Cai, Yong Chen, Xuelin Huang, and Mei-Cheng Wang. Semiparametric modelling and estimation of covariate-adjusted dependence between bivariate recurrent events. Biometrics, 76(4):1229– 1239, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hong:2020:TSS

[136] Chuan Hong, Georgia Salanti, Sally C. Morton, Richard D. Riley, Haitao Chu, Stephen E. Kimmel, and Yong Chen. Testing small study effects in multivariate meta-analysis. Biometrics, 76(4):1240–1250, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [137, 138, 139].

vanHouwelingen:2020:DTS

[137] Hans C. van Houwelingen. Discussion on “Testing small study effects in multivariate meta-analysis” by Chuan Hong, Georgia Salanti, Sally Morton, Richard Riley, Haitao Chu, Stephen E. Kimmel and Yong Chen. Biometrics, 76(4):1251–1254, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [136].

Noma:2020:DTS

[138] Hisashi Noma. Discussion on “Testing small study effects in multivariate meta-analysis” by Chuan Hong, Georgia Salanti, Sally Morton, Richard Riley, Haitao Chu, Stephen E. Kimmel, and Yong Chen. Biometrics, 76(4):1255–1259, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [136].

Carpenter:2020:DTS

[139] James R. Carpenter, Gerta Rücker, and Guido Schwarzer. Discussion on ‘Testing small study effects in multivariate meta-analysis’ by Chuan Hong, Georgia Salanti, Sally Morton, Richard Riley, Haitao Chu, Stephen E Kimmel and Yong Chen. Biometrics, 76(4):1260–1261, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [136].

McCaw:2020:OCR

[140] Zachary R. McCaw, Jacqueline M. Lane, Richa Saxena, Susan Redline, and Xihong Lin. Operating characteristics of the rank-based inverse normal transformation for quantitative trait analysis in genome-wide association studies. Biometrics, 76(4):1262–1272, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Backenroth:2020:NDF

[141] Daniel Backenroth, Russell T. Shinohara, Jennifer A. Schrack, and Jeff Goldsmith. Nonnegative decomposition of functional count data. Biometrics, 76(4):1273–1284, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dorazio:2020:OPD

[142] Robert M. Dorazio. Objective prior distributions for Jolly–Seber models of zero-augmented data. Biometrics, 76(4):1285–1296, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

deCastro:2020:BMR

[143] Mário de Castro, Ming-Hui Chen, Yuanye Zhang, and Anthony V. D’Amico. A Bayesian multi-risks survival (MRS) model in the presence of double censorings. Biometrics, 76(4):1297–1309, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2020:EIT

[144] Xinyang Huang and Jin Xu. Estimating individualized treatment rules with risk constraint. Biometrics, 76(4):1310–1318, December 2020. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2020:REU

[145] Ding-Geng Chen, Dungang Liu, Xiaoyi Min, and Heping Zhang. Relative efficiency of using summary versus individual data in random-effects meta-analysis. Biometrics, 76(4):1319–1329, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Han:2020:VSJ

[146] Dongxiao Han, Xiaogang Su, Liuquan Sun, Zhou Zhang, and Lei Liu. Variable selection in joint frailty models of recurrent and terminal events. Biometrics, 76(4):1330–1339, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baek:2020:USD

[147] Seungchul Baek, Yen-Yi Ho, and Yanyuan Ma. Using sufficient direction factor model to analyze latent activities associated with breast cancer survival. Biometrics, 76(4):1340–1350, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2020:RRC

[148] Feng-Chang Lin, Quefeng Li, and Jessica T. Lin. Relapse or reinfection: Classification of malaria infection using transition likelihoods. Biometrics, 76(4):1351–1363, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2020:PCC

[149] Yimei Li and Ying Yuan. PA-CRM: a continuous reassessment method for pediatric phase I oncology trials with concurrent adult trials. Biometrics, 76(4):1364–1373, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tanaka:2020:SOD

[150] Emi Tanaka. Simple outlier detection for a multi-environmental field trial. Biometrics, 76(4):1374–1382, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baker:2020:CMA

[151] Stuart G. Baker. CACE and meta-analysis (letter to the editor). Biometrics, 76(4):1383–1384, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [152].

Zhou:2020:RCM

[152] Jincheng Zhou, James S. Hodges, and Haitao Chu. Rejoinder to “CACE and meta-analysis (letter to the editor)” by Stuart Baker. Biometrics, 76(4):1385–1389, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [151].

Tattar:2020:BRT

[153] Prabhanjan Narayanachar Tattar. Book review: Time Series Clustering and Classification, Maharaj, Elizabeth Ann, D’Urso, Pierpaolo, and Caiado, Jorge. Boca Raton FL: CRC. Biometrics, 76(4):1390–1391, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McGowan:2020:BRR

[154] Lucy D’Agostino McGowan. Book review: Reproducible Research in R and R Studio, 3rd edition. By Gandrud, Christopher. Biometrics, 76(4):1391, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2020:BRA

[155] Li-Pang Chen. Book review: Artificial intelligence for drug development, precision medicine, and healthcare. Chang, Mark Boca Raton, FL: Chapman and Hall/CRC, 2020. Biometrics, 76(4):1392–1394, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2020:AR

[156] Anonymous. Acknowledgments referees 2020. Biometrics, 76(4):1395– 1398, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2021:IIa

[157] Anonymous. Issue information. Biometrics, 77(1):1–4, March 2021. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2021:RE

[158] Anonymous. Report of the Editors — 2020. Biometrics, 77(1):5–8, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williamson:2021:NVI

[159] Brian D. Williamson, Peter B. Gilbert, Marco Carone, and Noah Simon. Nonparametric variable importance assessment using machine learning techniques. Biometrics, 77(1):9–22, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [160, 161].

Lu:2021:DNV

[160] Min Lu and Hemant Ishwaran. Discussion on “Nonparametric variable importance assessment using machine learning techniques” by Brian D. Williamson, Peter B. Gilbert, Marco Carone, and Noah Simon. Biometrics, 77(1):23–27, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [159].

Williamson:2021:RNV

[161] Brian D. Williamson, Peter B. Gilbert, Marco Carone, and Noah Simon. Rejoinder to “Nonparametric variable importance assessment using machine learning techniques”. Biometrics, 77(1):28–30, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [159, 160].

Feng:2021:APM

[162] Jean Feng, Scott Emerson, and Noah Simon. Approval policies for modifications to machine learning-based software as a medical device: a study of bio-creep. Biometrics, 77(1):31–44, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [163, 164] and rejoinder [165].

Pennello:2021:DAP

[163] Gene Pennello, Berkman Sahiner, Alexej Gossmann, and Nicholas Petrick. Discussion on “Approval policies for modifications to machine learningbased software as a medical device: a study of bio-creep” by Jean Feng, Scott Emerson, and Noah Simon. Biometrics, 77(1):45–48, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [162].

Rose:2021:DAP

[164] Sherri Rose. Discussion on “Approval policies for modifications to machine learning-based software as a medical device: a study of biocreep” by Jean Feng, Scott Emerson, and Noah Simon. Biometrics, 77(1):49–51, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [162].

Feng:2021:RDA

[165] Jean Feng, Scott Emerson, and Noah Simon. Rejoinder to discussions on “Approval policies for modifications to machine learning-based software as a medical device: a study of bio-creep”. Biometrics, 77(1):52–53, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [163, 164, 162].

Yang:2021:AWD

[166] Yuchen Yang and Mei-Cheng Wang. Analyzing wearable device data using marked point processes. Biometrics, 77(1):54–66, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2021:CCE

[167] Lu Wang, Jill Schnall, Aeron Small, Rebecca A. Hubbard, Jason H. Moore, Scott M. Damrauer, and Jinbo Chen. Case contamination in electronic health records based case-control studies. Biometrics, 77(1):67–77, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Luo:2021:BLM

[168] Yu Luo, David A. Stephens, Aman Verma, and David L. Buckeridge. Bayesian latent multi-state modeling for nonequidistant longitudinal electronic health records. Biometrics, 77(1):78–90, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2021:ZIP

[169] Tianchen Xu, Ryan T. Demmer, and Gen Li. Zero-inflated Poisson factor model with application to microbiome read counts. Biometrics, 77(1):91– 101, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2021:RVP

[170] Yukun Liu, Pengfei Li, Lei Song, Kai Yu, and Jing Qin. Retrospective versus prospective score tests for genetic association with case-control data. Biometrics, 77(1):102–112, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2021:CSM

[171] Yudong Wang, Zhi-Sheng Ye, and Hongyuan Cao. On computation of semiparametric maximum likelihood estimators with shape constraints. Biometrics, 77(1):113–124, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Oganisian:2021:BNM

[172] Arman Oganisian, Nandita Mitra, and Jason A. Roy. A Bayesian nonparametric model for zero-inflated outcomes: Prediction, clustering, and causal estimation. Biometrics, 77(1):125–135, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Castelletti:2021:BIC

[173] Federico Castelletti and Guido Consonni. Bayesian inference of causal effects from observational data in Gaussian graphical models. Biometrics, 77(1):136–149, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2021:JMA

[174] Jie Zhou, Xin Chen, Xinyuan Song, and Liuquan Sun. A joint modeling approach for analyzing marker data in the presence of a terminal event. Biometrics, 77(1):150–161, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coulombe:2021:WRA

[175] Janie Coulombe, Erica E. M. Moodie, and Robert W. Platt. Weighted regression analysis to correct for informative monitoring times and confounders in longitudinal studies. Biometrics, 77(1):162–174, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Feifel:2021:DIG

[176] J. Feifel and D. Dobler. Dynamic inference in general nested case-control designs. Biometrics, 77(1):175–185, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Davison:2021:PED

[177] A. C. Davison, S. Hautphenne, and A. Kraus. Parameter estimation for discretely observed linear birth-and-death processes. Biometrics, 77(1): 186–196, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rudolph:2021:TSD

[178] Kara E. Rudolph, Jonathan Levy, and Mark J. van der Laan. Transporting stochastic direct and indirect effects to new populations. Biometrics, 77 (1):197–211, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tian:2021:PPC

[179] Zhaoyang Tian, Kun Liang, and Pengfei Li. A powerful procedure that controls the false discovery rate with directional information. Biometrics, 77(1):212–222, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Clairon:2021:ATR

[180] Q. Clairon, R. Henderson, N. J. Young, E. D. Wilson, and C. J. Taylor. Adaptive treatment and robust control. Biometrics, 77(1):223–236, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alfo:2021:UBE

[181] Marco Alfò, Dankmar Böhning, and Irene Rocchetti. Upper bound estimators of the population size based on ordinal models for capture–recapture experiments. Biometrics, 77(1):237–248, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sugasawa:2021:ESP

[182] Shonosuke Sugasawa and Hisashi Noma. Efficient screening of predictive biomarkers for individual treatment selection. Biometrics, 77(1):249–257, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2021:GRB

[183] Meng Xu, Philip T. Reiss, and Ivor Cribben. Generalized reliability based on distances. Biometrics, 77(1):258–270, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mitani:2021:MAM

[184] A. A. Mitani, E. K. Kaye, and K. P. Nelson. Marginal analysis of multiple outcomes with informative cluster size. Biometrics, 77(1):271–282, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ostrovnaya:2021:TTD

[185] Irina Ostrovnaya, Audrey Mauguen, Venkatraman E. Seshan, and Colin B. Begg. Testing tumors from different anatomic sites for clonal relatedness using somatic mutation data. Biometrics, 77(1):283–292, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jang:2021:ECS

[186] Ja-Yoon Jang, Hee-Seok Oh, Yaeji Lim, and Ying Kuen Cheung. Ensemble clustering for step data via binning. Biometrics, 77(1):293–304, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brownstein:2021:BAS

[187] Naomi C. Brownstein, Veronica Bunn, Luis M. Castro, and Debajyoti Sinha. Bayesian analysis of survival data with missing censoring indicators. Biometrics, 77(1):305–315, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2021:BFA

[188] Jingxiong Xu, Wei Xu, and Laurent Briollais. A Bayes factor approach with informative prior for rare genetic variant analysis from next generation sequencing data. Biometrics, 77(1):316–328, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kreif:2021:ENC

[189] Noémi Kreif, Oleg Sofrygin, Julie A. Schmittdiel, Alyce S. Adams, Richard W. Grant, Zheng Zhu, Mark J. van der Laan, and Romain Neugebauer. Exploiting nonsystematic covariate monitoring to broaden the scope of evidence about the causal effects of adaptive treatment strategies. Biometrics, 77(1):329–342, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Calhoun:2021:RMR

[190] Peter Calhoun, Richard A. Levine, and Juanjuan Fan. Repeated measures random forests (RMRF): Identifying factors associated with nocturnal hypoglycemia. Biometrics, 77(1):343–351, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Polansky:2021:IIN

[191] Leo Polansky, Ken B. Newman, and Lara Mitchell. Improving inference for nonlinear state-space models of animal population dynamics given biased sequential life stage data. Biometrics, 77(1):352–361, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2021:PSE

[192] Ting-Huei Chen and Hanaa Boughal. A penalized structural equation modeling method accounting for secondary phenotypes for variable selection on genetically regulated expression from PrediXcan for Alzheimer’s disease. Biometrics, 77(1):362–371, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gomez-Rubio:2021:BRD

[193] Virgilio Gómez-Rubio. Book review of Disease mapping: From foundations to multidimensional modeling. Biometrics, 77(1):372–373, March 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2021:IIb

[194] Anonymous. Issue information. Biometrics, 77(2):375–378, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tan:2021:EIM

[195] Kean Ming Tan, Junwei Lu, Tong Zhang, and Han Liu. Estimating and inferring the maximum degree of stimulus-locked time-varying brain connectivity networks. Biometrics, 77(2):379–390, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2021:BGS

[196] Kyoungjae Lee and Xuan Cao. Bayesian group selection in logistic regression with application to MRI data analysis. Biometrics, 77(2):391–400, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brown:2021:IME

[197] Roland Brown, Yingling Fan, Kirti Das, and Julian Wolfson. Iterated multisource exchangeability models for individualized inference with an application to mobile sensor data. Biometrics, 77(2):401–412, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2021:RES

[198] David Cheng, Ashwin N. Ananthakrishnan, and Tianxi Cai. Robust and efficient semi-supervised estimation of average treatment effects with application to electronic health records data. Biometrics, 77(2):413–423, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2021:NSM

[199] Kaiqiong Zhao, Karim Oualkacha, Lajmi Lakhal-Chaieb, Aurélie Labbe, Kathleen Klein, Antonio Ciampi, Marie Hudson, Inés Colmegna, Tomi Pastinen, Tieyuan Zhang, Denise Daley, and Celia M. T. Greenwood. A novel statistical method for modeling covariate effects in bisulfite sequencing derived measures of DNA methylation. Biometrics, 77(2):424–438, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kundu:2021:SBM

[200] Suprateek Kundu and Benjamin B. Risk. Scalable Bayesian matrix normal graphical models for brain functional networks. Biometrics, 77(2):439– 450, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2021:IMC

[201] Tong Wang and Wei Ma. The impact of misclassification on covariateadaptive randomized clinical trials. Biometrics, 77(2):451–464, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2021:RBC

[202] Yunan Wu and Lan Wang. Resampling-based confidence intervals for model-free robust inference on optimal treatment regimes. Biometrics, 77(2):465–476, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agniel:2021:ELS

[203] Denis Agniel and Layla Parast. Evaluation of longitudinal surrogate markers. Biometrics, 77(2):477–489, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anaya-Izquierdo:2021:SRS

[204] Karim Anaya-Izquierdo and Neal Alexander. Spatial regression and spillover effects in cluster randomized trials with count outcomes. Biometrics, 77(2):490–505, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2021:CSI

[205] Hyung Park, Eva Petkova, Thaddeus Tarpey, and R. Todd Ogden. A constrained single-index regression for estimating interactions between a treatment and covariates. Biometrics, 77(2):506–518, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2021:MRP

[206] Chixiang Chen, Biyi Shen, Aiyi Liu, Rongling Wu, and Ming Wang. A multiple robust propensity score method for longitudinal analysis with intermittent missing data. Biometrics, 77(2):519–532, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bakoyannis:2021:NAN

[207] Giorgos Bakoyannis. Nonparametric analysis of nonhomogeneous multistate processes with clustered observations. Biometrics, 77(2):533–546, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qadir:2021:SEC

[208] Ghulam A. Qadir and Ying Sun. Semiparametric estimation of crosscovariance functions for multivariate random fields. Biometrics, 77(2): 547–560, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2021:SRC

[209] Ching-Yun Wang and Xiao Song. Semiparametric regression calibration for general hazard models in survival analysis with covariate measurement error; surprising performance under linear hazard. Biometrics, 77(2):561– 572, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2021:SFE

[210] Yan Zhou, Peter X.-K. Song, and Xiaoquan Wen. Structural factor equation models for causal network construction via directed acyclic mixed graphs. Biometrics, 77(2):573–586, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Arfe:2021:OTP

[211] Andrea Arfè, Brian Alexander, and Lorenzo Trippa. Optimality of testing procedures for survival data in the nonproportional hazards setting. Biometrics, 77(2):587–598, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chan:2021:DER

[212] Stephanie Chan, Xuan Wang, Ina Jazić, Sarah Peskoe, Yingye Zheng, and Tianxi Cai. Developing and evaluating risk prediction models with panel current status data. Biometrics, 77(2):599–609, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2021:QRS

[213] Tonghui Yu, Liming Xiang, and Huixia Judy Wang. Quantile regression for survival data with covariates subject to detection limits. Biometrics, 77(2):610–621, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Denti:2021:TGP

[214] Francesco Denti, Michele Guindani, Fabrizio Leisen, Antonio Lijoi, William Duncan Wadsworth, and Marina Vannucci. Two-group Poisson– Dirichlet mixtures for multiple testing. Biometrics, 77(2):622–633, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morita:2021:SBA

[215] Satoshi Morita, Peter Müller, and Hiroyasu Abe. A semiparametric Bayesian approach to population finding with time-to-event and toxicity data in a randomized clinical trial. Biometrics, 77(2):634–648, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gu:2021:BHM

[216] Chenyang Gu, Haiden Huskamp, Julie Donohue, and Sharon-Lise Normand. A Bayesian hierarchical model for characterizing the diffusion of new antipsychotic drugs. Biometrics, 77(2):649–660, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kaplan:2021:BBO

[217] Adam Kaplan and Thomas A. Murray. Batch Bayesian optimization design for optimizing a neurostimulator. Biometrics, 77(2):661–674, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Allen:2021:BMM

[218] Carter Allen, Sara E. Benjamin-Neelon, and Brian Neelon. A Bayesian multivariate mixture model for skewed longitudinal data with intermittent missing observations: an application to infant motor development. Biometrics, 77(2):675–688, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tran:2021:LOU

[219] Trung Dung Tran, Emmanuel Lesaffre, Geert Verbeke, and Joke Duyck. Latent Ornstein–Uhlenbeck models for Bayesian analysis of multivariate longitudinal categorical responses. Biometrics, 77(2):689–701, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hagiwara:2021:SAS

[220] Yasuhiro Hagiwara, Tomohiro Shinozaki, Hirofumi Mukai, and Yutaka Matsuyama. Sensitivity analysis for subsequent treatments in confirmatory oncology clinical trials: a two-stage stochastic dynamic treatment regime approach. Biometrics, 77(2):702–714, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kansanen:2021:HTL

[221] Kasper Kansanen, Petteri Packalen, Matti Maltamo, and Lauri Mehtätalo. Horvitz–Thompson-like estimation with distance-based detection probabilities for circular plot sampling of forests. Biometrics, 77(2):715–728, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2021:EOT

[222] Xiaofei Chen, Daniel F. Heitjan, Gerald Greil, and Haekyung Jeon-Slaughter. Estimating the optimal timing of surgery from observational data. Biometrics, 77(2):729–739, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wen:2021:PGF

[223] Lan Wen, Jessica G. Young, James M. Robins, and Miguel A. Hernán. Parametric g-formula implementations for causal survival analyses. Biometrics, 77(2):740–753, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Su:2021:FLF

[224] Weiji Su, Xia Wang, and Rhonda D. Szczesniak. Flexible link functions in a joint hierarchical Gaussian process model. Biometrics, 77(2):754– 764, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kline:2021:EBO

[225] David Kline and Staci A. Hepler. Estimating the burden of the opioid epidemic for adults and adolescents in Ohio counties. Biometrics, 77(2): 765–775, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2021:BRI

[226] Ran Li and Usama Bilal. Book review: Interactive web-based data visualization with R, plotly, and shiny (Carson Sievert). Biometrics, 77(2): 776–777, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2021:BRS

[227] Chen Hu. Book review: Statistical methods for survival trial design-With applications to cancer clinical trials using R by Jianrong Wu, CRC Press, 2018, ISBN 978-0-367-73432-9. Biometrics, 77(2):777–778, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2021:CEL

[228] Anonymous. Correction to “Empirical-likelihood-based criteria for model selection on marginal analysis of longitudinal data with dropout missingness,” by Chen, C., Shen, B., Zhang, L., Xue, Y. and Wang, M.; 75(3), 950–965, 2019. Biometrics, 77(2):779, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [5].

Chen:2021:CCA

[229] Chixiang Chen, Biyi Shen, Lijun Zhang, Yuan Xue, and Ming Wang. Correction to “A cluster-adjusted rank-based test for a clinical trial concerning multiple endpoints with application to dietary intervention assessment,” by Zhang, W., Liu, A., Tang, L.L. and Li, Q; 75(3), 821–830, 2019. Biometrics, 75(3):780, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [229].

Anonymous:2021:IIc

[230] Anonymous. Issue information. Biometrics, 77(3):781–784, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stringer:2021:ABI

[231] Alex Stringer, Patrick Brown, and Jamie Stafford. Approximate Bayesian inference for case-crossover models. Biometrics, 77(3):785–795, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2021:BAP

[232] Yifei Zhang, Sha Cao, Chi Zhang, Ick Hoon Jin, and Yong Zang. A Bayesian adaptive phase I/II clinical trial design with late-onset competing risk outcomes. Biometrics, 77(3):796–808, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Klein:2021:BVS

[233] Nadja Klein and Michael Stanley Smith. Bayesian variable selection for non-Gaussian responses: a marginally calibrated copula approach. Biometrics, 77(3):809–823, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2021:BCR

[234] Liangliang Zhang, Yushu Shi, Robert R. Jenq, Kim-Anh Do, and Christine B. Peterson. Bayesian compositional regression with structured priors for microbiome feature selection. Biometrics, 77(3):824–838, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carroll:2021:CCR

[235] Cody Carroll, Hans-Georg Müller, and Alois Kneip. Cross-component registration for multivariate functional data, with application to growth curves. Biometrics, 77(3):839–851, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhong:2021:CNG

[236] Qingzhi Zhong, Huazhen Lin, and Yi Li. Cluster non-Gaussian functional data. Biometrics, 77(3):852–865, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Martinez-Hernandez:2021:NTE

[237] Israel Martínez-Hernández and Marc G. Genton. Nonparametric trend estimation in functional time series with application to annual mortality rates. Biometrics, 77(3):866–878, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2021:MND

[238] Yi Zhao, Lexin Li, and Brian S. Caffo. Multimodal neuroimaging data integration and pathway analysis. Biometrics, 77(3):879–889, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gao:2021:RMD

[239] Xu Gao, Weining Shen, Liwen Zhang, Jianhua Hu, Norbert J. Fortin, Ron D. Frostig, and Hernando Ombao. Regularized matrix data clustering and its application to image analysis. Biometrics, 77(3):890–902, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Green:2021:UHD

[240] Brittany Green, Heng Lian, Yan Yu, and Tianhai Zu. Ultra highdimensional semiparametric longitudinal data analysis. Biometrics, 77(3): 903–913, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2021:PFL

[241] Lu Tang and Peter X.-K. Song. Poststratification fusion learning in longitudinal data analysis. Biometrics, 77(3):914–928, September 2021. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Deng:2021:EIP

[242] Yuhao Deng, Chong You, Yukun Liu, Jing Qin, and Xiao-Hua Zhou. Estimation of incubation period and generation time based on observed lengthbiased epidemic cohort with censoring for COVID-19 outbreak in China. Biometrics, 77(3):929–941, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2021:TOP

[243] Rui Chen and Menggang Yu. Tailored optimal posttreatment surveillance for cancer recurrence. Biometrics, 77(3):942–955, September 2021. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2021:ANS

[244] Li-Pang Chen and Grace Y. Yi. Analysis of noisy survival data with graphical proportional hazards measurement error models. Biometrics, 77 (3):956–969, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2021:ECP

[245] Kaitlyn Cook and Rui Wang. Estimation of conditional power for clusterrandomized trials with interval-censored endpoints. Biometrics, 77(3): 970–983, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Srinivasan:2021:CKF

[246] Arun Srinivasan, Lingzhou Xue, and Xiang Zhan. Compositional knockoff filter for high-dimensional regression analysis of microbiome data. Biometrics, 77(3):984–995, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fang:2021:WSA

[247] Fang Fang, Jiwei Zhao, S. Ejaz Ahmed, and Annie Qu. A weak-signal-assisted procedure for variable selection and statistical inference with an informative subsample. Biometrics, 77(3):996–1010, September 2021. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gao:2021:CDF

[248] Xing Gao, Sungwon Lee, Gen Li, and Sungkyu Jung. Covariate-driven factorization by thresholding for multiblock data. Biometrics, 77(3):1011– 1023, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2021:CPC

[249] Ziqi Chen, Jing Ning, Yu Shen, and Jing Qin. Combining primary cohort data with external aggregate information without assuming comparability. Biometrics, 77(3):1024–1036, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hyun:2021:PSI

[250] Sangwon Hyun, Kevin Z. Lin, Max G’Sell, and Ryan J. Tibshirani. Postselection inference for changepoint detection algorithms with application to copy number variation data. Biometrics, 77(3):1037–1049, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2021:MLA

[251] Yang Liu, Yukun Liu, Pengfei Li, and Lin Zhu. Maximum likelihood abundance estimation from capture-recapture data when covariates are missing at random. Biometrics, 77(3):1050–1060, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2021:JPS

[252] Lihui Zhao, Tom Chen, Vladimir Novitsky, and Rui Wang. Joint penalized spline modeling of multivariate longitudinal data, with application to HIV-1 RNA load levels and CD4 cell counts. Biometrics, 77(3):1061–1074, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2021:BEA

[253] Chenguang Wang, Zhixin Wang, Gary L. Rosner, Warner K. Huh, Richard B. S. Roden, and Sejong Bae. A batch-effect adjusted Simon’s two-stage design for cancer vaccine clinical studies. Biometrics, 77(3):1075–1088, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Egleston:2021:SIN

[254] Brian L. Egleston, Tian Bai, Richard J. Bleicher, Stanford J. Taylor, Michael H. Lutz, and Slobodan Vucetic. Statistical inference for natural language processing algorithms with a demonstration using type 2 diabetes prediction from electronic health record notes. Biometrics, 77(3): 1089–1100, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shu:2021:VEI

[255] Di Shu, Jessica G. Young, Sengwee Toh, and Rui Wang. Variance estimation in inverse probability weighted Cox models. Biometrics, 77(3): 1101–1117, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hart:2021:SRL

[256] Kari R. Hart, Teng Fei, and John J. Hanfelt. Scalable and robust latent trajectory class analysis using artificial likelihood. Biometrics, 77(3):1118– 1128, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wei:2021:BRS

[257] Yu-Chung Wei. Book review: Statistics for making decisions by Nicholas T. Longford Boca Raton, FL: Chapman and Hall/CRC, 2021. Hard cover. pp. 307. $96. Biometrics, 77(3):1129, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tu:2021:BRS

[258] Yu-Kang Tu. Book review: Structural equation modeling with partial least squares using Stata and R by Mehmet Mehmetoglu and Sergio Venturini Boca Raton, FL, USA: Chapman and Hall/CRC, 2021. Hard cover. pp.

382. £99.99. Biometrics, 77(3):1130–1131, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dasgupta:2021:BRI

[259] Nairanjana Dasgupta. Book review: Introduction to Data Science: Data Analysis and Algorithms with R, By Rafael Irrizarry Boca Raton, FL: CRC Press, 2020. Hard cover. pp. 711. Biometrics, 77(3):1131–1132, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2021:BRS

[260] Li-Pang Chen. Book review: Statistical Foundations of Data Science, Jianqing Fan, Runze Li, Cun-Hui Zhang and Hui Zou Boca Raton, FL: Chapman and Hall/CRC, 2021. Hard cover. pp. 774. USD 130.00. Biometrics, 77(3):1132–1135, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Whiteman:2021:BRH

[261] Andrew Whiteman. Book review: Handbook of neuroimaging data analysis, by Hernando Ombao, Martin Lindquist, Wesley Thompson, John Aston Boca Raton, FL: CRC Press, 2017. Hard cover. pp. 702. 210.00 USD. Biometrics, 77(3):1135–1137, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2021:IId

[262] Anonymous. Issue information. Biometrics, 77(4):1139–1142, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2021:CMS

[263] Yen-Tsung Huang. Causal mediation of semicompeting risks. Biometrics, 77(4):1143–1154, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [229, 265, 266] and rejoinder [267].

Chan:2021:DCM

[264] Kwun Chuen Gary Chan, Fei Gao, and Fan Xia. Discussion on “Causal mediation of semicompeting risks” by Yen-Tsung Huang. Biometrics, 77(4):1155–1159, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [263, 267].

Stensrud:2021:DCM

[265] Mats J. Stensrud, Jessica G. Young, and Torben Martinussen. Discussion on “Causal mediation of semicompeting risks” by Yen-Tsung Huang. Biometrics, 77(4):1160–1164, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [263, 267].

Fulcher:2021:DCM

[266] Isabel R. Fulcher, Ilya Shpitser, Vanessa Didelez, Kali Zhou, and Daniel O. Scharfstein. Discussion on “Causal mediation of semicompeting risks” by Yen-Tsung Huang. Biometrics, 77(4):1165–1169, December 2021. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [263, 267].

Huang:2021:RCM

[267] Yen-Tsung Huang. Rejoinder to “Causal mediation of semicompeting risks”. Biometrics, 77(4):1170–1174, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [263, 264, 265, 266].

Wang:2021:SPC

[268] Bingkai Wang, Xi Luo, Yi Zhao, and Brian Caffo. Semiparametric partial common principal component analysis for covariance matrices. Biometrics, 77(4):1175–1186, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2021:SMI

[269] Jing Cheng and Dylan S. Small. Semiparametric models and inference for the effect of a treatment when the outcome is nonnegative with clumping at zero. Biometrics, 77(4):1187–1201, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bracher:2021:MMM

[270] Johannes Bracher and Leonhard Held. A marginal moment matching approach for fitting endemic-epidemic models to underreported disease surveillance counts. Biometrics, 77(4):1202–1214, December 2021. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Helgeson:2021:NCS

[271] Erika S. Helgeson, David M. Vock, and Eric Bair. Nonparametric cluster significance testing with reference to a unimodal null distribution. Biometrics, 77(4):1215–1226, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2021:NMR

[272] Wei Hu, Tianyu Pan, Dehan Kong, and Weining Shen. Nonparametric matrix response regression with application to brain imaging data analysis. Biometrics, 77(4):1227–1240, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hejazi:2021:ENI

[273] Nima S. Hejazi, Mark J. van der Laan, Holly E. Janes, Peter B. Gilbert, and David C. Benkeser. Efficient nonparametric inference on the effects of stochastic interventions under two-phase sampling, with applications to vaccine efficacy trials. Biometrics, 77(4):1241–1253, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2021:NBI

[274] Yiwang Zhou, Peter X. K. Song, and Haoda Fu. Net benefit index: Assessing the influence of a biomarker for individualized treatment rules. Biometrics, 77(4):1254–1264, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2021:CPW

[275] Lu Mao and Tuo Wang. A class of proportional win-fractions regression models for composite outcomes. Biometrics, 77(4):1265–1275, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2021:EIM

[276] Ruoqi Yu. Evaluating and improving a matched comparison of antidepressants and bone density. Biometrics, 77(4):1276–1288, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Boussari:2021:MEH

[277] Olayidé Boussari, Laurent Bordes, Gaëlle Romain, Marc Colonna, Nadine Bossard, Laurent Remontet, and Valérie Jooste. Modeling excess hazard with time-to-cure as a parameter. Biometrics, 77(4):1289–1302, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2021:PPL

[278] Peijie Wang, Danning Li, and Jianguo Sun. A pairwise pseudo-likelihood approach for left-truncated and interval-censored data under the Cox model. Biometrics, 77(4):1303–1314, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Parast:2021:EMS

[279] Layla Parast, Tianxi Cai, and Lu Tian. Evaluating multiple surrogate markers with censored data. Biometrics, 77(4):1315–1327, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dai:2021:MSL

[280] Xiongtao Dai, Zhenhua Lin, and Hans-Georg Müller. Modeling sparse longitudinal data on Riemannian manifolds. Biometrics, 77(4):1328–1341, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beesley:2021:SAC

[281] Lauren J. Beesley and Jeremy M. G. Taylor. A stacked approach for chained equations multiple imputation incorporating the substantive model. Biometrics, 77(4):1342–1354, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2021:PFC

[282] Peng Wang, Siqi Xu, Yi-Xin Wang, Baolin Wu, Wing Kam Fung, Guimin Gao, Zhijiang Liang, and Nianjun Liu. Penalized Fieller’s confidence interval for the ratio of bivariate normal means. Biometrics, 77(4):1355–1368, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kenney:2021:PPP

[283] Toby Kenney, Hong Gu, and Tianshu Huang. Poisson PCA: Poisson measurement error corrected PCA, with application to microbiome data. Biometrics, 77(4):1369–1384, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2021:RCM

[284] Lin Zhang, Andrew DiLernia, Karina Quevedo, Jazmin Camchong, Kelvin Lim, and Wei Pan. A random covariance model for bi-level graphical modeling with application to resting-state fMRI data. Biometrics, 77(4): 1385–1396, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2021:HIB

[285] Baihua He, Tingyan Zhong, Jian Huang, Yanyan Liu, Qingzhao Zhang, and Shuangge Ma. Histopathological imaging-based cancer heterogeneity analysis via penalized fusion with model averaging. Biometrics, 77(4):

1397–1408, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ji:2021:BCA

[286] Jiadong Ji, Yong He, Lei Liu, and Lei Xie. Brain connectivity alteration detection via matrix-variate differential network model. Biometrics, 77(4): 1409–1421, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Luckett:2021:ROC

[287] Daniel J. Luckett, Eric B. Laber, Samer S. El-Kamary, Cheng Fan, Ravi Jhaveri, Charles M. Perou, Fatma M. Shebl, and Michael R. Kosorok. Receiver operating characteristic curves and confidence bands for support vector machines. Biometrics, 77(4):1422–1430, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2021:BAR

[288] Zhenke Wu, Livia Casciola-Rosen, Antony Rosen, and Scott L. Zeger. A Bayesian approach to restricted latent class models for scientifically structured clustering of multivariate binary outcomes. Biometrics, 77(4): 1431–1444, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chakraborty:2021:UHG

[289] Saptarshi Chakraborty, Colin B. Begg, and Ronglai Shen. Using the “Hidden” genome to improve classification of cancer types. Biometrics, 77(4): 1445–1455, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wilson:2021:CME

[290] Katie Wilson and Jon Wakefield. Child mortality estimation incorporating summary birth history data. Biometrics, 77(4):1456–1466, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Benkeser:2021:IPP

[291] David Benkeser, Iván Díaz, Alex Luedtke, Jodi Segal, Daniel Scharfstein, and Michael Rosenblum. Improving precision and power in randomized trials for COVID-19 treatments using covariate adjustment, for binary, ordinal, and time-to-event outcomes. Biometrics, 77(4):1467–1481, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [292, 293, 294] and rejoinder [295].

Proschan:2021:DIP

[292] Michael A. Proschan. Discussion on “Improving precision and power in randomized trials for COVID-19 treatments using covariate adjustment for binary, ordinal, and time-to-event outcomes”. Biometrics, 77(4):1482– 1484, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [291, 295].

Zhang:2021:DIP

[293] Min Zhang and Baqun Zhang. Discussion of “Improving precision and power in randomized trials for COVID-19 treatments using covariate adjustment, for binary, ordinal, and time-to-event outcomes”. Biometrics, 77(4):1485–1488, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [291, 295].

LaVange:2021:DIP

[294] Lisa M. LaVange. Discussion on “Improving precision and power in randomized trials for COVID-19 treatments using covariate adjustment, for binary, ordinal, and time-to-event outcomes” by David Benkeser, Ivan Diaz, Alex Luedtke, Jodi Segal, Daniel Scharfstein, and Michael Rosenblum. Biometrics, 77(4):1489–1491, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [291, 295].

Benkeser:2021:RIP

[295] David Benkeser, Iván Díaz, Alex Luedtke, Jodi Segal, Daniel Scharfstein, and Michael Rosenblum. Rejoinder: Improving precision and power in randomized trials for COVID-19 treatments using covariate adjustment, for binary, ordinal, and time-to-event outcomes. Biometrics, 77(4):1492– 1494, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [291, 292, 293, 294].

Mukherjee:2021:BRR

[296] Rajarshi Mukherjee. Book review: Replication and evidence factors in observational studies by Paul R. Rosenbaum Boca Raton, FL: Chapman and Hall/CRC, 2021. pp. 276. Biometrics, 77(4):1495–1498, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2021:BRM

[297] Lin Liu. Book review: Matrix-based introduction to multivariate data analysis, by Kohei Adachi, 2nd edition. Singapore: Springer Nature, 2020. pp. 457. Biometrics, 77(4):1498–1500, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eskelson:2021:BRB

[298] Bianca N. I. Eskelson. Book review: Biometry for forestry and environmental data with examples in R. Mehtätalo, Lauri and Lappi, Juha. Boca Raton, FL: Chapman and Hall/CRC, 2020. pp. 426. Biometrics, 77(4): 1500–1502, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2021:BRD

[299] Mei-Hsien Lee. Book review: Data analysis with RStudio: an easygoing introduction by Franz Kronthaler and Silke Zöllner Berlin, Germany: Springer-Verlag GmbH, DE, 2021. pp. 131. Biometrics, 77(4):1502–1503, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lai:2021:BRD

[300] Yin-Ju Lai and Chuhsing Kate Hsiao. Book review: Data science and machine learning: Mathematical and statistical methods by Dirk P. Kroese, Zdravko Botev, Thomas Taimre, Radislav Vaisman Boca Raton, FL: Chapman and Hall/CRC, 2019. pp. 523. Biometrics, 77(4):1503–1504, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2021:AR

[301] Anonymous. Acknowledgments referees 2021. Biometrics, 77(4):1505– 1508, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2022:IIa

[302] Anonymous. Issue information. Biometrics, 78(1):1–4, March 2022. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2022:RE

[303] Anonymous. Report of the editors — 2021. Biometrics, 78(1):5–8, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Parast:2022:RMC

[304] Layla Parast, Tanya P. Garcia, Ross L. Prentice, and Raymond J. Carroll. Robust methods to correct for measurement error when evaluating a surrogate marker. Biometrics, 78(1):9–23, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kang:2022:TRT

[305] Hyunseung Kang, Youjin Lee, T. Tony Cai, and Dylan S. Small. Two robust tools for inference about causal effects with invalid instruments. Biometrics, 78(1):24–34, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alonso:2022:IIP

[306] Ariel Alonso, Wim Van der Elst, Lizet Sanchez, Patricia Luaces, and Geert Molenberghs. Identifying individual predictive factors for treatment efficacy. Biometrics, 78(1):35–45, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Loh:2022:NMA

[307] Wen Wei Loh, Beatrijs Moerkerke, Tom Loeys, and Stijn Vansteelandt. Nonlinear mediation analysis with high-dimensional mediators whose causal structure is unknown. Biometrics, 78(1):46–59, March 2022. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2022:BGS

[308] Yeonhee Park, Suyu Liu, Peter F. Thall, and Ying Yuan. Bayesian group sequential enrichment designs based on adaptive regression of response and survival time on baseline biomarkers. Biometrics, 78(1):60–71, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2022:SBL

[309] Cui Guo, Jian Kang, and Timothy D. Johnson. A spatial Bayesian latent factor model for image-on-image regression. Biometrics, 78(1):72–84, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hui:2022:AIS

[310] Francis K. C. Hui, Nicole A. Hill, and A. H. Welsh. Assuming independence in spatial latent variable models: Consequences and implications of misspecification. Biometrics, 78(1):85–99, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Antonelli:2022:CIH

[311] Joseph Antonelli, Georgia Papadogeorgou, and Francesca Dominici. Causal inference in high dimensions: a marriage between Bayesian modeling and good frequentist properties. Biometrics, 78(1):100–114, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yiu:2022:JCE

[312] Sean Yiu and Li Su. Joint calibrated estimation of inverse probability of treatment and censoring weights for marginal structural models. Biometrics, 78(1):115–127, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2022:OOC

[313] Yuxin Zhu and Mei-Cheng Wang. Obtaining optimal cutoff values for tree classifiers using multiple biomarkers. Biometrics, 78(1):128–140, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2022:TSP

[314] Jixiong Wang, Ashish Patel, James M. S. Wason, and Paul J. Newcombe. Two-stage penalized regression screening to detect biomarkertreatment interactions in randomized clinical trials. Biometrics, 78(1): 141–150, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yi:2022:SVS

[315] Fengting Yi, Niansheng Tang, and Jianguo Sun. Simultaneous variable selection and estimation for joint models of longitudinal and failure time data with interval censoring. Biometrics, 78(1):151–164, March 2022. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2022:ACI

[316] Chun Yin Lee, Kin Yau Wong, K. F. Lam, and Jinfeng Xu. Analysis of clustered interval-censored data using a class of semiparametric partly linear frailty transformation models. Biometrics, 78(1):165–178, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shin:2022:WCI

[317] Yei Eun Shin, Ruth M. Pfeiffer, Barry I. Graubard, and Mitchell H. Gail. Weight calibration to improve efficiency for estimating pure risks from the additive hazards model with the nested case-control design. Biometrics, 78(1):179–191, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhong:2022:RMS

[318] Yingchao Zhong and Douglas E. Schaubel. Restricted mean survival time as a function of restriction time. Biometrics, 78(1):192–201, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2022:NMT

[319] Hao Chen and Dylan S. Small. New multivariate tests for assessing covariate balance in matched observational studies. Biometrics, 78(1):202–213, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beesley:2022:SIA

[320] Lauren J. Beesley and Bhramar Mukherjee. Statistical inference for association studies using electronic health records: handling both selection bias and outcome misclassification. Biometrics, 78(1):214–226, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2022:SIU

[321] Danhyang Lee and Jae Kwang Kim. Semiparametric imputation using conditional Gaussian mixture models under item nonresponse. Biometrics, 78(1):227–237, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

AlMohamad:2022:SCI

[322] Diaa Al Mohamad, Jelle J. Goeman, and Erik W. van Zwet. Simultaneous confidence intervals for ranks with application to ranking institutions. Biometrics, 78(1):238–247, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DiMarzio:2022:DEC

[323] Marco Di Marzio, Stefania Fensore, Agnese Panzera, and Charles C. Taylor. Density estimation for circular data observed with errors. Biometrics, 78(1):248–260, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Deng:2022:SMC

[324] Yangqing Deng, Yinqiu He, Gongjun Xu, and Wei Pan. Speeding up Monte Carlo simulations for the adaptive sum of powered score test with importance sampling. Biometrics, 78(1):261–273, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Borchers:2022:LCH

[325] David L. Borchers, Peter Nightingale, Ben C. Stevenson, and Rachel M. Fewster. A latent capture history model for digital aerial surveys. Biometrics, 78(1):274–285, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Niu:2022:MGM

[326] Mu Niu, Fay Frost, Jordan E. Milner, Anna Skarin, and Paul G. Blackwell. Modelling group movement with behaviour switching in continuous time. Biometrics, 78(1):286–299, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gerard:2022:BDR

[327] Emma Gerard, Sarah Zohar, Hoai-Thu Thai, Christelle Lorenzato, Marie-Karelle Riviere, and Moreno Ursino. Bayesian dose regimen assessment in early phase oncology incorporating pharmacokinetics and pharmacodynamics. Biometrics, 78(1):300–312, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hart:2022:NBM

[328] Brian Hart, Michele Guindani, Stephen Malone, and Mark Fiecas. A nonparametric Bayesian model for estimating spectral densities of restingstate EEG twin data. Biometrics, 78(1):313–323, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choi:2022:SSS

[329] Young-Geun Choi, Lawrence P. Hanrahan, Derek Norton, and Ying-Qi Zhao. Simultaneous spatial smoothing and outlier detection using penalized regression, with application to childhood obesity surveillance from electronic health records. Biometrics, 78(1):324–336, March 2022. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2022:EOI

[330] Yizhe Xu, Tom H. Greene, Adam P. Bress, Brian C. Sauer, Brandon K. Bellows, Yue Zhang, William S. Weintraub, Andrew E. Moran, and Jincheng Shen. Estimating the optimal individualized treatment rule from a cost-effectiveness perspective. Biometrics, 78(1):337–351, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Blanche:2022:CMT

[331] Paul Blanche, Jean-François Dartigues, and Jérémie Riou. A closed max-t test for multiple comparisons of areas under the ROC curve. Biometrics, 78(1):352–363, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhong:2022:GMS

[332] Wujuan Zhong, Toni Darville, Xiaojing Zheng, Jason Fine, and Yun Li. Generalized multi-SNP mediation intersection-union test. Biometrics, 78 (1):364–375, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Isheden:2022:REM

[333] Gabriel Isheden, Kamila Czene, and Keith Humphreys. Random effects models of lymph node metastases in breast cancer: quantifying the roles of covariates and screening using a continuous growth model. Biometrics, 78(1):376–387, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shook-Sa:2022:PSS

[334] Bonnie E. Shook-Sa and Michael G. Hudgens. Power and sample size for observational studies of point exposure effects. Biometrics, 78(1):388– 398, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vonesh:2022:BES

[335] Edward F. Vonesh and Tom Greene. Biased estimation with shared parameter models in the presence of competing dropout mechanisms. Biometrics, 78(1):399–406, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [336].

Thomadakis:2022:RBE

[336] Christos Thomadakis, Loukia Meligkotsidou, Nikos Pantazis, and Giota Touloumi. Rejoinder to “Biased Estimation With Shared Parameter Models in the Presence of Competing Dropout Mechanisms”. Biometrics, 78 (1):407–408, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [335].

Chang:2022:BRM

[337] Howard H. Chang. Book review: Modelling spatial and spatial-temporal data: a Bayesian approach, Haining, Robert P. and Li, Guangquan Boca Raton, FL: Chapman and Hall/CRC, 2020. pp. 608. Biometrics, 78(1): 409–410, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fusfeld:2022:BRS

[338] Zachary Fusfeld. Book review: Statistical inference via data science: a modern dive into R and the Tidyverse, Ismay, Chester and Kim, Albert Y. Boca Raton, FL: Chapman and Hall/CRC, 2019. pp. 460. Biometrics, 78(1):410–412, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ozanne:2022:BRB

[339] Marie V. Ozanne. Book review: Bayesian analysis of infectious diseases — COVID-19 and beyond, by Lyle Broemeling Broemeling, Lyle D. Boca Raton, FL: Chapman and Hall/CRC, 2021. pp. 330. Biometrics, 78(1): 412–413, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Akkaya-Hocagil:2022:BRM

[340] Tugba Akkaya-Hocagil. Book review: Meta-analysis: Methods for health and experimental studies: Khan, Shahjahan Springer Nature Singapore Pte. Ltd. 2020. pp. 293. Biometrics, 78(1):414, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2022:BRO

[341] Amber W. Wang. Book review: Omic association studies with R and bioconductor: Gonzalez, Juan R. and Caceres, Alejandro Boca Raton, FL: Chapman and Hall/CRC, 2020. pp. 390. Biometrics, 78(1):415–416, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2022:IIb

[342] Anonymous. Issue information. Biometrics, 78(2):417–420, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eden:2022:NES

[343] Svetlana K. Eden, Chun Li, and Bryan E. Shepherd. Nonparametric estimation of Spearman’s rank correlation with bivariate survival data. Biometrics, 78(2):421–434, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2022:JMS

[344] Cai Li, Luo Xiao, and Sheng Luo. Joint model for survival and multivariate sparse functional data with application to a study of Alzheimer’s Disease. Biometrics, 78(2):435–447, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Han:2022:SEN

[345] Bo Han, Ingrid Van Keilegom, and Xiaoguang Wang. Semiparametric estimation of the nonmixture cure model with auxiliary survival information. Biometrics, 78(2):448–459, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rennert:2022:CRM

[346] Lior Rennert and Sharon X. Xie. Cox regression model under dependent truncation. Biometrics, 78(2):460–473, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hung:2022:VCF

[347] Ying Hung, Li-Hsiang Lin, and C. F. Jeff Wu. Varying coefficient frailty models with applications in single molecular experiments. Biometrics, 78(2):474–486, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2022:GRA

[348] Lin Zhang and Lei Sun. A generalized robust allele-based genetic association test. Biometrics, 78(2):487–498, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2022:PRS

[349] Bingxin Zhao and Fei Zou. On polygenic risk scores for complex traits prediction. Biometrics, 78(2):499–511, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yi:2022:IIG

[350] Huangdi Yi, Qingzhao Zhang, Cunjie Lin, and Shuangge Ma. Informationincorporated Gaussian graphical model for gene expression data. Biometrics, 78(2):512–523, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ren:2022:GGM

[351] Mingyang Ren, Sanguo Zhang, Qingzhao Zhang, and Shuangge Ma. Gaussian graphical model-based heterogeneity analysis via penalized fusion. Biometrics, 78(2):524–535, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Geng:2022:BSH

[352] Lijiang Geng and Guanyu Hu. Bayesian spatial homogeneity pursuit for survival data with an application to the SEER respiratory cancer data. Biometrics, 78(2):536–547, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lan:2022:GMP

[353] Zhou Lan, Brian J. Reich, Joseph Guinness, Dipankar Bandyopadhyay, Liangsuo Ma, and F. Gerard Moeller. Geostatistical modeling of positivedefinite matrices: an application to diffusion tensor imaging. Biometrics, 78(2):548–559, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2022:SFM

[354] Lu Zhang and Sudipto Banerjee. Spatial factor modeling: a Bayesian matrix-normal approach for misaligned data. Biometrics, 78(2):560–573, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2022:SEC

[355] Yujia Li, Xiangrui Zeng, Chien-Wei Lin, and George C. Tseng. Simultaneous estimation of cluster number and feature sparsity in high-dimensional cluster analysis. Biometrics, 78(2):574–585, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2022:KRM

[356] Yang Wang and Zhangsheng Yu. A kernel regression model for panel count data with nonparametric covariate functions. Biometrics, 78(2): 586–597, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hwang:2022:MAC

[357] Wen-Han Hwang, Richard Huggins, and Jakub Stoklosa. A model for analyzing clustered occurrence data. Biometrics, 78(2):598–611, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Safo:2022:SLD

[358] Sandra E. Safo, Eun Jeong Min, and Lillian Haine. Sparse linear discriminant analysis for multiview structured data. Biometrics, 78(2):612–623, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2022:CIT

[359] Jiabei Yang, Issa J. Dahabreh, and Jon A. Steingrimsson. Causal interaction trees: Finding subgroups with heterogeneous treatment effects in observational data. Biometrics, 78(2):624–635, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mountain:2022:RPM

[360] Rachael Mountain and Chris Sherlock. Recruitment prediction for multicenter clinical trials based on a hierarchical Poisson-gamma model:

Asymptotic analysis and improved intervals. Biometrics, 78(2):636–648, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scharfstein:2022:GSA

[361] Daniel O. Scharfstein, Jon Steingrimsson, Aidan McDermott, Chenguang Wang, Souvik Ray, Aimee Campbell, Edward Nunes, and Abigail Matthews. Global sensitivity analysis of randomized trials with nonmonotone missing binary outcomes: Application to studies of substance use disorders. Biometrics, 78(2):649–659, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liao:2022:VRI

[362] Jiangang Liao and Charles Rohde. Variance reduction in the inverse probability weighted estimators for the average treatment effect using the propensity score. Biometrics, 78(2):660–667, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2022:RBN

[363] Lan Liu and Eric Tchetgen Tchetgen. Regression-based negative control of homophily in dyadic peer effect analysis. Biometrics, 78(2):668–678, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sheng:2022:SEA

[364] Ying Sheng, Yifei Sun, Chiung-Yu Huang, and Mi-Ok Kim. Synthesizing external aggregated information in the presence of population heterogeneity: a penalized empirical likelihood approach. Biometrics, 78(2):679–690, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2022:DEO

[365] Ruoyu Wang and Qihua Wang. Determination and estimation of optimal quarantine duration for infectious diseases with application to data analysis of COVID-19. Biometrics, 78(2):691–700, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sauer:2022:SSI

[366] Sara Sauer, Bethany Hedt-Gauthier, Claudia Rivera-Rodriguez, and Sebastien Haneuse. Small-sample inference for cluster-based outcomedependent sampling schemes in resource-limited settings: Investigating low birthweight in rwanda. Biometrics, 78(2):701–715, June 2022. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2022:PSH

[367] Yan Li, Chun Yu, Yize Zhao, Weixin Yao, Robert H. Aseltine, and Kun Chen. Pursuing sources of heterogeneity in modeling clustered population. Biometrics, 78(2):716–729, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Comment:2022:BDF

[368] Leah Comment, Brent A. Coull, Corwin Zigler, and Linda Valeri. Bayesian data fusion: Probabilistic sensitivity analysis for unmeasured confounding using informative priors based on secondary data. Biometrics, 78(2):730– 741, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2022:BSM

[369] Yin Song, Shufei Ge, Jiguo Cao, Liangliang Wang, and Farouk S. Nathoo. A Bayesian spatial model for imaging genetics. Biometrics, 78(2):742– 753, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Marks-Anglin:2022:EEB

[370] Arielle Marks-Anglin, Chongliang Luo, Jin Piao, Mary Beth Connolly Gibbons, Christopher H. Schmid, Jing Ning, and Yong Chen. EMBRACE: an EM-based bias reduction approach through copas-model estimation for quantifying the evidence of selective publishing in network meta-analysis. Biometrics, 78(2):754–765, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2022:MDC

[371] Zhen Yang and Yen-Yi Ho. Modeling dynamic correlation in zero-inflated bivariate count data with applications to single-cell RNA sequencing data. Biometrics, 78(2):766–776, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chakladar:2022:IPW

[372] Sujatro Chakladar, Samuel Rosin, Michael G. Hudgens, M. Elizabeth Halloran, John D. Clemens, Mohammad Ali, and Michael E. Emch. Inverse probability weighted estimators of vaccine effects accommodating partial interference and censoring. Biometrics, 78(2):777–788, June 2022. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schaarschmidt:2022:TTT

[373] Frank Schaarschmidt, Christian Ritz, and Ludwig A. Hothorn. The Tukey trend test: Multiplicity adjustment using multiple marginal models. Biometrics, 78(2):789–797, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Benedetti:2022:IRI

[374] Marco H. Benedetti, Veronica J. Berrocal, and Naveen N. Narisetty. Identifying regions of inhomogeneities in spatial processes via an M-RA and mixture priors. Biometrics, 78(2):798–811, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gasparrini:2022:CPF

[375] Antonio Gasparrini, Fabian Scheipl, Ben Armstrong, and Michael G. Kenward. Correction to “A penalized framework for distributed lag nonlinear models” by Antonio Gasparrini, Fabian Scheipl, Ben Armstrong, and Michael G. Kenward; 73, 938–948, September 2017. Biometrics, 78 (2):812, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [3].

Dasgupta:2022:BRB

[376] Sayan Dasgupta. Book review: Bayesian compendium, Marcel Oijen, Cham: Springer Nature Switzerland AG. 2020. xiv, 204 p. Biometrics, 73(3):813–815, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wen:2022:BRU

[377] Shu-Hui Wen. Book review: Using R for biostatistics. T. W. MacFarland and J. M. Yates. Springer International Publishing. 2021. ISBN: 978-3-030-62403-3; 978-3-030-62404-0 (eBook). Biometrics, 73(3):815–816, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2022:BRP

[378] Yu-Jyun Huang. Book review: Predictive analytics: Parametric models for regression and classification using R, John Wiley & Sons. 2020. ISBN: 978-1-118-94889-7; 978-1-118-94890-3 (ebook). Biometrics, 73(3):816–817, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roychoudhury:2022:BRC

[379] Satrajit Roychoudhury. Book review: Cure models: Methods, applications, and implementation, Yingwei Peng Binbing Yu Boca Raton, FL: Chapman and Hall/CRC, 2021. pp. 268. Biometrics, 73(3):817, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2022:BRHa

[380] Li-Pang Chen. Book review: Handbook of meta-analysis Boca Raton, FL: Chapman and Hall/CRC. 2021. ISBN 978-1-138-10640-6, pp. 592. Biometrics, 73(3):819–820, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2022:IIc

[381] Anonymous. Issue information. Biometrics, 78(3):821–424, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsiatis:2022:EVE

[382] Anastasios A. Tsiatis and Marie Davidian. Estimating vaccine efficacy over time after a randomized study is unblinded. Biometrics, 78(3):825– 838, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [383, 384, 385] and rejoinder [386].

Halloran:2022:DEV

[383] M. Elizabeth Halloran. Discussion on “Estimating vaccine efficacy over time after a randomized study is unblinded” by Anastasios A. Tsiatis and Marie Davidian. Biometrics, 78(3):839–840, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [382].

Janes:2022:DEV

[384] Holly Janes, Fei Gao, and Alex Luedtke. Discussion on “Estimating vaccine efficacy over time after a randomized study is unblinded” by Anastasios A. Tsiatis and Marie Davidian. Biometrics, 78(3):841–843, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [382].

Follmann:2022:DEV

[385] Dean Follmann. Discussion on “Estimating vaccine efficacy over time after a randomized study is unblinded” by Anastasios A. Tsiatis and Marie Davidian. Biometrics, 78(3):844–847, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [382].

Tsiatis:2022:REV

[386] Anastasios A. Tsiatis and Marie Davidian. Rejoinder: Estimating vaccine efficacy over time after a randomized study is unblinded. Biometrics, 78 (3):848–851, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [382, 383, 384, 385].

Wang:2022:MSA

[387] Wei Wang, Shou-En Lu, Jerry Q. Cheng, Minge Xie, and John B. Kostis. Multivariate survival analysis in big data: a divide-and-combine approach. Biometrics, 78(3):852–866, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Edelmann:2022:CVD

[388] Dominic Edelmann, Thomas Welchowski, and Axel Benner. A consistent version of distance covariance for right-censored survival data and its application in hypothesis testing. Biometrics, 78(3):867–879, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Basak:2022:SAC

[389] Piyali Basak, Antonio Linero, Debajyoti Sinha, and Stuart Lipsitz. Semiparametric analysis of clustered interval-censored survival data using soft Bayesian additive regression trees (SBART). Biometrics, 78(3):880–893, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yi:2022:FSL

[390] Grace Y. Yi, Wenqing He, and Raymond. J. Carroll. Feature screening with large-scale and high-dimensional survival data. Biometrics, 78(3): 894–907, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morrison:2022:RIC

[391] Doug Morrison, Oliver Laeyendecker, and Ron Brookmeyer. Regression with interval-censored covariates: Application to cross-sectional incidence estimation. Biometrics, 78(3):908–921, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nevo:2022:MSC

[392] Daniel Nevo, Deborah Blacker, Eric B. Larson, and Sebastien Haneuse. Modeling semi-competing risks data as a longitudinal bivariate process. Biometrics, 78(3):922–936, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2022:SES

[393] Shu Yang. Semiparametric estimation of structural nested mean models with irregularly spaced longitudinal observations. Biometrics, 78(3):937– 949, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2022:SMB

[394] J. Liu, Xinlian Zhang, T. Chen, T. Wu, T. Lin, L. Jiang, S. Lang, L. Liu, L. Natarajan, J. X. Tu, T. Kosciolek, J. Morton, T. T. Nguyen, B. Schnabl, R. Knight, C. Feng, Y. Zhong, and X. M. Tu. A semiparametric model for between-subject attributes: Applications to beta-diversity of microbiome data. Biometrics, 78(3):950–962, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stevenson:2022:SCS

[395] Ben C. Stevenson, Rachel M. Fewster, and Koustubh Sharma. Spatial correlation structures for detections of individuals in spatial capturerecapture models. Biometrics, 78(3):963–973, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fiksel:2022:TFL

[396] Jacob Fiksel, Scott Zeger, and Abhirup Datta. A transformation-free linear regression for compositional outcomes and predictors. Biometrics, 78 (3):974–987, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jin:2022:BNA

[397] Wei Jin, Yang Ni, Leah H. Rubin, Amanda B. Spence, and Yanxun Xu. A Bayesian nonparametric approach for inferring drug combination effects on mental health in people with HIV. Biometrics, 78(3):988–1000, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Linero:2022:SBE

[398] Antonio R. Linero. Simulation-based estimators of analytically intractable causal effects. Biometrics, 78(3):1001–1017, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gao:2022:TAM

[399] Lucy L. Gao, Daniela Witten, and Jacob Bien. Testing for association in multiview network data. Biometrics, 78(3):1018–1030, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Luo:2022:RFF

[400] Ruiyan Luo and Xin Qi. Restricted function-on-function linear regression model. Biometrics, 78(3):1031–1044, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gao:2022:APH

[401] Xu Gao, Weining Shen, Jing Ning, Ziding Feng, and Jianhua Hu. Addressing patient heterogeneity in disease predictive model development. Biometrics, 78(3):1045–1055, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

David:2022:OSD

[402] Olivier David, Arnaud Le Rouzic, and Christine Dillmann. Optimization of sampling designs for pedigrees and association studies. Biometrics, 78(3):1056–1066, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Deng:2022:TEM

[403] Kai Deng and Xin Zhang. Tensor envelope mixture model for simultaneous clustering and multiway dimension reduction. Biometrics, 78(3):1067– 1079, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2022:DIC

[404] Ben Wu, Subhadip Pal, Jian Kang, and Ying Guo. Distributional independent component analysis for diverse neuroimaging modalities. Biometrics, 78(3):1092–1105, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [406, 407, 408, 409] and rejoinder [410].

Zhang:2022:IAM

[405] Han Zhang, Lu Deng, William Wheeler, Jing Qin, and Kai Yu. Integrative analysis of multiple case-control studies. Biometrics, 78(3):1080–1091, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shappell:2022:DDI

[406] Heather Shappell and Sean L. Simpson. Discussion on “Distributional independent component analysis for diverse neuroimaging modalities” by Ben Wu, Subhadip Pal, Jian Kang, and Ying Guo. Biometrics, 78(3): 1106–1108, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [404].

Mejia:2022:DDI

[407] Amanda F. Mejia. Discussion on “Distributional independent component analysis for diverse neuroimaging modalities” by Ben Wu, Subhadip Pal, Jian Kang, and Ying Guo. Biometrics, 78(3):1109–1112, September 2022.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [406].

Keeratimahat:2022:DDI

[408] Kan Keeratimahat and Thomas E. Nichols. Discussion on “Distributional independent component analysis for diverse neuroimaging modalities” by Ben Wu, Subhadip Pal, Jian Kang, and Ying Guo. Biometrics, 78(3): 1113–1117, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [406].

Moerkerke:2022:DDI

[409] Beatrijs Moerkerke and Ruth Seurinck. Discussion on “Distributional independent component analysis for diverse neuroimaging modalities” by Ben Wu, Subhadip Pal, Jian Kang, and Ying Guo. Biometrics, 78(3): 1118–1121, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [406].

Wu:2022:RDD

[410] Ben Wu, Subhadip Pal, Jian Kang, and Ying Guo. Rejoinder to discussions of “Distributional independent component analysis for diverse neuroimaging modalities”. Biometrics, 78(3):1122–1126, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [404, 406, 406, 407, 408, 409].

Wood:2022:IUC

[411] Simon N. Wood. Inferring UK COVID-19 fatal infection trajectories from daily mortality data: Were infections already in decline before the UK lockdowns? Biometrics, 78(3):1127–1140, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baldoni:2022:EDC

[412] Pedro L. Baldoni, Naim U. Rashid, and Joseph G. Ibrahim. Efficient detection and classification of epigenomic changes under multiple conditions. Biometrics, 78(3):1141–1154, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jiang:2022:USC

[413] Lingjing Jiang, Niina Haiminen, Anna-Paola Carrieri, Shi Huang, Yoshiki Vázquez-Baeza, Laxmi Parida, Ho-Cheol Kim, Austin D. Swafford, Rob Knight, and Loki Natarajan. Utilizing stability criteria in choosing feature selection methods yields reproducible results in microbiome data. Biometrics, 78(3):1155–1167, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eaton:2022:NEI

[414] Anne Eaton, Yifei Sun, James Neaton, and Xianghua Luo. Nonparametric estimation in an illness-death model with component-wise censoring. Biometrics, 78(3):1168–1180, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williamson:2022:MMR

[415] Brian D. Williamson, James P. Hughes, and Amy D. Willis. A multiview model for relative and absolute microbial abundances. Biometrics, 78(3): 1181–1194, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fisher:2022:PIS

[416] H. F. Fisher, R. J. Boys, C. S. Gillespie, C. J. Proctor, and A. Golightly. Parameter inference for a stochastic kinetic model of expanded polyglutamine proteins. Biometrics, 78(3):1195–1208, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chakraborty:2022:BAC

[417] Saptarshi Chakraborty, Tian Lan, Yiider Tseng, and Samuel W. K. Wong. Bayesian analysis of coupled cellular and nuclear trajectories for cell migration. Biometrics, 78(3):1209–1220, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2022:SCM

[418] Kevin He, Ji Zhu, Jian Kang, and Yi Li. Stratified Cox models with timevarying effects for national kidney transplant patients: a new blockwise steepest ascent method. Biometrics, 78(3):1221–1232, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2022:QDI

[419] Cheng Zheng and Lei Liu. Quantifying direct and indirect effect for longitudinal mediator and survival outcome using joint modeling approach. Biometrics, 78(3):1233–1243, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Golmakani:2022:NMC

[420] Marzieh K. Golmakani, Rebecca A. Hubbard, and Diana L. Miglioretti. Nonhomogeneous Markov chain for estimating the cumulative risk of multiple false positive screening tests. Biometrics, 78(3):1244–1256, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2022:SSF

[421] Lu Mao, KyungMann Kim, and Xinran Miao. Sample size formula for general win ratio analysis. Biometrics, 78(3):1257–1268, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2022:BRHb

[422] Li-Pang Chen. Book review: Handbook of measurement error models, Edited by Grace Y. Yi, Aurore Delaigle, and Paul Gustafson, Boca Raton, FL: Chapman and Hall/CRC, 2021. Biometrics, 78(3):1269–1270, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2022:BRS

[423] Han-Ming Wu. Book review: Supervised machine learning for text analysis in R, by Emil Hvitfeldt and Julia Silge. Biometrics, 78(3):1270–1272, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cui:2022:BRM

[424] Yuehua Cui. Book review: Multivariate data integration using R: Methods and applications with the mixOmics package. Kim-Anh Lê Cao and Zoe Marie Welham, New York: CRC Press, 2021. Biometrics, 78(3):1272– 1273, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2022:BRB

[425] Charlotte Wang. Book review: Behavior analysis with machine learning using R. Enrique Garcia Ceja. (2022). London; Boca Raton: CRC Press. DOI: 10.1201/978-1-003-20346-9. Biometrics, 78(3):1274, September 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2022:IId

[426] Anonymous. Issue information. Biometrics, 78(4):1275–1278, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dupont:2022:SNA

[427] Emiko Dupont, Simon N. Wood, and Nicole H. Augustin. Spatial+: a novel approach to spatial confounding. Biometrics, 78(4):1279–1290, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [428, 429, 430, 431] and rejoinder [432].

Reich:2022:DSN

[428] Brian J. Reich, Shu Yang, and Yawen Guan. Discussion on “Spatial+: a novel approach to spatial confounding” by Dupont, Wood, and Augustin. Biometrics, 78(4):1291–1294, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [427].

Marques:2022:DSN

[429] Isa Marques and Thomas Kneib. Discussion on “Spatial+: a novel approach to spatial confounding” by Emiko Dupont, Simon N. Wood, and Nicole H. Augustin. Biometrics, 78(4):1295–1299, December 2022. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [427].

Schmidt:2022:DSN

[430] Alexandra M. Schmidt. Discussion on “Spatial+: a novel approach to spatial confounding” by Emiko Dupont, Simon N. Wood, and Nicole H. Augustin. Biometrics, 78(4):1300–1304, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [427].

Papadogeorgou:2022:DSN

[431] Georgia Papadogeorgou. Discussion on “Spatial+: a novel approach to spatial confounding” by Emiko Dupont, Simon N. Wood, and Nicole H. Augustin. Biometrics, 78(4):1305–1308, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [427].

Dupont:2022:RDS

[432] Emiko Dupont, Simon N. Wood, and Nicole H. Augustin. Rejoinder to the discussions of “Spatial+: a novel approach to spatial confounding”. Biometrics, 78(4):1309–1312, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [427, 428, 429, 430, 431].

German:2022:WRS

[433] Christopher A. German, Janet S. Sinsheimer, Jin Zhou, and Hua Zhou. WiSER: Robust and scalable estimation and inference of within-subject variances from intensive longitudinal data. Biometrics, 78(4):1313–1327, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2022:VCR

[434] Dewei Wang, Xichen Mou, and Yan Liu. Varying-coefficient regression analysis for pooled biomonitoring. Biometrics, 78(4):1328–1341, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2022:ALS

[435] Heng Chen and Daniel F. Heitjan. Analysis of local sensitivity to nonignorability with missing outcomes and predictors. Biometrics, 78(4): 1342–1352, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2022:GST

[436] Jianghao Li and Sin-Ho Jung. Group sequential testing for cluster randomized trials with time-to-event endpoint. Biometrics, 78(4):1353–1364, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ventz:2022:ISD

[437] Steffen Ventz, Rahul Mazumder, and Lorenzo Trippa. Integration of survival data from multiple studies. Biometrics, 78(4):1365–1376, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2022:EOR

[438] Xin Chen, Rui Song, Jiajia Zhang, Swann Arp Adams, Liuquan Sun, and Wenbin Lu. On estimating optimal regime for treatment initiation time based on restricted mean residual lifetime. Biometrics, 78(4):1377–1389, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vakulenko-Lagun:2022:NES

[439] Bella Vakulenko-Lagun, Jing Qian, Sy Han Chiou, Nancy Wang, and Rebecca A. Betensky. Nonparametric estimation of the survival distribution under covariate-induced dependent truncation. Biometrics, 78(4): 1390–1401, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2022:SVS

[440] Liuquan Sun, Shuwei Li, Lianming Wang, Xinyuan Song, and Xuemei Sui. Simultaneous variable selection in regression analysis of multivariate interval-censored data. Biometrics, 78(4):1402–1413, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bussy:2022:BAC

[441] Simon Bussy, Mokhtar Z. Alaya, Anne-Sophie Jannot, and Agathe Guilloux. Binacox: automatic cut-point detection in high-dimensional Cox model with applications in genetics. Biometrics, 78(4):1414–1426, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scharf:2022:MBC

[442] Henry R. Scharf, Ann M. Raiho, Sierra Pugh, Carl A. Roland, David K. Swanson, Sarah E. Stehn, and Mevin B. Hooten. Multivariate Bayesian clustering using covariate-informed components with application to boreal vegetation sensitivity. Biometrics, 78(4):1427–1440, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2022:BAT

[443] Yusha Liu, John A. Kairalla, and Lindsay A. Renfro. Bayesian adaptive trial design for a continuous biomarker with possibly nonlinear or nonmonotone prognostic or predictive effects. Biometrics, 78(4):1441–1453, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fattorini:2022:DBP

[444] Lorenzo Fattorini, Marzia Marcheselli, Caterina Pisani, and Luca Pratelli. Design-based properties of the nearest neighbor spatial interpolator and its bootstrap mean squared error estimator. Biometrics, 78(4):1454–1463, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2022:SIR

[445] Chenguang Wang, Ao Yuan, Leslie Cope, and Jing Qin. A semiparametric isotonic regression model for skewed distributions with application to DNA–RNA–protein analysis. Biometrics, 78(4):1464–1474, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2022:PAV

[446] Baojiang Chen, Ao Yuan, and Jing Qin. Pool adjacent violators algorithmassisted learning with application on estimating optimal individualized treatment regimes. Biometrics, 78(4):1475–1488, December 2022. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Faraji:2022:ALR

[447] Nasrin Faraji, Mohammad Jafari Jozani, and Nader Nematollahi. Another look at regression analysis using ranked set samples with application to an osteoporosis study. Biometrics, 78(4):1489–1502, December 2022. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2022:EMP

[448] Hao Sun, Ashkan Ertefaie, and Brent A. Johnson. Estimating mean potential outcome under adaptive treatment length strategies in continuous time. Biometrics, 78(4):1503–1514, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2022:RCP

[449] Jiayin Zheng, Yingye Zheng, and Li Hsu. Re-calibrating pure risk integrating individual data from two-phase studies with external summary statistics. Biometrics, 78(4):1515–1529, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fintzi:2022:LNA

[450] Jonathan Fintzi, Jon Wakefield, and Vladimir N. Minin. A linear noise approximation for stochastic epidemic models fit to partially observed incidence counts. Biometrics, 78(4):1530–1541, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2022:MMM

[451] Yaqing Xu, Mengyun Wu, and Shuangge Ma. Multidimensional molecular measurements-environment interaction analysis for disease outcomes. Biometrics, 78(4):1542–1554, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2022:BSA

[452] Hao Sun, Emily Berg, and Zhengyuan Zhu. Bivariate small-area estimation for binary and Gaussian variables based on a conditionally specified model. Biometrics, 78(4):1555–1565, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2022:EBD

[453] Qiong Wu, Xiaoqi Huang, Adam J. Culbreth, James A. Waltz, L. Elliot Hong, and Shuo Chen. Extracting brain disease-related connectome subgraphs by adaptive dense subgraph discovery. Biometrics, 78(4):1566– 1578, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ren:2022:HCH

[454] Mingyang Ren, Qingzhao Zhang, Sanguo Zhang, Tingyan Zhong, Jian Huang, and Shuangge Ma. Hierarchical cancer heterogeneity analysis based on histopathological imaging features. Biometrics, 78(4):1579–1591, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Insolia:2022:SFS

[455] Luca Insolia, Ana Kenney, Francesca Chiaromonte, and Giovanni Felici. Simultaneous feature selection and outlier detection with optimality guarantees. Biometrics, 78(4):1592–1603, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Franks:2022:RSM

[456] Alexander M. Franks. Reducing subspace models for large-scale covariance regression. Biometrics, 78(4):1604–1613, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2022:JAC

[457] Yunfeng Zhang and Irina Gaynanova. Joint association and classification analysis of multi-view data. Biometrics, 78(4):1614–1625, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huling:2022:SDR

[458] Jared D. Huling and Menggang Yu. Sufficient dimension reduction for populations with structured heterogeneity. Biometrics, 78(4):1626–1638, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2022:BPB

[459] Bo Zhang, Siyu Heng, Emily J. MacKay, and Ting Ye. Bridging preference-based instrumental variable studies and cluster-randomized encouragement experiments: Study design, noncompliance, and average cluster effect ratio. Biometrics, 78(4):1639–1650, December 2022. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hashizume:2022:FUC

[460] Koichi Hashizume, Jun Tshuchida, and Takashi Sozu. Flexible use of copula-type model for dose-finding in drug combination clinical trials. Biometrics, 78(4):1651–1661, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dinart:2022:SSE

[461] Derek Dinart, Carine Bellera, and Virginie Rondeau. Sample size estimation for cancer randomized trials in the presence of heterogeneous populations. Biometrics, 78(4):1662–1673, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lotspeich:2022:EOR

[462] Sarah C. Lotspeich, Bryan E. Shepherd, Gustavo G. C. Amorim, Pamela A. Shaw, and Ran Tao. Efficient odds ratio estimation under two-phase sampling using error-prone data from a multi-national HIV research cohort. Biometrics, 78(4):1674–1685, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2022:JMZ

[463] Jiyuan Hu, Chan Wang, Martin J. Blaser, and Huilin Li. Joint modeling of zero-inflated longitudinal proportions and time-to-event data with application to a gut microbiome study. Biometrics, 78(4):1686–1698, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2022:WIR

[464] Sheng Wang and Hyunseung Kang. Weak-instrument robust tests in twosample summary-data Mendelian randomization. Biometrics, 78(4):1699– 1713, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tai:2022:BRF

[465] An-Shun Tai and Sheng-Hsuan Lin. Book review: Fundamentals of Causal Inference With R. Babette A. Brumback. 2021. New York : Chapman and Hall/CRC Press. 2021. Biometrics, 78(4):1714–1715, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Feng:2022:BRA

[466] Yen-Chen Anne Feng. Book review: Assessing COVID-19 and other pandemics and epidemics using computational modelling and data analysis, Subhendu Kumar Pani, Sujata Dash, Wellington P. dos Santos, Syed Ahmad Chan Bukhari, and Francesco Flammini, Switzerland: Springer Nature Switzerland AG. 2022. Biometrics, 78(4):1715–1716, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsu:2022:BRS

[467] Jason C. Hsu. Book review: Statistical issues in drug development, third edition. Stephen S. Senn New Jersey: John Wiley and Sons, Ltd., 2021. ISBN: 978-1-119-23857-7. Biometrics, 78(4):1716–1717, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2022:AR

[468] Anonymous. Acknowledgments referees 2022. Biometrics, 78(4):1721– 1724, December 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2023:IIa

[469] Anonymous. Issue information. Biometrics, 79(1):1–4, March 2023. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2023:RE

[470] Anonymous. Report of the Editors — 2022. Biometrics, 79(1):5–8, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stallard:2023:AED

[471] Nigel Stallard. Adaptive enrichment designs with a continuous biomarker. Biometrics, 79(1):9–19, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [472, 473, 474, 475] and rejoinder [476].

Phillips:2023:DAE

[472] Rachael V. Phillips and Mark J. van der Laan. Discussion on “Adaptive enrichment designs with a continuous biomarker” by Nigel Stallard. Biometrics, 79(1):20–22, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [471].

Wason:2023:DAE

[473] James M. S. Wason. Discussion on “Adaptive enrichment designs with a continuous biomarker” by Nigel Stallard. Biometrics, 79(1):23–25, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [471].

Jennison:2023:DAE

[474] Christopher Jennison. Discussion on “Adaptive enrichment designs with a continuous biomarker” by N. Stallard. Biometrics, 79(1):26–30, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [471].

Flournoy:2023:DAE

[475] Nancy Flournoy and Sergey Tarima. Discussion on “Adaptive enrichment designs with a continuous biomarker” by Nigel Stallard. Biometrics, 79 (1):31–35, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [471].

Stallard:2023:RDA

[476] Nigel Stallard. Rejoinder to discussion on “Adaptive enrichment designs with a continuous biomarker”. Biometrics, 79(1):36–38, March 2023. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [471, 472, 473, 474, 475].

Li:2023:CAC

[477] Ziyi Li, Yijian Huang, Dattatraya Patil, and Martin G. Sanda. Covariate adjustment in continuous biomarker assessment. Biometrics, 79(1):39– 48, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jiang:2023:EPD

[478] Liyun Jiang, Lei Nie, and Ying Yuan. Elastic priors to dynamically borrow information from historical data in clinical trials. Biometrics, 79(1):49– 60, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2023:RMT

[479] Lu Mao. On restricted mean time in favor of treatment. Biometrics, 79(1):61–72, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jiang:2023:DLS

[480] Jiakun Jiang, Wei Yang, Erin M. Schnellinger, Stephen E. Kimmel, and Wensheng Guo. Dynamic logistic state space prediction model for clinical decision making. Biometrics, 79(1):73–85, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barnett:2023:NST

[481] Helen Yvette Barnett, Sofía S. Villar, Helena Geys, and Thomas Jaki. A novel statistical test for treatment differences in clinical trials using a response-adaptive forward-looking Gittins Index Rule. Biometrics, 79(1): 86–97, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Davis-Plourde:2023:SSC

[482] Kendra Davis-Plourde, Monica Taljaard, and Fan Li. Sample size considerations for stepped wedge designs with subclusters. Biometrics, 79 (1):98–112, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2023:FAM

[483] Hyung Park, Eva Petkova, Thaddeus Tarpey, and R. Todd Ogden. Functional additive models for optimizing individualized treatment rules. Biometrics, 79(1):113–126, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Martinussen:2023:ESD

[484] Torben Martinussen and Mats Julius Stensrud. Estimation of separable direct and indirect effects in continuous time. Biometrics, 79(1):127–139, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gao:2023:NAR

[485] Fei Gao and K. C. G. Chan. Noniterative adjustment to regression estimators with population-based auxiliary information for semiparametric models. Biometrics, 79(1):140–150, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2023:BNQ

[486] Steven G. Xu and Brian J. Reich. Bayesian nonparametric quantile process regression and estimation of marginal quantile effects. Biometrics, 79(1): 151–164, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Seo:2023:AFT

[487] Byungtae Seo and Sangwook Kang. Accelerated failure time modeling via nonparametric mixtures. Biometrics, 79(1):165–177, March 2023. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:JMA

[488] Miaomiao Wang, Xinyu Zhang, Alan T. K. Wan, Kang You, and Guohua Zou. Jackknife model averaging for high-dimensional quantile regression. Biometrics, 79(1):178–189, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hou:2023:RPI

[489] Jue Hou, Stephanie F. Chan, Xuan Wang, and Tianxi Cai. Risk prediction with imperfect survival outcome information from electronic health records. Biometrics, 79(1):190–202, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ditzhaus:2023:CPI

[490] Marc Ditzhaus, Jon Genuneit, Arnold Janssen, and Markus Pauly. CASANOVA: Permutation inference in factorial survival designs. Biometrics, 79(1):203–215, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kapelner:2023:MPS

[491] Adam Kapelner and Abba Krieger. A matching procedure for sequential experiments that iteratively learns which covariates improve power. Biometrics, 79(1):216–229, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2023:SUF

[492] Shu Yang, Yilong Zhang, Guanghan Frank Liu, and Qian Guan. SMIM: a unified framework of survival sensitivity analysis using multiple imputation and martingale. Biometrics, 79(1):230–240, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kundu:2023:LRA

[493] Prosenjit Kundu and Nilanjan Chatterjee. Logistic regression analysis of two-phase studies using generalized method of moments. Biometrics, 79 (1):241–252, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2023:IVE

[494] Shuwei Li and Limin Peng. Instrumental variable estimation of complier causal treatment effect with interval-censored data. Biometrics, 79(1): 253–263, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2023:ISS

[495] Mengbing Li, Daniel E. Park, Maliha Aziz, Cindy M. Liu, Lance B. Price, and Zhenke Wu. Integrating sample similarities into latent class analysis: a tree-structured shrinkage approach. Biometrics, 79(1):264–279, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:LGD

[496] Wu Wang, Ying Sun, and Huixia Judy Wang. Latent group detection in functional partially linear regression models. Biometrics, 79(1):280– 291, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosal:2023:VSN

[497] Rahul Ghosal and Arnab Maity. Variable selection in nonlinear functionon-scalar regression. Biometrics, 79(1):292–303, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tuft:2023:SLR

[498] Marie Tuft, Martica H. Hall, and Robert T. Krafty. Spectra in low-rank localized layers (SpeLLL) for interpretable time-frequency analysis. Biometrics, 79(1):304–318, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Moss:2023:MPB

[499] Jonas Moss and Riccardo De Bin. Modelling publication bias and phacking. Biometrics, 79(1):319–331, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maronge:2023:GCC

[500] Jacob M. Maronge, Ran Tao, Jonathan S. Schildcrout, and Paul J. Rathouz. Generalized case-control sampling under generalized linear models. Biometrics, 79(1):332–343, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xia:2023:DLG

[501] Lu Xia, Bin Nan, and Yi Li. Debiased lasso for generalized linear models with a diverging number of covariates. Biometrics, 79(1):344–357, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lian:2023:APR

[502] Qinshu Lian, Jing Zhang, James S. Hodges, Yong Chen, and Haitao Chu. Accounting for post-randomization variables in meta-analysis: a joint meta-regression approach. Biometrics, 79(1):358–367, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2023:IMS

[503] Liwen Wu, Junyao Wang, and Abdus S. Wahed. Interim monitoring in sequential multiple assignment randomized trials. Biometrics, 79(1):368– 380, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Russo:2023:IRA

[504] Massimiliano Russo, Steffen Ventz, Victoria Wang, and Lorenzo Trippa. Inference in response-adaptive clinical trials when the enrolled population varies over time. Biometrics, 79(1):381–393, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2023:IEI

[505] Xinyu Li, Wang Miao, Fang Lu, and Xiao-Hua Zhou. Improving efficiency of inference in clinical trials with external control data. Biometrics, 79 (1):394–403, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mei:2023:HDC

[506] Hao Mei, Ruofan Jia, Guanzhong Qiao, Zhenqiu Lin, and Shuangge Ma. Human disease clinical treatment network for the elderly: analysis of the medicare inpatient length of stay and readmission data. Biometrics, 79 (1):404–416, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Verbeeck:2023:LMM

[507] Johan Verbeeck, Christel Faes, Thomas Neyens, Niel Hens, Geert Verbeke, Patrick Deboosere, and Geert Molenberghs. A linear mixed model to estimate COVID-19-induced excess mortality. Biometrics, 79(1):417– 425, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ward:2023:ILI

[508] Caitlin Ward, Grant D. Brown, and Jacob J. Oleson. An individual level infectious disease model in the presence of uncertainty from multiple, imperfect diagnostic tests. Biometrics, 79(1):426–436, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2023:SCS

[509] Xiao Song, Edward C. Chao, and Ching-Yun Wang. A smoothed corrected score approach for proportional hazards model with misclassified discretized covariates induced by error-contaminated continuous timedependent exposure. Biometrics, 79(1):437–448, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mork:2023:EPC

[510] Daniel Mork and Ander Wilson. Estimating perinatal critical windows of susceptibility to environmental mixtures via structured Bayesian regression tree pairs. Biometrics, 79(1):449–461, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McGee:2023:BMI

[511] Glen McGee, Ander Wilson, Thomas F. Webster, and Brent A. Coull. Bayesian multiple index models for environmental mixtures. Biometrics, 79(1):462–474, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:2023:SAI

[512] Paul R. Rosenbaum. Sensitivity analyses informed by tests for bias in observational studies. Biometrics, 79(1):475–487, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fei:2023:EAB

[513] Teng Fei, John Hanfelt, and Limin Peng. Evaluating the association between latent classes and competing risks outcomes with multiphenotype data. Biometrics, 79(1):488–501, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Luo:2023:CIO

[514] Shanshan Luo, Wei Li, and Yangbo He. Causal inference with outcomes truncated by death in multiarm studies. Biometrics, 79(1):502–513, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Balde:2023:RRO

[515] Ismaila Baldé, Yi Archer Yang, and Geneviève Lefebvre. Reader reaction to “Outcome-adaptive lasso: Variable selection for causal inference” by Shortreed and Ertefaie (2017). Biometrics, 79(1):514–520, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jones:2023:RRR

[516] Jeremiah Jones, Ashkan Ertefaie, and Susan M. Shortreed. Rejoinder to “Reader reaction to ‘Outcome-adaptive Lasso: Variable selection for causal inference’ by Shortreed and Ertefaie (2017)”. Biometrics, 79(1): 521–525, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chattopadhyay:2023:BRG

[517] Amrita Chattopadhyay. Book review: Gene expression data analysis: a statistical and machine learning perspective. Pankaj Barah, Dhruba Kumar Bhattacharyya, Jugal Kumar Kalita (2022). Boca Raton, Florida and London. CRC Press; Taylor and Francis. Biometrics, 79(1):526–528, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ting:2023:BRC

[518] Naitee Ting. Book review: Confidence intervals for discrete data in clinical research. By Vivek Pradhan, Ashis K. Gangopadhyay, Sandeep M. Menon, Cynthia Basu, Tathagata Banerjee, Chapman and Hall. 2021. pp.

226. $119.95. (hbk). ISBN: 978-1-138-04898-0. Biometrics, 79(1):528–531, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chang:2023:BRE

[519] Hung-Ching Chang and Michael T. Gorczyca. Book review: The effect: an introduction to research design and causality. By Nick Huntington-Klein (2022). New York. Chapman and Hall. Biometrics, 79(1):531–532, March 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2023:IIb

[520] Anonymous. Issue information. Biometrics, 79(2):533–538, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:IVE

[521] Linbo Wang, Eric Tchetgen Tchetgen, Torben Martinussen, and Stijn Vansteelandt. Instrumental variable estimation of the causal hazard ratio. Biometrics, 79(2):539–550, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [522, 523, 524, 525].

Frandsen:2023:DIV

[522] Brigham Russell Frandsen. Discussion on “Instrumental variable estimation of the causal hazard ratio,” by Linbo Wang, Eric Tchetgen Tchetgen, Torben Martinussen, Stijn Vansteelandt. Biometrics, 79(2):551–553, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [521].

Baer:2023:DIV

[523] Benjamin R. Baer, Robert L. Strawderman, and Ashkan Ertefaie. Discussion on “Instrumental variable estimation of the causal hazard ratio,” by Linbo Wang, Eric Tchetgen Tchetgen, Torben Martinussen, and Stijn Vansteelandt. Biometrics, 79(2):554–558, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [521].

OMalley:2023:DIV

[524] A. James O’Malley, Pablo Martínez-Camblor, and Todd A. MacKenzie. Discussion on “Instrumental variable estimation of the causal hazard ratio” by Linbo Wang, Eric Tchetgen Tchetgen, Torben Martinussen, and Stijn Vansteelandt. Biometrics, 79(2):559–563, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [521].

Wang:2023:RDI

[525] Linbo Wang, Eric Tchetgen Tchetgen, Torben Martinussen, and Stijn Vansteelandt. Rejoinder to discussions on “Instrumental variable estimation of the causal hazard ratio”. Biometrics, 79(2):564–568, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [522, 523, 524, 521].

Ye:2023:IDD

[526] Ting Ye, Ashkan Ertefaie, James Flory, Sean Hennessy, and Dylan S. Small. Instrumented difference-in-differences. Biometrics, 79(2):569–581, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [527, 528, 529, 530] and rejoinder [531].

Beyhum:2023:DID

[527] Jad Beyhum, Jean-Pierre Florens, and Ingrid Van Keilegom. Discussion on “Instrumented difference-in-differences” by Ting Ye, Ashkan Ertefaie, James Flory, Sean Hennessy, and Dylan S. Small. Biometrics, 79(2):582– 586, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [526].

Tan:2023:DID

[528] Zhiqiang Tan. Discussion on “Instrumented difference-in-differences” by Ye, Ertefaie, Flory, Hennessy, Small. Biometrics, 79(2):587–591, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [526].

Kang:2023:DID

[529] Hyunseung Kang. Discussion on “Instrumented difference-in-differences” by Ting Ye, Ashkan Ertefaie, James Flory, Sean Hennessy & Dylan S. Small. Biometrics, 79(2):592–596, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [526].

DiazOrdaz:2023:DID

[530] Karla DiazOrdaz. Discussion on: Instrumented difference-in-differences, by Ting Ye, Ashkan Ertefaie, James Flory, Sean Hennessy and Dylan S. Small. Biometrics, 79(2):597–600, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [526].

Ye:2023:RID

[531] Ting Ye, Ashkan Ertefaie, James Flory, Sean Hennessy, and Dylan S. Small. Rejoinder to “Instrumented difference-in-differences”. Biometrics, 79(2):601–603, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [527, 528, 529, 530, 526].

Masotti:2023:NBF

[532] Maria Masotti, Lin Zhang, Ethan Leng, Gregory J. Metzger, and Joseph S. Koopmeiners. A novel Bayesian functional spatial partitioning method with application to prostate cancer lesion detection using MRI. Biometrics, 79(2):604–615, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2023:BSM

[533] Cheng-Han Yu, Raquel Prado, Hernando Ombao, and Daniel Rowe. Bayesian spatiotemporal modeling on complex-valued fMRI signals via kernel convolutions. Biometrics, 79(2):616–628, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2023:BIS

[534] Cheng-Han Yu, Meng Li, Colin Noe, Simon Fischer-Baum, and Marina Vannucci. Bayesian inference for stationary points in Gaussian process regression models for event-related potentials analysis. Biometrics, 79(2): 629–641, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2023:BOI

[535] Peng Yu, Spencer Ericksen, Anthony Gitter, and Michael A. Newton. Bayes optimal informer sets for early-stage drug discovery. Biometrics, 79(2):642–654, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2023:BIS

[536] Yize Zhao, Ben Wu, and Jian Kang. Bayesian interaction selection model for multimodal neuroimaging data analysis. Biometrics, 79(2):655–668, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2023:BSS

[537] Haiyan Zheng, Thomas Jaki, and James M. S. Wason. Bayesian sample size determination using commensurate priors to leverage preexperimental data. Biometrics, 79(2):669–683, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ren:2023:RBV

[538] Jie Ren, Fei Zhou, Xiaoxi Li, Shuangge Ma, Yu Jiang, and Cen Wu. Robust Bayesian variable selection for gene-environment interactions. Biometrics, 79(2):684–694, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2023:SAT

[539] Yanqing Sun, Qiong Shou, Peter B. Gilbert, Fei Heng, and Xiyuan Qian. Semiparametric additive time-varying coefficients model for longitudinal data with censored time origin. Biometrics, 79(2):695–710, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mandel:2023:NNC

[540] Francesca Mandel, Riddhi Pratim Ghosh, and Ian Barnett. Neural networks for clustered and longitudinal data using mixed effects models. Biometrics, 79(2):711–721, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Weaver:2023:FDA

[541] Caleb Weaver, Luo Xiao, and Wenbin Lu. Functional data analysis for longitudinal data with informative observation times. Biometrics, 79(2): 722–733, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hoque:2023:THD

[542] Md Erfanul Hoque, Elif F. Acar, and Mahmoud Torabi. A timeheterogeneous D-vine copula model for unbalanced and unequally spaced longitudinal data. Biometrics, 79(2):734–746, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wan:2023:MQR

[543] Chuang Wan, Wei Zhong, Wenyang Zhang, and Changliang Zou. Multikink quantile regression for longitudinal data with application to progesterone data analysis. Biometrics, 79(2):747–760, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2023:MBC

[544] Luoying Yang and Tong Tong Wu. Model-based clustering of highdimensional longitudinal data via regularization. Biometrics, 79(2):761– 774, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:CML

[545] Linbo Wang, Xiang Meng, Thomas S. Richardson, and James M. Robins. Coherent modeling of longitudinal causal effects on binary outcomes. Biometrics, 79(2):775–787, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:RAC

[546] Xuan Wang, Layla Parast, Larry Han, Lu Tian, and Tianxi Cai. Robust approach to combining multiple markers to improve surrogacy. Biometrics, 79(2):788–798, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Parast:2023:THU

[547] Layla Parast, Tianxi Cai, and Lu Tian. Testing for heterogeneity in the utility of a surrogate marker. Biometrics, 79(2):799–810, June 2023. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Feng:2023:SPS

[548] Jean Feng, Arjun Sondhi, Jessica Perry, and Noah Simon. Selective prediction-set models with coverage rate guarantees. Biometrics, 79(2): 811–825, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Do:2023:JFM

[549] Hyungrok Do, Shinjini Nandi, Preston Putzel, Padhraic Smyth, and Judy Zhong. A joint fairness model with applications to risk predictions for underrepresented populations. Biometrics, 79(2):826–840, June 2023. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2023:CTP

[550] Bingxin Zhao, Fei Zou, and Hongtu Zhu. Cross-trait prediction accuracy of summary statistics in genome-wide association studies. Biometrics, 79(2):841–853, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heiling:2023:ECT

[551] Hillary M. Heiling, Douglas R. Wilson, Naim U. Rashid, Wei Sun, and Joseph G. Ibrahim. Estimating cell type composition using isoform expression one gene at a time. Biometrics, 79(2):854–865, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2023:MSC

[552] Lin Lin, Wei Shi, Jianbo Ye, and Jia Li. Multisource single-cell data integration by MAW barycenter for Gaussian mixture models. Biometrics, 79(2):866–877, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2023:FSL

[553] Congran Yu, Wenwen Guo, Xinyuan Song, and Hengjian Cui. Feature screening with latent responses. Biometrics, 79(2):878–890, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2023:ERA

[554] Yuyang Xu, Zhonghua Liu, and Jianfeng Yao. An eigenvalue ratio approach to inferring population structure from whole genome sequencing data. Biometrics, 79(2):891–902, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2023:UHD

[555] Dingke Tang, Dehan Kong, Wenliang Pan, and Linbo Wang. Ultra-high dimensional variable selection for doubly robust causal inference. Biometrics, 79(2):903–914, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dong:2023:JGN

[556] Meichen Dong, Yiping He, Yuchao Jiang, and Fei Zou. Joint gene network construction by single-cell RNA sequencing data. Biometrics, 79(2):915– 925, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nghiem:2023:SML

[557] Linh H. Nghiem, Francis K. C. Hui, Samuel Müller, and A. H. Welsh. Screening methods for linear errors-in-variables models in high dimensions. Biometrics, 79(2):926–939, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2023:CHD

[558] Tianqi Liu, Yu Lu, Biqing Zhu, and Hongyu Zhao. Clustering highdimensional data via feature selection. Biometrics, 79(2):940–950, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2023:GFN

[559] Hang Yu, Yuanjia Wang, and Donglin Zeng. A general framework of nonparametric feature selection in high-dimensional data. Biometrics, 79 (2):951–963, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2023:RPE

[560] Fuli Zhang and Kung-Sik Chan. Random projection ensemble classification with high-dimensional time series. Biometrics, 79(2):964–974, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsiatis:2023:EOR

[561] Anastasios A. Tsiatis, Marie Davidian, and Shannon T. Holloway. Estimation of the odds ratio in a proportional odds model with censored time-lagged outcome in a randomized clinical trial. Biometrics, 79(2): 975–987, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bian:2023:VSR

[562] Zeyu Bian, Erica E. M. Moodie, Susan M. Shortreed, and Sahir Bhatnagar. Variable selection in regression-based estimation of dynamic treatment regimes. Biometrics, 79(2):988–999, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Betensky:2023:NSE

[563] Rebecca A. Betensky, Jing Qian, and Jingyao Hou. Nonparametric and semiparametric estimation with sequentially truncated survival data. Biometrics, 79(2):1000–1013, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

vanderLaan:2023:NEC

[564] Lars van der Laan, Wenbo Zhang, and Peter B. Gilbert. Nonparametric estimation of the causal effect of a stochastic threshold-based intervention. Biometrics, 79(2):1014–1028, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ertefaie:2023:NIP

[565] Ashkan Ertefaie, Nima S. Hejazi, and Mark J. van der Laan. Nonparametric inverse-probability-weighted estimators based on the highly adaptive lasso. Biometrics, 79(2):1029–1041, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hong:2023:PCC

[566] Guanglei Hong, Fan Yang, and Xu Qin. Posttreatment confounding in causal mediation studies: a cutting-edge problem and a novel solution via sensitivity analysis. Biometrics, 79(2):1042–1056, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dahabreh:2023:ERM

[567] Issa J. Dahabreh, Sarah E. Robertson, Lucia C. Petito, Miguel A. Hernán, and Jon A. Steingrimsson. Efficient and robust methods for causally interpretable meta-analysis: Transporting inferences from multiple randomized trials to a target population. Biometrics, 79(2):1057–1072, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2023:GNS

[568] Qihuang Zhang and Grace Y. Yi. Generalized network structured models with mixed responses subject to measurement error and misclassification. Biometrics, 79(2):1073–1088, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2023:ZIP

[569] Qihuang Zhang and Grace Y. Yi. Zero-inflated Poisson models with measurement error in the response. Biometrics, 79(2):1089–1102, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2023:CTG

[570] Ningning Xu, Aldo Solari, and Jelle J. Goeman. Closed testing with Globaltest, with application in metabolomics. Biometrics, 79(2):1103– 1113, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Proschan:2023:NFE

[571] Michael A. Proschan and Dean A. Follmann. A note on familywise error rate for a primary and secondary endpoint. Biometrics, 79(2):1114–1118, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Abramowicz:2023:DSF

[572] Konrad Abramowicz, Alessia Pini, Lina Schelin, Sara Sjöstedt de Luna, Aymeric Stamm, and Simone Vantini. Domain selection and familywise error rate for functional data: a unified framework. Biometrics, 79(2): 1119–1132, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hawila:2023:ECC

[573] Nour Hawila and Arthur Berg. Exact-corrected confidence interval for risk difference in noninferiority binomial trials. Biometrics, 79(2):1133– 1144, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yan:2023:EQI

[574] Feifei Yan, Yanyan Liu, Jianwen Cai, and Haibo Zhou. Estimated quadratic inference function for correlated failure time data. Biometrics, 79(2):1145–1158, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2023:GFC

[575] Hong Zhang and Zheyang Wu. The generalized Fisher’s combination and accurate p-value calculation under dependence. Biometrics, 79(2):1159– 1172, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2023:INP

[576] Haoyan Hu and Yumou Qiu. Inference for nonparanormal partial correlation via regularized rank-based nodewise regression. Biometrics, 79 (2):1173–1186, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2023:DVM

[577] Yutong Liu, Toni Darville, Xiaojing Zheng, and Quefeng Li. Decomposition of variation of mixed variables by a latent mixed Gaussian copula model. Biometrics, 79(2):1187–1200, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xin:2023:CDA

[578] Huiqin Xin and Sihai Dave Zhao. A compound decision approach to covariance matrix estimation. Biometrics, 79(2):1201–1212, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2023:ITG

[579] Dasom Lee, Shu Yang, Lin Dong, Xiaofei Wang, Donglin Zeng, and Jianwen Cai. Improving trial generalizability using observational studies. Biometrics, 79(2):1213–1225, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:FGB

[580] Zhengjia Wang, John Magnotti, Michael S. Beauchamp, and Meng Li. Functional group bridge for simultaneous regression and support estimation. Biometrics, 79(2):1226–1238, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:RFP

[581] Guangxing Wang, Sisheng Liu, Fang Han, and Chong-Zhi Di. Robust functional principal component analysis via a functional pairwise spatial sign operator. Biometrics, 79(2):1239–1253, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Altieri:2023:CTI

[582] Linda Altieri, Alessio Farcomeni, and Danilo Alunni Fegatelli. Continuous time-interaction processes for population size estimation, with an application to drug dealing in Italy. Biometrics, 79(2):1254–1267, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:STM

[583] Hairu Wang, Zhiping Lu, and Yukun Liu. Score test for missing at random or not under logistic missingness models. Biometrics, 79(2):1268– 1279, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tran:2023:CVS

[584] Lam Tran, Kevin He, Di Wang, and Hui Jiang. A cross-validation statistical framework for asymmetric data integration. Biometrics, 79(2): 1280–1292, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2023:PAC

[585] Siyun Yang, Mirjam Moerbeek, Monica Taljaard, and Fan Li. Power analysis for cluster randomized trials with continuous coprimary endpoints. Biometrics, 79(2):1293–1305, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2023:TDH

[586] Anthony Cheng, Disheng Mao, Yuping Zhang, Joseph Glaz, and Zhengqing Ouyang. Translocation detection from Hi-C data via scan statistics. Biometrics, 79(2):1306–1317, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2023:ARR

[587] Gen Li, Yan Li, and Kun Chen. It’s all relative: Regression analysis with compositional predictors. Biometrics, 79(2):1318–1329, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shahn:2023:FCI

[588] Zach Shahn, Miguel A. Hernán, and James M. Robins. A formal causal interpretation of the case-crossover design. Biometrics, 79(2):1330–1343, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [589, 590, 591] and rejoinder [592].

Andersen:2023:DFC

[589] Per Kragh Andersen and Torben Martinussen. Discussion of “A formal causal interpretation of the case-crossover design”. Biometrics, 79(2): 1344–1345, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [588].

Pfeiffer:2023:DFC

[590] Ruth M. Pfeiffer and Mitchell H. Gail. Discussion of “A formal causal interpretation of the case-crossover design” by Zach Shahn, Miguel A. Hernan, and James M. Robins. Biometrics, 79(2):1346–1348, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [588].

Lumley:2023:DFC

[591] Thomas Lumley. Discussion on “A formal causal interpretation of the case-crossover design” by Zach Shahn, Miguel A. Hernán, and James M. Robins. Biometrics, 79(2):1349–1350, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [588].

Shahn:2023:RFC

[592] Zach Shahn, Miguel A. Hernán, and James M. Robins. Rejoinder: A formal causal interpretation of the case-crossover design. Biometrics, 79 (2):1351–1358, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [588, 589, 590, 591].

Jiang:2023:STD

[593] Shu Jiang, Jiguo Cao, Bernard Rosner, and Graham A. Colditz. Supervised two-dimensional functional principal component analysis with time-to-event outcomes and mammogram imaging data. Biometrics, 79 (2):1359–1369, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DAngelo:2023:BNA

[594] Laura D’Angelo, Antonio Canale, Zhaoxia Yu, and Michele Guindani. Bayesian nonparametric analysis for the detection of spikes in noisy calcium imaging data. Biometrics, 79(2):1370–1382, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2023:BNA

[595] Chenyang Zhang and Guosheng Yin. Bayesian nonparametric analysis of restricted mean survival time. Biometrics, 79(2):1383–1396, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Parker:2023:BFD

[596] Paul A. Parker and Scott H. Holan. A Bayesian functional data model for surveys collected under informative sampling with application to mortality estimation using NHANES. Biometrics, 79(2):1397–1408, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alexandria:2023:AIE

[597] Shaina J. Alexandria, Michael G. Hudgens, and Allison E. Aiello. Assessing intervention effects in a randomized trial within a social network. Biometrics, 79(2):1409–1419, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DiGravio:2023:DAT

[598] Chiara Di Gravio, Ran Tao, and Jonathan S. Schildcrout. Design and analysis of two-phase studies with multivariate longitudinal data. Biometrics, 79(2):1420–1432, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Proper:2023:AME

[599] Jennifer Proper and Thomas A. Murray. An alternative metric for evaluating the potential patient benefit of response-adaptive randomization procedures. Biometrics, 79(2):1433–1445, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:BMA

[600] Chenguang Wang, Min Lin, Gary L. Rosner, and Guoxing Soon. A Bayesian model with application for adaptive platform trials having temporal changes. Biometrics, 79(2):1446–1458, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2023:BPT

[601] Yujie Zhao, Rui (Sammi) Tang, Yeting Du, and Ying Yuan. A Bayesian platform trial design to simultaneously evaluate multiple drugs in multiple indications with mixed endpoints. Biometrics, 79(2):1459–1471, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McCaw:2023:LSO

[602] Zachary R. McCaw, Sheila M. Gaynor, Ryan Sun, and Xihong Lin. Leveraging a surrogate outcome to improve inference on a partially missing target outcome. Biometrics, 79(2):1472–1484, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sloan:2023:RMA

[603] Abigail Sloan, Chao Cheng, Bernard Rosner, Regina G. Ziegler, Stephanie A. Smith-Warner, and Molin Wang. A repeated measures approach to pooled and calibrated biomarker data. Biometrics, 79(2): 1485–1495, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jeffries:2023:ETE

[604] Neal O. Jeffries, James F. Troendle, and Nancy L. Geller. Evaluating treatment effects in group sequential multivariate longitudinal studies with covariate adjustment. Biometrics, 79(2):1496–1506, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2023:HMA

[605] Yuzi Zhang, Howard H. Chang, Qu Cheng, Philip A. Collender, Ting Li, Jinge He, and Justin V. Remais. A hierarchical model for analyzing multisite individual-level disease surveillance data from multiple systems. Biometrics, 79(2):1507–1519, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kowal:2023:SCD

[606] Daniel R. Kowal and Bohan Wu. Semiparametric count data regression for self-reported mental health. Biometrics, 79(2):1520–1533, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Etievant:2023:IER

[607] Lola Etievant, Joshua N. Sampson, and Mitchell H. Gail. Increasing efficiency and reducing bias when assessing HPV vaccination efficacy by using nontargeted HPV strains. Biometrics, 79(2):1534–1545, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Drew:2023:SED

[608] Clara Drew, Moses Badio, Dehkontee Dennis, Lisa Hensley, Elizabeth Higgs, Michael Sneller, Mosoka Fallah, and Cavan Reilly. Simplifying the estimation of diagnostic testing accuracy over time for high specificity tests in the absence of a gold standard. Biometrics, 79(2):1546–1558, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ma:2023:FCM

[609] Zichen Ma, Shannon W. Davis, and Yen-Yi Ho. Flexible copula model for integrating correlated multi-omics data from single-cell experiments. Biometrics, 79(2):1559–1572, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2023:ISB

[610] Ryan Sun, Liang Zhu, Yimei Li, Yutaka Yasui, and Leslie Robison. Inference for set-based effects in genetic association studies with intervalcensored outcomes. Biometrics, 79(2):1573–1585, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Banks:2023:BRS

[611] David Banks. Book review: Statistics in the public interest: In memory of Stephen E. Fienberg. Alicia L. Carriquiry, Judith M. Tanur, William F. Eddy, Margaret L. Smykla (Eds.), New York City : Springer. 2022. Biometrics, 79(2):1586–1587, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hodges:2023:BRW

[612] James S. Hodges. Book review: Writing grant proposals in epidemiology, preventive medicine, and biostatistics. Lisa Chasan-Taber, CRC Press: Boca Raton FL. 2022. Biometrics, 79(2):1587–1589, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsiao:2023:BRP

[613] Chuhsing Kate Hsiao. Book review: Principles of biostatistics (3rd ed), Marcello Pagano, Kimberlee Gauvreau, Heather Mattie (2022). Boca Raton, FL : CRC Press. Biometrics, 79(2):1589–1590, June 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2023:IIc

[614] Anonymous. Issue information. Biometrics, 79(3):1591–1596, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ying:2023:SCS

[615] Andrew Ying and Eric J. Tchetgen Tchetgen. Structural cumulative survival models for estimation of treatment effects accounting for treatment switching in randomized experiments. Biometrics, 79(3):1597–1609, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Molstad:2023:DRI

[616] Aaron J. Molstad and Rohit K. Patra. Dimension reduction for integrative survival analysis. Biometrics, 79(3):1610–1623, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hartman:2023:CIL

[617] Nicholas Hartman, Sehee Kim, Kevin He, and John D. Kalbfleisch. Concordance indices with left-truncated and right-censored data. Biometrics, 79(3):1624–1634, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wen:2023:JIC

[618] Jiyang Wen, Chen Hu, and Mei-Cheng Wang. Joint inference for competing risks data using multiple endpoints. Biometrics, 79(3):1635–1645, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2023:MLE

[619] Chengyuan Lu, Jelle Goeman, and Hein Putter. Maximum likelihood estimation in the additive hazards model. Biometrics, 79(3):1646–1656, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reeder:2023:PEF

[620] Harrison T. Reeder, Junwei Lu, and Sebastien Haneuse. Penalized estimation of frailty-based illness-death models for semi-competing risks. Biometrics, 79(3):1657–1669, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2023:MPH

[621] Kaitlyn Cook, Wenbin Lu, and Rui Wang. Marginal proportional hazards models for clustered interval-censored data with time-dependent covariates. Biometrics, 79(3):1670–1685, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2023:ISE

[622] Ming-Yueh Huang and Chiung-Yu Huang. Improved semiparametric estimation of the proportional rate model with recurrent event data. Biometrics, 79(3):1686–1700, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mandal:2023:NPE

[623] S. Mandal, J. Qin, and R. M. Pfeiffer. Non-parametric estimation of the age-at-onset distribution from a cross-sectional sample. Biometrics, 79(3): 1701–1712, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2023:IRB

[624] Tao Sun, Yu Cheng, and Ying Ding. An information ratio-based goodnessof-fit test for copula models on censored data. Biometrics, 79(3):1713– 1725, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Garcia-Donato:2023:MUQ

[625] Gonzalo García-Donato, Stefano Cabras, and María Eugenia Castellanos. Model uncertainty quantification in Cox regression. Biometrics, 79(3): 1726–1736, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ruhl:2023:GIC

[626] Jasmin Rühl, Jan Beyersmann, and Sarah Friedrich. General independent censoring in event-driven trials with staggered entry. Biometrics, 79(3): 1737–1748, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2023:NIG

[627] Lu Mao. Nonparametric inference of general while-alive estimands for recurrent events. Biometrics, 79(3):1749–1760, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qin:2023:TLB

[628] Xing Qin, Shuangge Ma, and Mengyun Wu. Two-level Bayesian interaction analysis for survival data incorporating pathway information. Biometrics, 79(3):1761–1774, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Allen:2023:BMM

[629] Carter Allen, Yuzhou Chang, Brian Neelon, Won Chang, Hang J. Kim, Zihai Li, Qin Ma, and Dongjun Chung. A Bayesian multivariate mixture model for high throughput spatial transcriptomics. Biometrics, 79(3): 1775–1787, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2023:TBS

[630] Zhongwei Zhang, Reinaldo B. Arellano-Valle, Marc G. Genton, and Raphaël Huser. Tractable Bayes of skew-elliptical link models for correlated binary data. Biometrics, 79(3):1788–1800, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Panigrahi:2023:IBM

[631] Snigdha Panigrahi, Shariq Mohammed, Arvind Rao, and Veerabhadran Baladandayuthapani. Integrative Bayesian models using post-selective inference: a case study in radiogenomics. Biometrics, 79(3):1801–1813, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2023:BRA

[632] Inkoo Lee, Debajyoti Sinha, Qing Mai, Xin Zhang, and Dipankar Bandyopadhyay. Bayesian regression analysis of skewed tensor responses. Biometrics, 79(3):1814–1825, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:ABS

[633] Yakun Wang, Zeda Li, and Scott A. Bruce. Adaptive Bayesian sum of trees model for covariate-dependent spectral analysis. Biometrics, 79(3): 1826–1839, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roberts:2023:SSV

[634] Emily K. Roberts, Michael R. Elliott, and Jeremy M. G. Taylor. Solutions for surrogacy validation with longitudinal outcomes for a gene therapy. Biometrics, 79(3):1840–1852, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kowal:2023:SSL

[635] Daniel R. Kowal. Subset selection for linear mixed models. Biometrics, 79(3):1853–1867, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ito:2023:GGE

[636] Tsubasa Ito and Shonosuke Sugasawa. Grouped generalized estimating equations for longitudinal data analysis. Biometrics, 79(3):1868–1879, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2023:AFL

[637] Ting Li, Huichen Zhu, Tengfei Li, and Hongtu Zhu. Asynchronous functional linear regression models for longitudinal data in reproducing kernel Hilbert space. Biometrics, 79(3):1880–1895, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matthews:2023:CPS

[638] John N. S. Matthews, Sofia Bazakou, Robin Henderson, and Linda D. Sharples. Contrasting principal stratum and hypothetical strategy estimands in multi-period crossover trials with incomplete data. Biometrics, 79(3):1896–1907, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heller:2023:OMT

[639] Ruth Heller, Abba Krieger, and Saharon Rosset. Optimal multiple testing and design in clinical trials. Biometrics, 79(3):1908–1919, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jin:2023:CPA

[640] Peng Jin, Wenbin Lu, Yu Chen, and Mengling Liu. Change-plane analysis for subgroup detection with a continuous treatment. Biometrics, 79(3): 1920–1933, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wei:2023:ETL

[641] Waverly Wei, Maya Petersen, Mark J. van der Laan, Zeyu Zheng, Chong Wu, and Jingshen Wang. Efficient targeted learning of heterogeneous treatment effects for multiple subgroups. Biometrics, 79(3):1934–1946, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wei:2023:TRQ

[642] Bo Wei, Limin Peng, Ying Guo, Amita Manatunga, and Jennifer Stevens. Tensor response quantile regression with neuroimaging data. Biometrics, 79(3):1947–1958, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhong:2023:JSM

[643] Weibin Zhong and Guoqing Diao. Joint semiparametric models for casecohort designs. Biometrics, 79(3):1959–1971, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rodriguez-Alvarez:2023:MAS

[644] María Xosé Rodríguez-Álvarez, María Durbán, Paul H. C. Eilers, Dae-Jin Lee, and Francisco Gonzalez. Multidimensional adaptive P -splines with application to neurons’ activity studies. Biometrics, 79(3):1972–1985, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Daly-Grafstein:2023:CPN

[645] Daniel Daly-Grafstein and Paul Gustafson. Combining parametric and nonparametric models to estimate treatment effects in observational studies. Biometrics, 79(3):1986–1995, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2023:SET

[646] Yu-Jen Cheng, Yen-Chun Liu, Chang-Yu Tsai, and Chiung-Yu Huang. Semiparametric estimation of the transformation model by leveraging external aggregate data in the presence of population heterogeneity. Biometrics, 79(3):1996–2009, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2023:SJM

[647] Chun Yin Lee, Kin Yau Wong, Kwok Fai Lam, and Dipankar Bandyopadhyay. A semiparametric joint model for cluster size and subunitspecific interval-censored outcomes. Biometrics, 79(3):2010–2022, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dai:2023:RAE

[648] Guorong Dai, Yanyuan Ma, Jill Hasler, Jinbo Chen, and Raymond J. Carroll. A robust approach for electronic health record-based casecontrol studies with contaminated case pools. Biometrics, 79(3):2023– 2035, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2023:QRN

[649] Aiai Yu, Yujie Zhong, Xingdong Feng, and Ying Wei. Quantile regression for nonignorable missing data with its application of analyzing electronic medical records. Biometrics, 79(3):2036–2049, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2023:FMA

[650] Huihang Liu and Xinyu Zhang. Frequentist model averaging for undirected Gaussian graphical models. Biometrics, 79(3):2050–2062, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Todem:2023:NST

[651] David Todem, Wei-Wen Hsu, and KyungMann Kim. Nonparametric scanning tests of homogeneity for hierarchical models with continuous covariates. Biometrics, 79(3):2063–2075, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mollenhoff:2023:IAC

[652] Kathrin Möllenhoff, Kirsten Schorning, and Franziska Kappenberg. Identifying alert concentrations using a model-based bootstrap approach. Biometrics, 79(3):2076–2088, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2023:APB

[653] Ao Huang, Kosuke Morikawa, Tim Friede, and Satoshi Hattori. Adjusting for publication bias in meta-analysis via inverse probability weighting using clinical trial registries. Biometrics, 79(3):2089–2102, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steyer:2023:EAI

[654] Lisa Steyer, Almond Stöcker, and Sonja Greven. Elastic analysis of irregularly or sparsely sampled curves. Biometrics, 79(3):2103–2115, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2023:GFS

[655] Dana Johnson, Wenbin Lu, and Marie Davidian. A general framework for subgroup detection via one-step value difference estimation. Biometrics, 79(3):2116–2126, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2023:PSR

[656] Ke Zhu and Hanzhong Liu. Pair-switching rerandomization. Biometrics, 79(3):2127–2142, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gerard:2023:DRE

[657] David Gerard. Double reduction estimation and equilibrium tests in natural autopolyploid populations. Biometrics, 79(3):2143–2156, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2023:CAT

[658] Ye He, Ling Zhou, Yingcun Xia, and Huazhen Lin. Center-augmented l_{2}- type regularization for subgroup learning. Biometrics, 79(3):2157–2170, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diana:2023:GMF

[659] Alex Diana, Eleni Matechou, Jim Griffin, Todd Arnold, Simone Tenan, and Stefano Volponi. A general modeling framework for open wildlife populations based on the Polya tree prior. Biometrics, 79(3):2171–2183, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2023:NPI

[660] Siqi Xu, Peng Wang, Wing Kam Fung, and Zhonghua Liu. A novel penalized inverse-variance weighted estimator for Mendelian randomization with applications to COVID-19 outcomes. Biometrics, 79(3):2184–2195, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fogarty:2023:TWN

[661] Colin B. Fogarty. Testing weak nulls in matched observational studies. Biometrics, 79(3):2196–2207, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2023:MRM

[662] Zhonghua Liu, Ting Ye, Baoluo Sun, Mary Schooling, and Eric Tchetgen Tchetgen. Mendelian randomization mixed-scale treatment effect robust identification and estimation for causal inference. Biometrics, 79(3):2208– 2219, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Giffin:2023:GPS

[663] A. Giffin, B. J. Reich, S. Yang, and A. G. Rappold. Generalized propensity score approach to causal inference with spatial interference. Biometrics, 79(3):2220–2231, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2023:FDA

[664] Chenlin Zhang, Huazhen Lin, Li Liu, Jin Liu, and Yi Li. Functional data analysis with covariate-dependent mean and covariance structures. Biometrics, 79(3):2232–2245, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2023:SCS

[665] Dong Liu, Changwei Zhao, Yong He, Lei Liu, Ying Guo, and Xinsheng Zhang. Simultaneous cluster structure learning and estimation of heterogeneous graphs for matrix-variate fMRI data. Biometrics, 79(3):2246– 2259, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2023:ETB

[666] Nina Zhou, Lu Wang, and Daniel Almirall. Estimating tree-based dynamic treatment regimes using observational data with restricted treatment sequences. Biometrics, 79(3):2260–2271, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2023:SCC

[667] Feipeng Zhang and Qunhua Li. Segmented correspondence curve regression for quantifying covariate effects on the reproducibility of highthroughput experiments. Biometrics, 79(3):2272–2285, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ekvall:2023:CLB

[668] K. O. Ekvall and M. Bottai. Concave likelihood-based regression with finite-support response variables. Biometrics, 79(3):2286–2297, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hans:2023:BDC

[669] Nicolai Hans, Nadja Klein, Florian Faschingbauer, Michael Schneider, and Andreas Mayr. Boosting distributional copula regression. Biometrics, 79(3):2298–2310, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ma:2023:GLF

[670] Ting Fung Ma, Fangfang Wang, and Jun Zhu. On generalized latent factor modeling and inference for high-dimensional binomial data. Biometrics, 79(3):2311–2320, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

LeBlanc:2023:MSL

[671] Patrick LeBlanc and Li Ma. Microbiome subcommunity learning with logistic-tree normal latent Dirichlet allocation. Biometrics, 79(3):2321– 2332, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2023:IBH

[672] Yi Zhao, Bingkai Wang, Chin-Fu Liu, Andreia V. Faria, Michael I. Miller, Brian S. Caffo, and Xi Luo. Identifying brain hierarchical structures associated with Alzheimer’s disease using a regularized regression method with tree predictors. Biometrics, 79(3):2333–2345, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2023:HWC

[673] Ruoqi Yu. How well can fine balance work for covariate balancing. Biometrics, 79(3):2346–2356, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chang:2023:CCE

[674] Changgee Chang, Zhiqi Bu, and Qi Long. CEDAR: communication efficient distributed analysis for regressions. Biometrics, 79(3):2357–2369, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jiang:2023:SIP

[675] Zhichao Jiang, Kosuke Imai, and Anup Malani. Statistical inference and power analysis for direct and spillover effects in two-stage randomized experiments. Biometrics, 79(3):2370–2381, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2023:EAU

[676] Bing Li, Constantine Gatsonis, Issa J. Dahabreh, and Jon A. Steingrimsson. Estimating the area under the ROC curve when transporting a prediction model to a target population. Biometrics, 79(3):2382–2393, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2023:IDR

[677] Shih-Hao Huang, Kerby Shedden, and Hsin wen Chang. Inference for the dimension of a regression relationship using pseudo-covariates. Biometrics, 79(3):2394–2403, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ren:2023:CEN

[678] Mingyang Ren, Sanguo Zhang, and Junhui Wang. Consistent estimation of the number of communities via regularized network embedding. Biometrics, 79(3):2404–2416, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tu:2023:AAL

[679] Danni Tu, Manu S. Goyal, Jordan D. Dworkin, Samuel Kampondeni, Lorenna Vidal, Eric Biondo-Savin, Sandeep Juvvadi, Prashant Raghavan, Jennifer Nicholas, Karen Chetcuti, Kelly Clark, Timothy Robert-Fitzgerald, Theodore D. Satterthwaite, Paul Yushkevich, Christos Davatzikos, Guray Erus, Nicholas J. Tustison, Douglas G. Postels, Terrie E. Taylor, Dylan S. Small, and Russell T. Shinohara. Automated analysis of low-field brain MRI in cerebral malaria. Biometrics, 79(3):2417–2429, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:HDM

[680] Jade Xiaoqing Wang, Yimei Li, Wilburn E. Reddick, Heather M. Conklin, John O. Glass, Arzu Onar-Thomas, Amar Gajjar, Cheng Cheng, and Zhao-Hua Lu. A high-dimensional mediation model for a neuroimaging mediator: Integrating clinical, neuroimaging, and neurocognitive data to mitigate late effects in pediatric cancer. Biometrics, 79(3):2430–2443, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:LSS

[681] Qinxia Wang, Ji Meng Loh, Xiaofu He, and Yuanjia Wang. A latent state space model for estimating brain dynamics from electroencephalogram (EEG) data. Biometrics, 79(3):2444–2457, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2023:BTS

[682] Juhee Lee, Peter F. Thall, and Pavlos Msaouel. Bayesian treatment screening and selection using subgroup-specific utilities of response and toxicity. Biometrics, 79(3):2458–2473, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Das:2023:BHQ

[683] Priyam Das, Christine B. Peterson, Yang Ni, Alexandre Reuben, Jiexin Zhang, Jianjun Zhang, Kim-Anh Do, and Veerabhadran Baladandayuthapani. Bayesian hierarchical quantile regression with application to characterizing the immune architecture of lung cancer. Biometrics, 79(3): 2474–2488, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Turchetta:2023:BSS

[684] Armando Turchetta, Erica E. M. Moodie, David A. Stephens, and Sylvie D. Lambert. Bayesian sample size calculations for comparing two strategies in SMART studies. Biometrics, 79(3):2489–2502, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diana:2023:FBI

[685] Alex Diana, Emily Beth Dennis, Eleni Matechou, and Byron John Treharne Morgan. Fast Bayesian inference for large occupancy datasets. Biometrics, 79(3):2503–2515, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stijven:2023:CCI

[686] Florian Stijven, Johan Verbeeck, and Geert Molenberghs. Comparing COVID-19 incidences longitudinally per economic sector against the background of preventive measures and vaccination. Biometrics, 79(3):2516– 2524, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tan:2023:ARM

[687] Jianbin Tan, Ye Shen, Yang Ge, Leonardo Martinez, and Hui Huang. Agerelated model for estimating the symptomatic and asymptomatic transmissibility of COVID-19 patients. Biometrics, 79(3):2525–2536, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stoner:2023:CDR

[688] Oliver Stoner, Alba Halliday, and Theo Economou. Correcting delayed reporting of COVID-19 using the generalized-Dirichlet-multinomial method. Biometrics, 79(3):2537–2550, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maleyeff:2023:AET

[689] Lara Maleyeff, Fan Li, Sebastien Haneuse, and Rui Wang. Assessing exposure-time treatment effect heterogeneity in stepped-wedge cluster randomized trials. Biometrics, 79(3):2551–2564, September 2023. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frieri:2023:DCT

[690] Rosamarie Frieri, William Fisher Rosenberger, Nancy Flournoy, and Zhantao Lin. Design considerations for two-stage enrichment clinical trials. Biometrics, 79(3):2565–2576, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Montoya:2023:ERA

[691] Lina M. Montoya, Michael R. Kosorok, Elvin H. Geng, Joshua Schwab, Thomas A. Odeny, and Maya L. Petersen. Efficient and robust approaches for analysis of sequential multiple assignment randomized trials: Illustration using the ADAPT-R trial. Biometrics, 79(3):2577–2591, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hoskovec:2023:IHM

[692] Lauren Hoskovec, Matthew D. Koslovsky, Kirsten Koehler, Nicholas Good, Jennifer L. Peel, John Volckens, and Ander Wilson. Infinite hidden Markov models for multiple multivariate time series with missing data. Biometrics, 79(3):2592–2604, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2023:SDM

[693] Fangya Mao and Richard J. Cook. Spatial dependence modeling of latent susceptibility and time to joint damage in psoriatic arthritis. Biometrics, 79(3):2605–2618, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2023:SDL

[694] Yuyan Wang, Akhgar Ghassabian, Bo Gu, Yelena Afanasyeva, Yiwei Li, Leonardo Trasande, and Mengling Liu. Semiparametric distributed lag quantile regression for modeling time-dependent exposure mixtures. Biometrics, 79(3):2619–2632, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2023:MRH

[695] Yuxin Zhu, Zheyu Wang, and David Newman-Toker. Misdiagnosis-related harm quantification through mixture models and harm measures. Biometrics, 79(3):2633–2648, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shepherd:2023:MVS

[696] Bryan E. Shepherd, Kyunghee Han, Tong Chen, Aihua Bian, Shannon Pugh, Stephany N. Duda, Thomas Lumley, William J. Heerman, and Pamela A. Shaw. Multiwave validation sampling for error-prone electronic health records. Biometrics, 79(3):2649–2663, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2023:PCP

[697] Laura Y. Zhou, Fei Zou, and Wei Sun. Prioritizing candidate peptides for cancer vaccines through predicting peptide presentation by HLA-I proteins. Biometrics, 79(3):2664–2676, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2023:NNI

[698] Tao Sun and Ying Ding. Neural network on interval-censored data with application to the prediction of Alzheimer’s disease. Biometrics, 79(3): 2677–2690, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2023:DSC

[699] Duncan Lee and Craig Anderson. Delivering spatially comparable inference on the risks of multiple severities of respiratory disease from spatially misaligned disease count data. Biometrics, 79(3):2691–2704, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Little:2023:ASM

[700] Paul Little, Li Hsu, and Wei Sun. Associating somatic mutation with clinical outcomes through kernel regression and optimal transport. Biometrics, 79(3):2705–2718, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bothwell:2023:PBC

[701] Samantha Bothwell, Alex Kaizer, Ryan Peterson, Danielle Ostendorf, Victoria Catenacci, and Julia Wrobel. Pattern-based clustering of daily weighin trajectories using dynamic time warping. Biometrics, 79(3):2719–2731, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2023:LMM

[702] Wei Zhang, Simon J. Bonner, and Rachel S. McCrea. Latent multinomial models for extended batch-mark data. Biometrics, 79(3):2732–2742, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Axelrod:2023:SAA

[703] Rachel Axelrod and Daniel Nevo. A sensitivity analysis approach for the causal hazard ratio in randomized and observational studies. Biometrics, 79(3):2743–2756, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hengelbrock:2023:HPU

[704] Johannes Hengelbrock, Johannes Rauh, Jona Cederbaum, Maximilian Kähler, and Michael Höhle. Hospital profiling using Bayesian decision theory. Biometrics, 79(3):2757–2769, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liao:2023:BRP

[705] Chen-Po Liao. Book review: Probability and random variables: theory and applications. By Iickho Song, So Ryoung Park, Seokho Yoon (2022). Springer Cham. ISBN: 978-3-030-97678-1; 978-3-030-97679-8 (eBook). Biometrics, 79(3):2770, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2023:BRM

[706] Chia-Yen Chen. Book review: Mendelian randomization: methods for causal inference using genetic variants, 2nd edition By Stephen Burgess and Simon G. Thompson. New York : Chapman & Hall. Biometrics, 79(3):2771–2772, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2023:BRF

[707] Li-Pang Chen. Book review: Fundamentals of high-dimensional statistics: with exercises and R Labs. By Johannes Lederer, Springer International Publishing, 2021. pp. 355. ISBN: 978-3-030-73791-7. Biometrics, 79(3): 2772–2773, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).